### v1

# Wake Vision: A Large-scale, Diverse Dataset and Benchmark Suite for TinyML Person Detection

Abstract
Machine learning applications on extremely low-power devices, commonly referred to as tiny machine learning (TinyML), promises a smarter and more connected world. However, the advancement of current TinyML research is hindered by the limited size and quality of pertinent datasets. To address this challenge, we introduce Wake Vision, a large-scale, diverse dataset tailored for person detection—the canonical task for TinyML visual sensing. Wake Vision comprises over 6 million images, which is a hundredfold increase compared to the previous standard, and has undergone thorough quality filtering. Using Wake Vision for training results in a 2.41% increase in accuracy compared to the established benchmark. Alongside the dataset, we provide a collection of five detailed benchmark sets that assess model performance on specific segments of the test data, such as varying lighting conditions, distances from the camera, and demographic characteristics of subjects. These novel fine-grained benchmarks facilitate the evaluation of model quality in challenging real-world scenarios that are often ignored when focusing solely on overall accuracy. Through an evaluation of a MobileNetV2 TinyML model on the benchmarks, we show that the input resolution plays a more crucial role than the model width in detecting distant subjects and that the impact of quantization on model robustness is minimal, thanks to the dataset quality. These findings underscore the importance of a detailed evaluation to identify essential factors for model development. The dataset, benchmark suite, code, and models are publicly available under the CC-BY 4.0 license, enabling their use for commercial use cases.
Keywords:
Dataset Person Detection TinyML Benchmarks
**footnotetext: These authors contributed equally to this work
1 Introduction
To reduce the energy consumption and cost of deploying Machine Learning (ML) deployments, researchers have pioneered a new field of ultra-low-power machine learning (TinyML) [3, 35, 1, 8]. TinyML aims to co-locate ML models with the sensors by deploying the model onto always-on microcontrollers (MCUs) and TinyML accelerators, thereby reducing the energy cost of transmitting data or constantly running larger processors. Deploying models on these tiny devices imposes severe constraints, typically confining the model’s size to a few hundred kilobytes, nearly four orders of magnitude less memory than mobile phones [2].
The field of TinyML requires large-scale datasets to facilitate high-quality research. However, the extreme constraints imposed by the hardware necessitate that TinyML models be compact and efficient, rendering them unsuitable for traditional, complex ML tasks. As a result, conventional ML datasets like ImageNet [9], designed for more resource-intensive models, are often ill-suited for TinyML research endeavors. While existing TinyML datasets [7, 34] have been valuable resources, they tend to be limited in scale, making it challenging to train production-grade TinyML models solely on these datasets.
To enable and mature the field of TinyML research, we present the Wake Vision dataset, a large dataset for person detection, the quintessential vision use case for TinyML [1]. Person detection is a binary image classification task where a model detects whether a person is present. While simple, this task has many important use cases, such as occupancy detection [29], smart HVAC/lighting [36], or acting as an always-on ‘wake model’ in a larger ML system [14].
Wake Vision has over 6M images, ~100x more than the prior state-of-the-art person detection dataset, Visual Wake Words (VWW) [7]. Wake Vision is derived from the Open Images v7 dataset [21, 19] and is permissively licensed. Wake Vision demonstrates a 2.41% accuracy improvement over VWW.
Due to the limited capacity of TinyML models, data quality is critical [26], however techniques, such as knowledge distillation [15], can leverage large, noisy datasets more effectively. For this reason, we choose to release two training sets: one that prioritizes dataset size, Wake Vision (Large), and one that prioritizes data quality, Wake Vision (Quality). We find that Wake Vision (Quality) out performs Wake Vision (Large), but we release both training sets to enable research into the tradeoffs between quantity and quality of data for TinyML.
Person detection systems are deployed, and expected to work, in challenging settings, such as low lighting. However, because these challenging settings are uncommon in images sourced from the internet, high-level metrics, like test set accuracy, may overestimate a model’s real world performance. To address this issue we introduce a suite of five fine-grain benchmarks, which contain only images of a particular demographic or scenario. Our benchmarks cover distance, lighting, depictions, and perceived gender and age. Finally, we use our benchmark suite to conduct a case study on the performance effects of typical TinyML compression. This analysis demonstrates the usefulness of our benchmark and can inform future TinyML model design.
In summary, we present the following novel contributions:
  • A new person detection dataset, called Wake Vision, which is an easily accessible, permissively licensed dataset that is ~100x larger than the prior standard.
  • A suite of five fine-grain benchmark sets that evaluate a model’s fairness and robustness in challenging settings.
  • A case study on the effect of TinyML scaling the highlights the importance of fine grain analysis.
2 Related Work
The current de facto standard dataset for person detection is the VWW dataset [7, 8, 23, 2, 1]. To our knowledge, this is the only current open-source dataset explicitly meant for person detection. Unfortunately, the VWW dataset is small and cannot be directly downloaded. Instead, it must be regenerated from MS-COCO [24], thus limiting its accessibility and usefulness.
Apart from dedicated person detection datasets, some general image classification datasets contain a person label, which makes it possible to create person detection models based on them. Examples include the Cifar-100 dataset [20], and the PASCAL Visual Object Classes dataset [11]. However, using general image classification datasets can lead to poor perceived performance of TinyML models, as they do not adequately represent the open “no-person” class [7].
Tab. 1 shows that our Wake Vision dataset contains up to almost two orders of magnitude more images than any other public and permissively licensed dataset. Furthermore, Wake Vision is the only person detection dataset to provide a fine-grained benchmark suite and be officially distributed to popular dataset services such as TensorFlow Datasets (TFDS) and Hugging Face Datasets.
Furthermore, Wake Vision integrates fine-grained benchmarks, which are a novel addition to existing vision data sets for TinyML. These fine-grained benchmark sets, which include distance, lighting, depictions, and perceived gender and age variations, allow for a deeper investigation into the effects of input and model compression on specific demographic or scenario subsets. Several research papers report little performance degradation from model compression on top-line metrics such as top-1 test set accuracy [6, 13, 10, 18]. However, recent research suggests that this focus on top-line metrics can hide poor performance on critical subsets of a dataset, e.g., subsets relating to sensitive information and fairness [16, 17].
Table 1: A comparison of person detection and image classification datasets.
Total # of Person Images Fine-Grain Suitable for
Dataset Images Train Validation Test Filtering TinyML
Wake Vision (Quality) 1,423,690 675,411 9,106 27,328
Wake Vision (Large) 6,477,906 3,238,953 - -
Visual Wake Words [7] 123,287 36,000 3,926 19,107
CIFAR-100 [20] 60,000 2,500 - 500
PASCAL VOC 2012 [11] 11,530 1,994 2,093 -
3 Wake Vision Dataset
We introduce the Wake Vision dataset, the largest dataset for TinyML person detection. The dataset’s size, quality, and detailed metadata enable new avenues of TinyML research. Wake Vision is two orders of magnitude larger than prior comparable datasets and can be adapted based on the target use case. Fig. 1 illustrates examples of images in the dataset. The dataset is permissively licensed for research and commercial use. Images are sourced from Flickr [12] images listed as having a CC-BY 2.0 license Labels are derived from Open Images and licensed under the CC-BY 4.0 license. In this section, we discuss our methodologies for label generation, data filtering, and error correction, as well as discussing the usability of the dataset and directly comparing it to the Visual Wake Dataset.
Older Person
Near Person
Bright Image
Female Person
Depicted Person
Young Person
Figure 1: Sample images from the Wake Vision fine-grained benchmark suite.
3.1 Label Generation
While a sufficiently large person detection dataset is indispensable for TinyML research, it can be challenging to obtain. Manually labelling millions of images would be prohibitively expensive for a nascent research field like TinyML.
Therefore, we leverage existing large-scale data efforts to to generate Wake Vision automatically. The base label in Wake Vision is a binary person/non-person label. We derive these labels from Open Images, which contain both image-level and bounding box labels. Image-level labels describe which objects are present in an image but not localized to a specific point in the image.
An image can contain several image-level labels, where as bounding box labels provide a set of four coordinates forming a box around the object. The bounding box label classes of Open Images are hierarchically structured so that one class can be a subcategory or a part of another class, e.g., a “Woman” is a subcategory of a “Person,” and a “Human Hand” is a part of “Person.”
We observe that the bounding box labels are generally more accurate and provide more information for data filtering but are less numerous. The Open Images training set has over 3 million images with image-level person labels, but only ~ 700,000 images with a person bounding box. The Open Images validation and test sets are fully labeled with image-level and bounding box labels. Consequently, this presents a trade-off between more data (image-level labels) or higher-quality labels (bounding box labels), as different use cases may have different requirements for the training data (e.g., active learning).
We provide two Wake Vision training sets: Wake Vision (Large) and Wake Vision (Quality). One larger set, labeled via Open Images image-level labels, and one smaller set, labeled via Open Images bounding box labels (Tab. 1). In our experiments, the Wake Vision (Quality) training set produce more accurate models than the Wake Vision (Large) training set, strongly suggesting that label quality is more important than quantity in our setting.
3.1.1 Large Scale Training Set from Image-Level Labels:
Image-level labels in Open Images contain little additional information in addition to the labels. The only additionally relevant information is a confidence property, which represents how certain it is that a label is correct. The confidence ranges from 0-1 for labels generated by machines, and is either 0 or 1 for labels negatively or positively verified by humans. We use this property to exclude images from Wake Vision if they contain only low confidence person labels.
3.1.2 High Quality Training Set From Bounding Box Labels:
Bounding Box labels in Open Images, in contrast to image-level labels, are all verified and localized by humans. This ensures that bounding box labels are less likely to be false positives. Bounding box labels can also be used to calculate an approximation of the area taken up by a person. We use this as a proxy for the distance of a person in the image, and exclude images where all person labels are far away, i.e., take up only a small portion of the image. An example of an image with a far away person can be found in Fig. 2(a). Finally, bounding box attributes include a depiction flag, which represents whether a label is a depiction. See an example of an image containing a person with the depiction flag set in Fig. 2(b). By default we use this flag to make sure that we do not consider a depiction a person.
We allow users to adapt the Wake Vision dataset to meet the needs of specific usecases by adding several configurable options in the open-source dataset creation script. A comprehensive list of Open Images labels related to Wake Vision can be found in Appendix 0.I.
(a) An example of a person far away
(b) An example of a depicted person
Figure 2: Examples of challenging outlier images.
3.1.3 Training Set Comparison
To compare the performance of the two training dataset sets, we train a MobileNetV2-0.25 [31] model on each dataset for an equal number of steps and compare its performance on the common test set (Tab. 2). Given the amount of label errors in the Wake Vision test set shown in Tab. 4, a perfect person detection model would achieve an accuracy of ~93%. The filtered training set generated from the bounding box labels outperforms the larger training set generated from the image-level labels by 2.65% test accuracy. This indicates that label quality is more important than quantity in this setting. Interestingly, although the gap between the two is reduced to just 0.36% when training exclusively on soft labels from a teacher model, the bonding box training set is still superior. This indicates that the difference in quality extends beyond just the labels and that our data filtering leads to an improvement on it’s own.
Table 2: Accuracy on Wake Vision Test set of a MobileNetV2-0.25 model trained on image-level and bounding box labels, with and without distillation. A MobileNetV2-1.0 model trained on the bonding box set is used as the teacher in distillation.
Training set Label Source Dataset Size Accuracy Distilled Accuracy
Wake Vision (Large) Image-Level 6,477,906 86.00% 88.96%
Wake Vision (Quality) Bounding Box 1,350,822 88.65% 89.32%
Despite the worse performance in this setting, the larger image-level training set may be useful for training higher capacity models for longer or when doing active learning [30] or distillation [5]. As our distillation results indicate, further data-centric filtering techniques may be needed to fully leverage the image-level training set. We experimented with tripling the model size and doubling the training time, but the difference in accuracy remained relatively consistent.
The two training sets provide a useful foundation for future work on the importance of data quality vs. quantity. Wake Vision can enable further research on data-centric TinyML [25, 26], specifically uncovering the relationship between model capacity and the sensitivity to training data quality.
3.2 Error Correction
Labeling errors are a challenge in computer vision [28, 4] that limit progress and obscures true performance. This is especially prevalent in large datasets that use complex labeling pipelines to scale while managing costs. Recognizing the importance of label quality, we take measures to estimate and improve the accuracy of its labels, aligning with best practices in the field of computer vision.
Wake Vision’s labels are derived from the Open Images labels; therefore, the errors that occur in the Open Images labeling pipeline are inherited by Wake Vision. All labels in the Open Images dataset are originally machine-generated candidate labels. Many of these labels are later verified by human annotators, and a subset of them are given bounding boxes and additional annotations.
While the machine-generated label phase aims to identify objects efficiently, any instances missed during this initial step are unlikely to be captured in downstream phases. Consequently, the Open Images dataset may contain numerous false negative labels, referring to images where an object is present, but lacks the corresponding label. Therefore, additional measures are necessary to identify and correct such labeling omissions to ensure the dataset’s integrity.
Given the substantial scale of Wake Vision, manually correcting label errors across the entire dataset becomes an arduous undertaking. To address this challenge, we employ the Confident Learning technique [27] to intelligently identify potential label errors. We selected Confident Learning as prior work has demonstrated its capability to find label errors in large datasets [28]. Once these suspected mislabeled instances are flagged, we methodically inspect and correct them through a manual verification process.
Table 3: Amount of label errors identified and corrected using Confident Learning.
Dataset Total Suggested Corrected Suggestion
Split Size Errors Errors Accuracy
Validation 16,964 632 81 12.82%
Test 51,282 1672 267 15.97%
As shown in Tab. 3, the confident learning process identified a large amount of possible label errors in Wake Vision’s Validation and Test sets. However, only between 12 and 16% of these possible label errors were legitimate errors. We correct these identified label errors in the Wake Vision validation and test sets.
Given this low acceptance rate of label issues identified by Confident Learning, we conclude that we cannot automate label cleaning to a point where no human in the loop is needed. Based on that we also conclude that this strategy too human-intensive to be applied to the much larger training sets.
We report our estimated label error rate after label corrections versus VWW in Tab. 4. We find that after our corrections, the VWW dataset has a top-level error rate of 7.8%, which is comparable to Wake Vision’s top-level error rate of 6.8%. Notably, about a third of the Wake Vision label errors stem from depictions missing the depiction flag. As we by default do not consider depictions as persons, missing this flag results in a label error. Not considering the depiction label errors, Wake Vision has closer to half the number of errors of VWW.
VWW does not, to the authors knowledge, include depictions. Thus no depiction errors are present for the VWW dataset. In real world settings, depictions are everywhere, therefore we expect that Wake Vision will prove to be a more realistic dataset. ML systems based on Wake Vision will therefore likely be able to handle real world scenarios better than equivalent systems based on VWW.
We are in the process of crowd-sourcing label corrections for the validation and test sets to eliminate the remaining errors. However, despite these efforts, Wake Vision, in its current state, already offers a significant improvement over the VWW dataset, as detailed in Section 3.4. Although further enhancements to label quality are underway, Wake Vision presents a valuable and directly improved resource compared to its predecessor VWW dataset.
Table 4: Percentage (and count) of label errors in a 500 sample subset of the VWW and Wake Vision test sets. Percent of label errors caused by misclassified depictions are reported.
Dataset Label Error Rate % Errors From Depictions
Visual Wake Words 7.8% (39) 0% (0)
Wake Vision 6.8% (34) 38.2% (13)
3.3 Usability
Wake Vision is available through TensorFlow Datasets [33] and HuggingFace Datasets [22] to be easily accessed and used by the community. The images and labels are rehosted to ensure the dataset will not be changed over time due to dead links, making it a more viable benchmark. The rehosted labels are generated according to our default dataset configuration, further described in Appendix 0.F.
3.4 Comparison to Visual Wake Words
Wake Vision serves as a drop-in improvement over VWW, thanks to its size and our extensive filtering. To compare the benefits of Wake Vision against VWW, we train two identical models using the same recipe for the same number of steps, one utilizing VWW’s training set and the other employing Wake Vision’s training set. For this comparative evaluation, both trained models are MobileNetV2 architectures with a width modifier of 0.25, which have previously served as benchmark TinyML models [7]. After training these identical models using the respective datasets’ training sets with the same recipe for an equal number of steps, we cross-evaluate their performance on the corresponding test sets.
In Tab. 5 we demonstrate a direct 0.5% improvement on VWW’s own test set, which illustrates that Wake Vision is a direct improvement over VWW, not simply a domain shift. We additionally achieve a 2.41% improvement over VWW on Wake Vision’s Test set, indicating that the new test set is more challenging.
Table 5: Accuracy on the Wake Vision and VWW test sets by models trained on each dataset’s respective training set.
Train
VWW Wake Vision
Test VWW 88.00% 88.50%
Wake Vision 86.27% 88.68%
4 Fine Grain Benchmark Suite
Test sets generated directly from images posted to the internet are not reflective of real world use cases as the distribution is biased towards images people deems as worth sharing (e.g. well lit and framed). In contrast, person detection systems are deployed, and expected to work, in more challenging settings, such as low lighting. Due to this gap, a model that achieves high test accuracy may end up performing poorly in real world scenarios, after the design stage has ended.
To address this issue, we present a benchmark suite for TinyML person detection that tests a model’s robustness in challenging settings, enabling better analysis at the design stage. Additionally the benchmark suite measures the model’s performance across demographics, to ensure the model does not exhibit clear bias. The suite consists of five fine-grain benchmark sets, which we elaborate on below. Each of the five fine-grained benchmarks have been chosen based on a combination of its relevance to TinyML use cases and the availability of requisite metadata to generate the sets. Each benchmark set is a subset of the validation or test set filtered based on the criteria under test (e.g. persons far away from the camera).
The typical usage of the benchmark is to determine if a model is sufficiently accurate in the planned deployment setting. For example, a model designer may make different design choices for a use case where the subject is close to the camera and well lit, vs. the inverse setting. In this section, we describe the design of each individual benchmark and use the suite to compare a VWW model to a Wake Vision model. Tab. 6 reports the number of samples in each of the individual benchmark sets and the F1 scores of a Wake Vision and VWW model on each set.
4.1 Perceived Gender and Age
The Perceived Gender and Age fine grained datasets are generated from the Open Images More Inclusive Annotations for People (MIAP) extended fairness labels [32]. Underrepresented sub-groups typically constitute a small portion of a generic test set; therefore, top-line metrics often obscure a bias in a model until it is deployed [16]. This benchmark aims to evaluate a model separately on demographics that are underrepresented in the underlying dataset distribution, identifying bias. These labels are based on a perceived gender and age representation, and are not necessarily representative of a persons gender identity or age. Like the authors of the original MIAP labels, we do not support or condone the creation or deployment of gender or age specific classifiers based on this dataset.
Table 6: Wake Vision Fine Grain Benchmark Suite. We report the samples in each set as well as the F1 score of a Wake Vision model and a VWW model on the test set.
Set Size F1-Score
Benchmark Val Test Wake Vision VWW
Gender Female 675 2,152 0.94 0.89
Male 1,238 3,748 0.93 0.89
Unknown 1,718 5,231 0.78 0.79
Age Young 271 875 0.94 0.90
Middle 2,035 6,371 0.93 0.89
Older 91 271 0.96 0.90
Unknown 1,388 4,120 0.73 0.76
Distance Near 5,888 17,536 0.90 0.88
Medium 2,304 7,168 0.83 0.83
Far 1,152 3,328 0.39 0.49
Lighting Dark 3,036 8,872 0.88 0.83
Normal 12,992 39,327 0.88 0.86
Bright 936 3,083 0.85 0.85
Depictions Person 179 501 0.83 0.77
Non-Person 313 957 0.89 0.85
No Depiction 7,990 24,183 0.90 0.87
4.2 Distance
The distance fine grained datasets aim to test how the distance of persons in images impact the performance of ML systems. This benchmark is critical for correctly predicting a model’s deployed performance. If a person detection system is intended to recognize subjects at great distances, the system’s performance on the far away dataset will be more informative than its performance on the top-level test set. We create three datasets based on the percentage of the image the subject bounding box covers. The three sets are near (>60%), at a medium distance (10-60%), and far away (<10%).
4.3 Lighting
The lighting fine grained datasets aim to test the performance of ML systems across different lighting conditions. For instance, in scenarios such as surveillance cameras, outdoor robotics, or augmented reality applications, models need to be robust to varying lighting conditions, including low-light environments. As with distance, this benchmark provides insight into possible causes of significant performance degradation once deployed, such as low-light settings. We create three fine grained datasets of this type for dark, normal and bright lighting conditions respectively. All images in the original Wake Vision Vision validation and test set are classified into these datasets according to their lighting condition. We quantify lighting conditions by the average pixel values of images in greyscale, which is a simple but effective method for distinguishing lighting conditions [37]. We define low as an average pixel value less than 85, normal as between 85 and 170, and bright lighting conditions as greater than 170.
4.4 Depictions
A particularly challenging task for a person detection model is to correctly reject depictions of people. In many use cases a person detection model can not falsely trigger on a depiction. For example, a room occupancy detector that incorrectly identifies a painting on the wall as a person. This benchmark measures the models accuracy on three related sets of non-person samples: depictions of people, depictions that are not of people, and images do not contain a depiction of any kind. Depictions of people can range from photo-realistic to crude stick figures.
4.5 Benchmark Results
Tab. 6 compares the results of a model trained on the Wake Vision (Quality) training set to an identical model trained on VWW. The Wake Vision model exhibits superior robustness across the challenging settings exercised by our benchmarking suite, further demonstrating the effectiveness of the dataset. For instance, on the “Depictions” benchmark, which evaluates performance on images containing persons, non-person objects, or no depictions, the Wake Vision model achieves an F1 score of 0.83 for person depictions, outperforming the VWW model’s 0.77. Similarly, for non-person depictions, the Wake Vision model scores 0.89 compared to the VWW model’s 0.85.
The Wake Vision model also showcases improved performance on the “Age” benchmark, with F1 scores of 0.94 for young subjects, 0.93 for middle-aged, and a notable 0.96 for older individuals, surpassing the VWW model’s scores of 0.90, 0.89, and 0.90, respectively. This highlights the Wake Vision dataset’s effectiveness in enhancing model robustness for detecting people across various age groups, particularly the elderly demographic. In addition to age and depictions, the Wake Vision model demonstrates resilience to challenging lighting conditions, achieving F1 scores of 0.88 for both dark and normal lighting scenarios, compared to the VWW model’s 0.83 and 0.86, respectively.
Collectively, these results underscore the significance of Wake Vision’s large-scale, diverse data and fine-grained benchmarks, enabling the development of more robust and reliable TinyML person detection models across a wide range of realistic and real-world scenarios.
Figure 3: Effects of scaling the image size vs. the model width on Wake Vision test accuracy and the Distance-Far benchmark set. The Distance-Far F1 score (right) is far more sensitive to changes in the image size than the overall test accuracy (left). Image sizes refer to square image resolutions (e.g. 96x96) and model size refers to the MobileNetV2 width multiplier. When not specified the image size is 160 and the width multiplier is 1.0.
5 Model Design Case Study
Basic test set performance can often misrepresent a model’s performance, given the typical domain shift between images scraped from the internet and real-world use cases. This issue is exacerbated when ML practitioners must trade off accuracy for model performance and size, which is necessary for TinyML use cases. A design decision might have seemingly little impact on the test accuracy but may destroy real-world performance depending on the deployment environment. For example, a person detection system may only operate in dark lighting conditions, but the test dataset has an insignificant number of dark samples; therefore the test accuracy will not reflect the real world accuracy.
The benchmark suite enables more holistic analysis during the design phase. To show this use we perform a series of scaling experiments employing typical TinyML compression techniques and identifying under which circumstances these techniques are appropriate. While these results can inform ML practitioners, our intention is to demonstrate the usefulness of the benchmark suite.
5.1 Image Size vs. Model Width Scaling
We train two series of MobileNetV2 models: one series that sweeps the input image size [64-256] and one that sweeps the width multiplier of a model [0.1-1.5]. We then benchmark these models on the Wake Vision test set as well as the far distance benchmark. We plot these results against the number of multiply accumulate (MAC) operations in the model as a proxy for on-device latency [2].
The results in Fig. 3 (left) show that when looking exclusively at the high-level metric (i.e., test accuracy), scaling the input image size has a similar impact as scaling the model size. However, as shown in Fig. 3 (right), when we consider only samples where the person is far away from the camera (i.e., the distance benchmark), we observe a much more significant impact when scaling the image size. In the case of distant subjects, the image size becomes the bottleneck.
These findings suggest that for ML developers targeting use cases where the subject is likely to be far from the camera, prioritizing larger input image sizes over wider models may be more beneficial. However, this critical design consideration could be obscured when solely relying on high-level metrics like overall test accuracy. The distance benchmark in Wake Vision effectively unveils the disproportionate impact of image size on model performance for distant subjects, enabling more informed decision-making during model optimization.
5.2 Quantization
Quantization is a crucial technique for deploying efficient TinyML models, offering substantial benefits in terms of reduced latency, memory footprint, and model size. However, prior work has suggested that quantization can disproportionately impact the performance of models on underrepresented subsets of data [17]. To assess the implications of quantization in the context of person detection, we investigate the impacts of int8 quantization on a model’s benchmark results across Wake Vision’s fine-grained benchmarks.
[…]
5.3 Grayscale
Converting a model’s input image channels to grayscale from RGB is a commonly employed optimization in the TinyML field [2] as it can substantially reduce a model’s memory consumption. We observed, however, that the grayscale optimization disproportionately impacts images on the brighter end of the spectrum as illustrated in Fig. 4. This further demonstrates the importance of fine grained analysis, as some real world deployment environments might be far brighter than the average Wake Vision test sample.
Figure 4: Impact of grayscale input images on the lighting benchmarks: dark, normal light, and bright. Models that use grayscale input images are more sensitive to bright lighting conditions than RGB.
6 Ethical Considerations
The primary purpose of the Wake Vision dataset and benchmark suite is to mature the field of TinyML by giving researchers access to large-scale datasets and enabling fine-grain analysis. Beyond this purpose, Wake Vision can directly improve the quality of person detection systems that benefit society, such as saving energy by only turning on the lights in a room when a person is present. Additionally, TinyML person detection protects user privacy by running inference on-device and not transmitting private data over the network. However, it should be noted that person detection models can also be used for malicious purposes, such as weapons targeting systems or large-scale surveillance.
The images that make up Wake Vision are entirely sourced from Flickr. While we attempt to only use images under the consent given by the CC-BY 2.0 license (by sourcing images through Open Images[21, 19]), we make no representations or warranties regarding the license status of each image. It is possible that some of the millions of images have been uploaded to Flickr by a user without the right to distribute the images with a CC-BY 2.0 license.
Futhermore, the benchmark suite is intended to ensure the fairness and robustness of person detection models. While the goal of these benchmark suites is to build more robust and fair models, our demographic benchmarks could be misused to create systems to classify gender and age, potentially breaching personal privacy or perpetuating discrimination. We have attempted to mitigate this risk by not releasing training sets for the fine-grained benchmark datasets.
7 Conclusions
The field of TinyML research is currently bottlenecked by the lack of large, high-quality datasets. To alleviate this bottleneck we introduced Wake Vision—a large scale dataset and benchmark suite for person detection. Compared to VWW, the current state of the art dataset for person detection, Wake Vision is ~100x larger, easier to use, has a comparable amount of label errors, and is more reflective of real world use cases. These factors contribute to Wake Vision being a drop-in improvement over VWW. Fairness across demographics and robustness to challenging settings is essential for many person detection use cases. To facilitate evaluation of this, we released a Wake Vision benchmark suite of six fine-grained benchmark sets to test the fairness and robustness of person detection models. Lastly we demonstrate how the Wake Vision benchmark suite can be used to investigate the effects of different types of input and model scaling and highlight the importance of fine grain analysis when designing an model.
Appendix 0.A Anonymized Repo
The code used for the paper is available at this anonymous repository: https://anonymous.4open.science/r/Wake_Vision-0D15/README.md
This repo contains the code to generate Wake Vision and the benchmark suite, as well as the code to train and evaluate models. This code is sufficient to reproduce all results in the paper.
Appendix 0.B Flowchart of Bounding Box Filtering Process
Fig. 5 illustrates the filtering process for Wake Vision when using Open Images’ bounding box labels as the label source.
Figure 5: A flowchart of the filtering of an image according to default Bounding Box filtering rules.
Appendix 0.C Fine Grained Benchmark Images
Fig. 6 gives example images for each of the benchmarks in the suite.
Gender
Female
Male
Gender Unknown
Age
Young
Middle
Older
Age Unknown
Distance
Near
Medium
Far
Lighting
Dark
Normal
Bright
Depictions
Person
Non-Person
No Depiction
Figure 6: Images from each fine grained benchmark dataset.
Appendix 0.D Open Image Label Distribution
Wake Vision is derived from Open Images V7, therefore the diversity of subjects in the images should follow the distribution of labels in Open Images. The Label distribution of Open images can be found here: https://storage.googleapis.com/openimages/web/factsfigures_v7.html#statistics
Appendix 0.E Dataset Access and Organization
Wake Vision is available directly online, via TFDS, or Hugging Face Datasets. In order to maintain author anonymity, we do not link to the dataset.
The dataset is organized as a set of compressed tar files containing images, and a series of label CSVs. The label CSVs are organized such that the file name of the image is the identifying index. The person label is provided as is a flag to indicate if the image is of a depiction. The Validation and Test label CSVs also have flags that denote a sample’s inclusion into a fine-grain benchmark set (e.g. Distance-Near). This structure makes it easy to access just the required data without requiring a full download. It also ensures the dataset can be easily updated as new versions are introduced.
Appendix 0.F Label Generation Details
Person Labels. The most straightforward Open Images label classes to label as person in Wake Vision are the "person" label and its subcategories (listed in Appendix 0.I). All of these are relabelled as persons in Wake Vision. These labels are present as both image-level labels and bounding box labels.
We furthermore inspect the image level label classes for synonyms and umbrella terms for all the person related labels. This search resulted in an additional six person related label classes. These person related label classes will only be used in the image level label configuration.
Person Body Part Labels. Body parts are more challenging to relabel, as it is dependent on the use case whether a body part should be considered a person.
For example a camera that detects whether a person is inside a room to decide if the light should be switched on would want to consider body parts as a person, as this will keep the lights on even when the person is only partly in the camera frame. For waking up electronics, however, it may not make sense to consider body parts as persons. This could, e.g., mean that a computer would turn on when detecting a foot.
To cater to both use cases we include a flag in our open-source dataset creation code to set whether body parts should be considered persons. By default we consider body parts as persons.
Depictions. Open Images bounding box labels contain metadata about whether an object is a depiction, e.g., a painting or a photograph of a person. This presents the challenge of how to handle depictions. While most use cases would not consider a depiction a person, it could make sense to either exclude them to make training easier, or include them as non-persons to make a model resistant to seeing depictions when deployed.
In line with how we handle body parts, we therefore include a flag for our open-source dataset creation code to set whether depictions should be excluded or considered non-persons. By default depictions are considered non-persons.
Bounding Box Size. The VWW dataset only considered a Common Objects in Context (COCO) person to be a person if the bounding box around the person took up at least 5% of the image [7]. If the person took up less than 5% of the image, the image was excluded from the dataset. To make our dataset work as a plug in replacement for VWW, we adopt the same defaults in Wake Vision. For different requirements, users can change a configuration parameter in our open-source dataset creation code.
Appendix 0.G Standardized Evaluation
VWW has no standardized way to evaluate performance on the dataset. This makes it challenging to compare works based on the dataset, since performance difference can come down to choices outside the contribution of the work. E.g., two works could be contributing with model improvements, but use different pre-processing pipelines that skews results.
To allow for both data-centric and model-centric improvements, we provide a standard model for data-centric contributions, and a standard pre-processing pipeline for model-centric contributions. Both types of contributions are expected to use accuracy as the primary metric for the overall test and validation set, and F1-score as the primary metric for the fine grained benchmarks introduced in Sec. 4.
Therefore, for data-centric contributions we propose to use a Mobilenet v2 model with a width modifier***Also known as alpha. of 0.25 [31]. For model-centric contributions we propose the following pre-processing pipeline:
  1. 1.
    Cast image pixel value datatype from 8-bit integers into 32-bit floating points
  2. 2.
    Resize image such that the shortest side matches the model input size
  3. 3.
    Perform a center crop on the image such that the longest size matches the model input size
  4. 4.
    Normalize pixel values to between -1 and 1 sample wise
  5. 5.
    Use image tensor as input features and person label as target feature
Appendix 0.H Dataset Download
The full Open Images v7 dataset is not hosted by the dataset authors. Rather it is provided as a collection of flickr image Uniform Resource Locators (URLs) and their associated labels. As an unfortunate result of this, the dataset is not static over time as image owners can delete their images from the flickr platform. As a result we were only able to download a subset of the original Open Images v7 dataset as shown in Tab. 7.
Table 7: Number of images downloaded from Open Images v7. Download occurred between the 28th of November to the 5th of December
Train Validation Test
Downloaded 7,936,979 36,406 109,305
Errors 1,055,669 5,214 16,131
Appendix 0.I Person Label Classes
We consider the following Open Images v7 labels to be a person for the Wake Vision dataset:
  • Person
  • Woman (Subcategory of Person)
  • Man (Subcategory of Person)
  • Girl (Subcategory of Person)
  • Boy (Subcategory of Person)
  • Human body (Part of Person)
  • Human face (Part of Person)
  • Human head (Part of Person)
  • Human (Person synonym - Only in Image Level Label Configuration)
  • Female person (Woman synonym - Only in Image Level Label Configuration)
  • Male person (Man synonym - Only in Image Level Label Configuration)
  • Child (Umbrella term for Girl & Boy - Only in Image Level Label Configuration)
  • Adolescent (Umbrella term for Girl & Boy - Only in Image Level Label Configuration)
  • Youth (Umbrella term for Girl & Boy - Only in Image Level Label Configuration)
Images containing the following Open Images v7 labels and no other person related labels are excluded from the Wake Vision dataset:
  • Human eye (Part of Person)
  • Skull (Part of Person)
  • Human mouth (Part of Person)
  • Human ear (Part of Person)
  • Human nose (Part of Person)
  • Human hair (Part of Person)
  • Human hand (Part of Person)
  • Human foot (Part of Person)
  • Human arm (Part of Person)
  • Human leg (Part of Person)
  • Beard (Part of Person)
Appendix 0.J Wake Vision Dataset Size
Table 8: Amount of images in the Wake Vision dataset
Person Images Non-Person Images Excluded
Image Level Label Train Dataset 2,880,214 2,880,214 2,176,551
Bounding Box Label Train Dataset 624,115 624,115 6,688,749
Validation Dataset 8,482 8,482 19,442
Test Dataset 25,641 25,641 58,023

### v2

@@
-Wake Vision: A Large-scale, Diverse Dataset and Benchmark Suite for TinyML Person Detection
+Wake Vision: A Large-scale, High-Quality Dataset and Benchmark Suite for TinyML Applications
 Abstract
-Machine learning applications on extremely low-power devices, commonly referred to as tiny machine learning (TinyML), promises a smarter and more connected world.
-However, the advancement of current TinyML research is hindered by the limited size and quality of pertinent datasets.
-To address this challenge, we introduce Wake Vision, a large-scale, diverse dataset tailored for person detection—the canonical task for TinyML visual sensing.
-Wake Vision comprises over 6 million images, which is a hundredfold increase compared to the previous standard, and has undergone thorough quality filtering.
-Using Wake Vision for training results in a 2.41% increase in accuracy compared to the established benchmark.
-Alongside the dataset, we provide a collection of five detailed benchmark sets that assess model performance on specific segments of the test data, such as varying lighting conditions, distances from the camera, and demographic characteristics of subjects.
-These novel fine-grained benchmarks facilitate the evaluation of model quality in challenging real-world scenarios that are often ignored when focusing solely on overall accuracy.
-Through an evaluation of a MobileNetV2 TinyML model on the benchmarks, we show that the input resolution plays a more crucial role than the model width in detecting distant subjects and that the impact of quantization on model robustness is minimal, thanks to the dataset quality.
-These findings underscore the importance of a detailed evaluation to identify essential factors for model development.
-The dataset, benchmark suite, code, and models are publicly available under the CC-BY 4.0 license, enabling their use for commercial use cases.
-Keywords:
-Dataset Person Detection TinyML Benchmarks
-**footnotetext: These authors contributed equally to this work
+Tiny machine learning (TinyML), which enables machine learning applications on extremely low-power devices, suffers from limited size and quality of relevant datasets.
+To address this issue, we introduce Wake Vision, a large-scale, diverse dataset tailored for person detection—the canonical task for TinyML visual sensing.
+Wake Vision comprises over 6 million images, representing a hundredfold increase compared to the previous standard, and has undergone thorough quality filtering.
+We provide two Wake Vision training sets: Wake Vision (Large) and Wake Vision (Quality), a smaller set with higher-quality labels.
+Our results demonstrate that using the Wake Vision (Quality) training set produces more accurate models than the Wake Vision (Large) training set, strongly suggesting that label quality is more important than quantity in our setting.
+We find use for the large training set for pre-training and knowledge distillation.
+To minimize label errors that can obscure true model performance, we manually label the validation and test sets, improving the test set error rate from 7.8% in the prior standard to only 2.2%.
+In addition to the dataset, we provide a collection of five detailed benchmark sets to facilitate the evaluation of model quality in challenging real-world scenarios that are often ignored when focusing solely on overall accuracy.
+These novel fine-grained benchmarks assess model performance on specific segments of the test data, such as varying lighting conditions, distances from the camera, and demographic characteristics of subjects.
+These benchmarks enable researchers and practitioners to better understand the strengths and limitations of their models in real-world scenarios.
+Our results demonstrate that using Wake Vision for training results in a 2.49% increase in accuracy compared to the established dataset.
+We also show the importance of dataset quality for low-capacity models and the value of dataset size for high-capacity models.
+The dataset, benchmark suite, code, and models are publicly available under the CC-BY 4.0 license at https://wakevision.ai/.
+**footnotetext: These authors contributed equally to this work.11footnotetext: Work done while a visiting researcher at Harvard University.
 1 Introduction
-To reduce the energy consumption and cost of deploying Machine Learning (ML) deployments, researchers have pioneered a new field of ultra-low-power machine learning (TinyML) [3, 35, 1, 8].
-TinyML aims to co-locate ML models with the sensors by deploying the model onto always-on microcontrollers (MCUs) and TinyML accelerators, thereby reducing the energy cost of transmitting data or constantly running larger processors.
-Deploying models on these tiny devices imposes severe constraints, typically confining the model’s size to a few hundred kilobytes, nearly four orders of magnitude less memory than mobile phones [2].
-The field of TinyML requires large-scale datasets to facilitate high-quality research.
+To reduce the energy consumption and cost of deploying Machine Learning (ML) deployments, researchers have pioneered ultra-low-power machine learning (TinyML) [3, 30, 1, 6].
+TinyML co-locates ML models with the sensors by deploying the model onto always-on microcontrollers (MCUs) and TinyML accelerators, thereby reducing the energy cost of transmitting data or constantly running larger processors.
+Deploying models on these tiny devices imposes severe constraints, confining the model’s size to $\sim$100 kilobytes, nearly four orders of magnitude less memory than mobile phones [2].
+TinyML requires large-scale datasets to facilitate high-quality research.
 However, the extreme constraints imposed by the hardware necessitate that TinyML models be compact and efficient, rendering them unsuitable for traditional, complex ML tasks.
 As a result, conventional ML datasets like ImageNet [7], designed for more resource-intensive models, are often ill-suited for TinyML research endeavors.
-While existing TinyML datasets [7, 34] have been valuable resources, they tend to be limited in scale, making it challenging to train production-grade TinyML models solely on these datasets.
-To enable and mature the field of TinyML research, we present the Wake Vision dataset, a large dataset for person detection, the quintessential vision use case for TinyML [1].
+While existing TinyML datasets [5, 29] have been valuable, they tend to be limited in scale, making it challenging to train production-grade TinyML models on these datasets.
+To enable and mature the field of TinyML research, we present the Wake Vision dataset, a large and high-quality dataset for person detection, the quintessential vision use case for TinyML [1].
 Person detection is a binary image classification task where a model detects whether a person is present.
 While simple, this task has many important use cases, such as occupancy detection [25], smart HVAC/lighting [31], or acting as an always-on ‘wake model’ in a larger ML system [11].
-Wake Vision has over 6M images, ~100x more than the prior state-of-the-art person detection dataset, Visual Wake Words (VWW) [7].
+Wake Vision has around  6M images, ~100x more than the prior state-of-the-art person detection dataset, Visual Wake Words (VWW) [5].
 Wake Vision is derived from the Open Images v7 dataset [16, 14] and is permissively licensed.
-Wake Vision demonstrates a 2.41% accuracy improvement over VWW.
-Due to the limited capacity of TinyML models, data quality is critical [26], however techniques, such as knowledge distillation [15], can leverage large, noisy datasets more effectively.
-For this reason, we choose to release two training sets: one that prioritizes dataset size, Wake Vision (Large), and one that prioritizes data quality, Wake Vision (Quality).
-We find that Wake Vision (Quality) out performs Wake Vision (Large), but we release both training sets to enable research into the tradeoffs between quantity and quality of data for TinyML.
-Person detection systems are deployed, and expected to work, in challenging settings, such as low lighting.
-However, because these challenging settings are uncommon in images sourced from the internet, high-level metrics, like test set accuracy, may overestimate a model’s real world performance.
-To address this issue we introduce a suite of five fine-grain benchmarks, which contain only images of a particular demographic or scenario.
-Our benchmarks cover distance, lighting, depictions, and perceived gender and age.
-Finally, we use our benchmark suite to conduct a case study on the performance effects of typical TinyML compression.
-This analysis demonstrates the usefulness of our benchmark and can inform future TinyML model design.
+Wake Vision demonstrates a 2.21% accuracy improvement over VWW.
+We leverage the Wake Vision dataset for an insightful study on the trade-offs between data quality and quantity in TinyML.
+We connect the dataset’s unique characteristics with the limitations of TinyML models.
+Prior work suggests that large, overparameterized models, which have more learnable parameters than data samples, can identify and adapt to errors in their training sets [32, 9].
+However, TinyML models are rarely overparameterized, leading us to hypothesize that the trade-off between data quality and quantity may be skewed towards quality for these much smaller ML models.
+To explore this hypothesis and showcase the value of the Wake Vision dataset in the TinyML domain, we release two training sets as part of the Wake Vision visual corpus: Wake Vision (Large), which prioritizes dataset size, and Wake Vision (Quality), which prioritizes data quality.
+Using a naive training script, we find that Wake Vision (Quality) outperforms Wake Vision (Large).
+However, our best-performing Wake Vision model is created by using the two sets in unison, employing Wake Vision (Large) for pre-training and Wake Vision (Quality) for fine-tuning.
+This approach demonstrates the potential of the Wake Vision dataset to advance TinyML research by leveraging both data quantity and quality.
+In addition to presenting these findings, we conduct several experiments designed to investigate the trade-offs between dataset quality and quantity across a range of models.
+To facilitate further research into these trade-offs for TinyML, we release both training sets.
+TinyML person detection systems are deployed and expected to function effectively in challenging settings, such as low lighting conditions.
+However, these challenging settings are often underrepresented in images sourced from the internet, which can lead to high-level metrics, like overall test set accuracy, overestimating a model’s real-world performance.
+To address this issue, we introduce a suite of five fine-grained benchmarks, each containing images of a specific demographic or scenario.
+The benchmarks are designed to provide a more comprehensive evaluation of a person detection model’s performance across a range of real-world conditions.
+Our benchmarks cover several factors, including distance, which evaluates the model’s performance in detecting people at various distances from the camera; lighting, which assesses the model’s accuracy in detecting individuals in both well-lit and poorly-lit environments; depictions, which measures the model’s ability to handle different representations of people, such as drawings or sculptures; perceived gender, which investigates potential biases in the model’s performance when detecting people of different perceived genders; and age, which examines the model’s effectiveness in detecting individuals across various age groups.
+By evaluating models using these fine-grained benchmarks contained within Wake Vision, we can identify potential biases, limitations, and areas for improvement that may not be apparent when relying solely on high-level metrics.
+This approach ensures that the person detection systems we develop are not only accurate on standard test sets but also robust and reliable across a diverse range of real-world settings and demographics.
 In summary, we present the following novel contributions:
-  • A new person detection dataset, called Wake Vision, which is an easily accessible, permissively licensed dataset that is ~100x larger than the prior standard.
-  • A suite of five fine-grain benchmark sets that evaluate a model’s fairness and robustness in challenging settings.
-  • A case study on the effect of TinyML scaling the highlights the importance of fine grain analysis.
+  • A new person detection dataset, called Wake Vision, which is an easily accessible, permissively licensed dataset that is two orders of magnitude larger than the prior standard.
+  • An improved, high-quality test set that reduces the estimated label error rate to only 2.2% from the prior standard of 7.8%
+  • A suite of five fine-grained benchmark sets that evaluate a model’s fairness and robustness in challenging settings.
 2 Related Work
 The current de facto standard dataset for person detection is the VWW dataset [5, 6, 18, 2, 1].
-To our knowledge, this is the only current open-source dataset explicitly meant for person detection.
+To our knowledge, this is the only current open-source dataset designed explicitly for person detection.
 Unfortunately, the VWW dataset is small and cannot be directly downloaded.
@@
 However, using general image classification datasets can lead to poor perceived performance of TinyML models, as they do not adequately represent the open “no-person” class [5].
-Tab.
-1 shows that our Wake Vision dataset contains up to almost two orders of magnitude more images than any other public and permissively licensed dataset.
+Table 1 shows that our Wake Vision dataset contains up to almost two orders of magnitude more images than any other public and permissively licensed dataset.
 Furthermore, Wake Vision is the only person detection dataset to provide a fine-grained benchmark suite and be officially distributed to popular dataset services such as TensorFlow Datasets (TFDS) and Hugging Face Datasets.
-Furthermore, Wake Vision integrates fine-grained benchmarks, which are a novel addition to existing vision data sets for TinyML.
-These fine-grained benchmark sets, which include distance, lighting, depictions, and perceived gender and age variations, allow for a deeper investigation into the effects of input and model compression on specific demographic or scenario subsets.
-Several research papers report little performance degradation from model compression on top-line metrics such as top-1 test set accuracy [6, 13, 10, 18].
-However, recent research suggests that this focus on top-line metrics can hide poor performance on critical subsets of a dataset, e.g., subsets relating to sensitive information and fairness [16, 17].
+However, recent research suggests that this focus on top-line metrics can hide poor performance on critical subsets of a dataset, such as subsets relating to sensitive information and fairness [12, 13].
 Table 1: A comparison of person detection and image classification datasets.
-Total # of Person Images Fine-Grain Suitable for
+Total # of Person Images Fine-Grained Suitable for
 Dataset Images Train Validation Test Filtering TinyML
-Wake Vision (Quality) 1,423,690 675,411 9,106 27,328
-Wake Vision (Large) 6,477,906 3,238,953 - -
+Wake Vision (Quality) 1,322,574 624,115 9,2091 27,881
+Wake Vision (Large) 5,760,428 2,880,214 - -
 Visual Wake Words [5] 123,287 36,000 3,926 19,107
@@
 Wake Vision is two orders of magnitude larger than prior comparable datasets and can be adapted based on the target use case.
+We also demonstrate how its quality is significantly improved compared to existing datsets.
 Fig.
 1 illustrates examples of images in the dataset.
-The dataset is permissively licensed for research and commercial use.
-Images are sourced from Flickr [12] images listed as having a CC-BY 2.0 license Labels are derived from Open Images and licensed under the CC-BY 4.0 license.
-In this section, we discuss our methodologies for label generation, data filtering, and error correction, as well as discussing the usability of the dataset and directly comparing it to the Visual Wake Dataset.
-Older Person
-Near Person
-Bright Image
-Female Person
-Depicted Person
-Young Person
+We discuss our methodologies for label generation, data filtering, and error correction, as well as discussing the usability of the dataset and comparing it to the Visual Wake Words Dataset.
+(a) Older Person
+(b) Near Person
+(c) Bright Image
+(d) Female Person
+(e) Depicted Person
+(f) Young Person
 Figure 1: Sample images from the Wake Vision fine-grained benchmark suite.
 3.1 Label Generation
-While a sufficiently large person detection dataset is indispensable for TinyML research, it can be challenging to obtain.
-Manually labelling millions of images would be prohibitively expensive for a nascent research field like TinyML.
-Therefore, we leverage existing large-scale data efforts to to generate Wake Vision automatically.
+While a large person detection dataset is indispensable for TinyML research, it can be challenging to obtain.
+Manually labelling millions of images would be prohibitively expensive for a nascent field like TinyML.
+Therefore, we use existing large-scale data efforts to generate Wake Vision automatically.
 The base label in Wake Vision is a binary person/non-person label.
@@
 Image-level labels describe which objects are present in an image but not localized to a specific point in the image.
-An image can contain several image-level labels, where as bounding box labels provide a set of four coordinates forming a box around the object.
-The bounding box label classes of Open Images are hierarchically structured so that one class can be a subcategory or a part of another class, e.g., a “Woman” is a subcategory of a “Person,” and a “Human Hand” is a part of “Person.”
+Bounding box labels provide a set of four coordinates forming a box around the object.
+The bounding box label classes are hierarchically structured so that one class can be a subcategory or a part of another class.
+For example a “Woman” is a subcategory of a “Person,” and a “Human Hand” is a part of “Person.”
 We observe that the bounding box labels are generally more accurate and provide more information for data filtering but are less numerous.
-The Open Images training set has over 3 million images with image-level person labels, but only ~ 700,000 images with a person bounding box.
+The Open Images training set has ~9 million images with image-level person labels, but only ~1.7 million images with a person bounding box.
 The Open Images validation and test sets are fully labeled with image-level and bounding box labels.
-Consequently, this presents a trade-off between more data (image-level labels) or higher-quality labels (bounding box labels), as different use cases may have different requirements for the training data (e.g., active learning).
-We provide two Wake Vision training sets: Wake Vision (Large) and Wake Vision (Quality).
-One larger set, labeled via Open Images image-level labels, and one smaller set, labeled via Open Images bounding box labels (Tab.
-1).
-In our experiments, the Wake Vision (Quality) training set produce more accurate models than the Wake Vision (Large) training set, strongly suggesting that label quality is more important than quantity in our setting.
-3.1.1 Large Scale Training Set from Image-Level Labels:
+Consequently, this presents a trade-off between more data (image-level labels) or higher-quality labels (bounding box labels).
+Because of this trade-off, we provide two Wake Vision training sets: Wake Vision (Large) and Wake Vision (Quality).
+One larger set, labeled via Open Images image-level labels, and one smaller set, labeled via Open Images bounding box labels (Table 1).
+Large Training Set from Image-Level Labels
 Image-level labels in Open Images contain little additional information in addition to the labels.
 The only additionally relevant information is a confidence property, which represents how certain it is that a label is correct.
-The confidence ranges from 0-1 for labels generated by machines, and is either 0 or 1 for labels negatively or positively verified by humans.
+The confidence ranges from 0 to 1 for labels generated by machines, and is either 0 or 1 for labels negatively or positively verified by humans respectively.
+See section 3.2 for more information.
 We use this property to exclude images from Wake Vision if they contain only low confidence person labels.
-3.1.2 High Quality Training Set From Bounding Box Labels:
+Quality Training Set From Bounding Box Labels
 Bounding Box labels in Open Images, in contrast to image-level labels, are all verified and localized by humans.
@@
 We allow users to adapt the Wake Vision dataset to meet the needs of specific usecases by adding several configurable options in the open-source dataset creation script.
-A comprehensive list of Open Images labels related to Wake Vision can be found in Appendix 0.I.
+For example a configuration option exists to change whether depictions are labelled as non-persons or excluded from the dataset.
+A comprehensive list of Open Images labels related to Wake Vision can be found in Sec.
+L.
 (a) An example of a person far away
@@
 Figure 2: Examples of challenging outlier images.
-3.1.3 Training Set Comparison
-To compare the performance of the two training dataset sets, we train a MobileNetV2-0.25 [31] model on each dataset for an equal number of steps and compare its performance on the common test set (Tab.
-2).
-Given the amount of label errors in the Wake Vision test set shown in Tab.
-4, a perfect person detection model would achieve an accuracy of ~93%.
-The filtered training set generated from the bounding box labels outperforms the larger training set generated from the image-level labels by 2.65% test accuracy.
+Training Set Comparison
+To compare the performance of the two Wake Vision training sets, we train a MobileNetV2-0.25 [26] model on each dataset for an equal number of steps and compare its performance on the common test set (Table 2).
+The filtered training set generated from the bounding box labels outperforms the larger training set generated from the image-level labels by 4.09% test accuracy.
 This indicates that label quality is more important than quantity in this setting.
-Interestingly, although the gap between the two is reduced to just 0.36% when training exclusively on soft labels from a teacher model, the bonding box training set is still superior.
-This indicates that the difference in quality extends beyond just the labels and that our data filtering leads to an improvement on it’s own.
-Table 2: Accuracy on Wake Vision Test set of a MobileNetV2-0.25 model trained on image-level and bounding box labels, with and without distillation.
-A MobileNetV2-1.0 model trained on the bonding box set is used as the teacher in distillation.
-Training set Label Source Dataset Size Accuracy Distilled Accuracy
-Wake Vision (Large) Image-Level 6,477,906 86.00% 88.96%
-Wake Vision (Quality) Bounding Box 1,350,822 88.65% 89.32%
-Despite the worse performance in this setting, the larger image-level training set may be useful for training higher capacity models for longer or when doing active learning [30] or distillation [5].
+The gap between the two is reduced to just 0.17% when training exclusively on soft labels from a teacher model (MobileNetV2-1.0) trained on bounding box labels.
+Combining the two training sets to train a model where Wake Vision (Large) acts as a pre-training set and Wake Vision (Quality) acts as a fine-tuning set achieves our best performance at a mean top-level test accuracy of 85.72%.
+As the images in Wake Vision (Quality) is a subset of the images in Wake Vision (Large), we do not train on more images than for Wake Vision (Large).
+Table 2: Accuracy on test set trained on image-level and bounding box labels, with and without distillation. A MobileNetV2-1.0 model trained on the bounding box set is used as the teacher in KD.
+Training set Label Source Training Set Size Accuracy Distilled Accuracy
+Wake Vision (Large) Image-Level 5,760,428 80.80$\pm$0.1885.00$\pm$0.20%
+Wake Vision (Quality) Bounding Box 1,248,230 84.89$\pm$0.1185.17$\pm$0.18%
+Wake Vision (Combined) Both 5,760,428 85.72$\pm$0.04-
 As our distillation results indicate, further data-centric filtering techniques may be needed to fully leverage the image-level training set.
@@
 Wake Vision can enable further research on data-centric TinyML [21, 22], specifically uncovering the relationship between model capacity and the sensitivity to training data quality.
-3.2 Error Correction
-Labeling errors are a challenge in computer vision [28, 4] that limit progress and obscures true performance.
+3.2 Label Correction
+Label errors are a challenge in computer vision [24, 4] that limit progress and obscures true performance.
 This is especially prevalent in large datasets that use complex labeling pipelines to scale while managing costs.
@@
 Therefore, additional measures are necessary to identify and correct such labeling omissions to ensure the dataset’s integrity.
-Given the substantial scale of Wake Vision, manually correcting label errors across the entire dataset becomes an arduous undertaking.
-To address this challenge, we employ the Confident Learning technique [27] to intelligently identify potential label errors.
-We selected Confident Learning as prior work has demonstrated its capability to find label errors in large datasets [28].
-Once these suspected mislabeled instances are flagged, we methodically inspect and correct them through a manual verification process.
-Table 3: Amount of label errors identified and corrected using Confident Learning.
-Dataset Total Suggested Corrected Suggestion
-Split Size Errors Errors Accuracy
-Validation 16,964 632 81 12.82%
-Test 51,282 1672 267 15.97%
-As shown in Tab.
-3, the confident learning process identified a large amount of possible label errors in Wake Vision’s Validation and Test sets.
-However, only between 12 and 16% of these possible label errors were legitimate errors.
-We correct these identified label errors in the Wake Vision validation and test sets.
-Given this low acceptance rate of label issues identified by Confident Learning, we conclude that we cannot automate label cleaning to a point where no human in the loop is needed.
-Based on that we also conclude that this strategy too human-intensive to be applied to the much larger training sets.
-We report our estimated label error rate after label corrections versus VWW in Tab.
-4.
-We find that after our corrections, the VWW dataset has a top-level error rate of 7.8%, which is comparable to Wake Vision’s top-level error rate of 6.8%.
-Notably, about a third of the Wake Vision label errors stem from depictions missing the depiction flag.
-As we by default do not consider depictions as persons, missing this flag results in a label error.
-Not considering the depiction label errors, Wake Vision has closer to half the number of errors of VWW.
-VWW does not, to the authors knowledge, include depictions.
-Thus no depiction errors are present for the VWW dataset.
-In real world settings, depictions are everywhere, therefore we expect that Wake Vision will prove to be a more realistic dataset.
-ML systems based on Wake Vision will therefore likely be able to handle real world scenarios better than equivalent systems based on VWW.
-We are in the process of crowd-sourcing label corrections for the validation and test sets to eliminate the remaining errors.
-However, despite these efforts, Wake Vision, in its current state, already offers a significant improvement over the VWW dataset, as detailed in Section 3.4.
-Although further enhancements to label quality are underway, Wake Vision presents a valuable and directly improved resource compared to its predecessor VWW dataset.
-Table 4: Percentage (and count) of label errors in a 500 sample subset of the VWW and Wake Vision test sets. Percent of label errors caused by misclassified depictions are reported.
-Dataset Label Error Rate % Errors From Depictions
-Visual Wake Words 7.8% (39) 0% (0)
-Wake Vision 6.8% (34) 38.2% (13)
+Given Wake Vision’s size, manually correcting label errors across the entire dataset becomes an arduous undertaking.
+Therefore, we prioritize labeling the Wake Vision validation and test sets so that the dataset accurately represents the model’s performance.
+To do so we make use of the Scale AI platform to crowd-source manual label corrections (see Appendix).
+With the ground truth established in the validation and test sets, Wake Vision becomes an ideal foundation for research on automated data cleaning techniques [23].
+We perform an initial exploration of these methods in appendix N.
+We report our estimated label error rate after label corrections versus VWW in Table 3.
+The Wake Vision validation and test set is considerably lower than that of the VWW dataset top-level error rate.
+Table 3: Estimated label error rate of VWW and Wake Vision, based on a 500 sample subset.
+The Wake Vision estimated error rate is broken down by set.
+The Wake Vision Train (Quality) and Train (Large) have different error rates due to the different sources of their labels.
+The Wake Vision Validation and Test sets have a lower error rate due to manual relabeling.
+Dataset Wake Vision Set Label Error Rate
+Visual Wake Words Train, Val, & Test 7.8%
+Wake Vision Train (Large) 15.2%
+Wake Vision Train (Quality) 6.8%
+Wake Vision Val & Test 2.2%
 3.3 Usability
-Wake Vision is available through TensorFlow Datasets [33] and HuggingFace Datasets [22] to be easily accessed and used by the community.
-The images and labels are rehosted to ensure the dataset will not be changed over time due to dead links, making it a more viable benchmark.
-The rehosted labels are generated according to our default dataset configuration, further described in Appendix 0.F.
+Wake Vision is available through TensorFlow Datasets [28] and HuggingFace Datasets [17] to enable easy access and use by the community.
+The images and labels are rehosted to ensure the dataset will not be changed over time due to dead links, making it a more stable benchmark.
+The rehosted labels are generated according to our default dataset configuration, further described in Sec.
+I.
 3.4 Comparison to Visual Wake Words
 Wake Vision serves as a drop-in improvement over VWW, thanks to its size and our extensive filtering.
-To compare the benefits of Wake Vision against VWW, we train two identical models using the same recipe for the same number of steps, one utilizing VWW’s training set and the other employing Wake Vision’s training set.
-For this comparative evaluation, both trained models are MobileNetV2 architectures with a width modifier of 0.25, which have previously served as benchmark TinyML models [7].
+To compare the benefits of Wake Vision against VWW, we train two identical MobileNetv2 [26] models with a width modifier of 0.25 for 200,000 steps on 224x224x3 images using AdamW [20].
+We use a lr of 0.002 with a cosine decay and a weight decay of 4E-6.
+One model uses VWW’s training set, and the other employs Wake Vision’s Quality training set.
 After training these identical models using the respective datasets’ training sets with the same recipe for an equal number of steps, we cross-evaluate their performance on the corresponding test sets.
-In Tab.
-5 we demonstrate a direct 0.5% improvement on VWW’s own test set, which illustrates that Wake Vision is a direct improvement over VWW, not simply a domain shift.
-We additionally achieve a 2.41% improvement over VWW on Wake Vision’s Test set, indicating that the new test set is more challenging.
-Table 5: Accuracy on the Wake Vision and VWW test sets by models trained on each dataset’s respective training set.
+Table 4 shows a 0.26% improvement on VWW’s own test set, which illustrates that Wake Vision is a direct improvement over VWW, not simply a domain shift.
+We additionally achieve a 1.1% improvement over VWW on Wake Vision’s Test set, indicating that the new test set is more challenging.
+Table 3 also shows the performance of a the Wake Vision (Combined) model described in section 3.1.
+This model achieves an average 0.75% improvement over the Wake Vision (Quality) trained model on the VWW test set and an average 0.83% improvement on the Wake Vision test set.
+The Wake Vision (Combined) model is trained for longer than the VWW and Wake Vision (Quality) models, but shows the value of using the training sets in unison even on the VWW test set.
+Table 4: Accuracy on the Wake Vision and VWW test sets by models trained on the VWW, Wake Vision (Quality) and Wake Vision (Combined) training sets.
 Train
-VWW Wake Vision
-Test VWW 88.00% 88.50%
-Wake Vision 86.27% 88.68%
+VWW Wake Vision (Quality) Wake Vision (Combined)
+Test VWW 88.33$\pm$0.29 88.59$\pm$0.1789.34$\pm$0.02%
+Wake Vision 83.79$\pm$0.2384.89$\pm$0.1185.72$\pm$0.04%
+3.5 Dataset Quality Vs. Size
+We investigate the impact of dataset quality and size of Resnet [10] style models of varying capacity.
+We measure dataset quality by the approximate rate of label errors.
+The Wake Vision (Quality) training set has an estimated error rate of around 7%.
+We simulate higher error rate datasets (15% and 30%) by flipping the binary label with a certain probability.
+Appendix D gives the derivation for the flip probability.
+We also sweep the dataset size by taking a slice of Wake Vision (Quality).
+Each model is trained for 50,000 steps on 224x224x3 images using AdamW [20], a learning rate of 0.001 with a cosine schedule and a weight decay of 0.004.
+Figure 3 illustrates the experiment’s results.
+We observe that smaller models (leftmost figure) are more sensitive to higher error rates than large models.
+The smallest model’s accuracy drop is 1.3%, going from the base train (quality) error rate of around 7% to 15% compared to only a 0.5% accuracy loss for the largest model.
+In contrast, large models benefit more from big datasets and start to over fit on the smaller slices of the training data.
+Figure 3: Comparing the impact of dataset error rate and size on ResNet models of varying capacity.
+The $y$-axis is the Wake Vision test accuracy and the $x$-axis is the percentage of Wake Vision Train (Quality) used in training.
+The error % indicates the expected rate of incorrectly labeled samples.
 4 Fine Grain Benchmark Suite
@@
 In contrast, person detection systems are deployed, and expected to work, in more challenging settings, such as low lighting.
-Due to this gap, a model that achieves high test accuracy may end up performing poorly in real world scenarios, after the design stage has ended.
-To address this issue, we present a benchmark suite for TinyML person detection that tests a model’s robustness in challenging settings, enabling better analysis at the design stage.
-Additionally the benchmark suite measures the model’s performance across demographics, to ensure the model does not exhibit clear bias.
-The suite consists of five fine-grain benchmark sets, which we elaborate on below.
+Due to this gap, a model that achieves high test accuracy may perform poorly in real-world scenarios after the design stage has ended.
+To address this issue, we present a benchmark suite for person detection that tests a model’s robustness in challenging settings, enabling better analysis at the design stage.
+Additionally, the suite measures the model’s performance across demographics to ensure it does not exhibit clear bias.
+The suite consists of five fine-grained benchmark sets.
 Each of the five fine-grained benchmarks have been chosen based on a combination of its relevance to TinyML use cases and the availability of requisite metadata to generate the sets.
-Each benchmark set is a subset of the validation or test set filtered based on the criteria under test (e.g.
-persons far away from the camera).
-The typical usage of the benchmark is to determine if a model is sufficiently accurate in the planned deployment setting.
-For example, a model designer may make different design choices for a use case where the subject is close to the camera and well lit, vs.
-the inverse setting.
-In this section, we describe the design of each individual benchmark and use the suite to compare a VWW model to a Wake Vision model.
-Tab.
-6 reports the number of samples in each of the individual benchmark sets and the F1 scores of a Wake Vision and VWW model on each set.
-4.1 Perceived Gender and Age
-The Perceived Gender and Age fine grained datasets are generated from the Open Images More Inclusive Annotations for People (MIAP) extended fairness labels [32].
+Each benchmark set is a subset of the respective validation or test set filtered based on the criteria under test.
+The benchmark determines whether a model is sufficiently accurate in the planned deployment setting.
+For example, a model designer may make different design choices for a use case where the subject is close to the camera and well-lit compared to the inverse setting.
+Appendix C contains a case study showing how the fine grained evaluation sets can be used to design robust models.
+Table 5 reports the number of samples in each of the individual benchmark sets and the F1 scores of a Wake Vision and VWW model on each set.
+Perceived Gender and Age
+These fine grained datasets are generated from the Open Images More Inclusive Annotations for People (MIAP) extended fairness labels [27].
 Underrepresented sub-groups typically constitute a small portion of a generic test set; therefore, top-line metrics often obscure a bias in a model until it is deployed [12].
 This benchmark aims to evaluate a model separately on demographics that are underrepresented in the underlying dataset distribution, identifying bias.
-These labels are based on a perceived gender and age representation, and are not necessarily representative of a persons gender identity or age.
-Like the authors of the original MIAP labels, we do not support or condone the creation or deployment of gender or age specific classifiers based on this dataset.
-Table 6: Wake Vision Fine Grain Benchmark Suite. We report the samples in each set as well as the F1 score of a Wake Vision model and a VWW model on the test set.
+These labels are based on a perceived gender and age representation and are not necessarily representative.
+Table 5: Wake Vision Fine Grain Benchmark Suite. We report the samples in each set and the average F1 score across three Wake Vision models and three VWW models on the benchmark set.
 Set Size F1-Score
 Benchmark Val Test Wake Vision VWW
-Gender Female 675 2,152 0.94 0.89
-Male 1,238 3,748 0.93 0.89
-Unknown 1,718 5,231 0.78 0.79
-Age Young 271 875 0.94 0.90
-Middle 2,035 6,371 0.93 0.89
-Older 91 271 0.96 0.90
-Unknown 1,388 4,120 0.73 0.76
-Distance Near 5,888 17,536 0.90 0.88
-Medium 2,304 7,168 0.83 0.83
-Far 1,152 3,328 0.39 0.49
-Lighting Dark 3,036 8,872 0.88 0.83
-Normal 12,992 39,327 0.88 0.86
-Bright 936 3,083 0.85 0.85
-Depictions Person 179 501 0.83 0.77
-Non-Person 313 957 0.89 0.85
-No Depiction 7,990 24,183 0.90 0.87
-4.2 Distance
-The distance fine grained datasets aim to test how the distance of persons in images impact the performance of ML systems.
+Gender Female 684 2,157 0.93 0.89
+Male 1,310 3,918 0.91 0.88
+Unknown 1,612 4,940 0.77 0.78
+Age Young 275 884 0.94 0.90
+Middle 2,133 6,595 0.91 0.88
+Older 90 276 0.94 0.89
+Unknown 1,299 3,837 0.71 0.74
+Distance Near 5,457 16,333 0.91 0.89
+Medium 2,213 6,876 0.85 0.84
+Far 398 1,140 0.59 0.67
+Lighting Dark 3,255 9,420 0.85 0.81
+Normal 14,315 43,010 0.82 0.82
+Bright 1,012 3,332 0.82 0.82
+Depictions Person 356 978 0.71 0.66
+Non-Person 352 1101 0.86 0.82
+No Depiction 8,583 25,802 0.87 0.85
+Distance
+The distance datasets test how the distance of persons in images impact the performance of ML systems.
 This benchmark is critical for correctly predicting a model’s deployed performance.
@@
 The three sets are near (>60%), at a medium distance (10-60%), and far away (<10%).
-4.3 Lighting
-The lighting fine grained datasets aim to test the performance of ML systems across different lighting conditions.
-For instance, in scenarios such as surveillance cameras, outdoor robotics, or augmented reality applications, models need to be robust to varying lighting conditions, including low-light environments.
-As with distance, this benchmark provides insight into possible causes of significant performance degradation once deployed, such as low-light settings.
-We create three fine grained datasets of this type for dark, normal and bright lighting conditions respectively.
-All images in the original Wake Vision Vision validation and test set are classified into these datasets according to their lighting condition.
+Lighting
+The lighting datasets aim to test the performance of ML systems across different lighting conditions.
+In scenarios like security cameras, outdoor robotics, or augmented reality applications, models must be robust to varying lighting conditions, including low-light environments.
+We create three fine grained datasets of this type for dark, normal, and bright lighting conditions, respectively.
 We quantify lighting conditions by the average pixel values of images in greyscale, which is a simple but effective method for distinguishing lighting conditions [33].
 We define low as an average pixel value less than 85, normal as between 85 and 170, and bright lighting conditions as greater than 170.
-4.4 Depictions
+Depictions
 A particularly challenging task for a person detection model is to correctly reject depictions of people.
@@
 4.1 Benchmark Results
-Tab.
-6 compares the results of a model trained on the Wake Vision (Quality) training set to an identical model trained on VWW.
+Table 5 compares the results of a model trained on the Wake Vision (Quality) training set to an identical model trained on VWW.
 The Wake Vision model exhibits superior robustness across the challenging settings exercised by our benchmarking suite, further demonstrating the effectiveness of the dataset.
-For instance, on the “Depictions” benchmark, which evaluates performance on images containing persons, non-person objects, or no depictions, the Wake Vision model achieves an F1 score of 0.83 for person depictions, outperforming the VWW model’s 0.77.
-Similarly, for non-person depictions, the Wake Vision model scores 0.89 compared to the VWW model’s 0.85.
-The Wake Vision model also showcases improved performance on the “Age” benchmark, with F1 scores of 0.94 for young subjects, 0.93 for middle-aged, and a notable 0.96 for older individuals, surpassing the VWW model’s scores of 0.90, 0.89, and 0.90, respectively.
-This highlights the Wake Vision dataset’s effectiveness in enhancing model robustness for detecting people across various age groups, particularly the elderly demographic.
-In addition to age and depictions, the Wake Vision model demonstrates resilience to challenging lighting conditions, achieving F1 scores of 0.88 for both dark and normal lighting scenarios, compared to the VWW model’s 0.83 and 0.86, respectively.
+For instance, on the “Depictions” benchmark, which evaluates performance on images containing persons, non-person objects, or no depictions, the Wake Vision model achieves an F1 score of 0.71 for person depictions, outperforming the VWW model’s 0.66.
+Similarly, the Wake Vision model scores 0.86 for non-person depictions compared to the VWW model’s 0.82.
+The Wake Vision model also showcases improved performance on the “Age” benchmark, with F1 scores of 0.94, 0.91, and 0.94 for young, middle, and older individuals, surpassing the VWW model’s scores of 0.90, 0.88, and 0.89, respectively.
+This highlights the dataset’s effectiveness in enhancing model robustness for detecting people across various age groups, particularly the elderly demographic.
+In addition, the Wake Vision model demonstrates resilience to challenging lighting conditions, achieving F1 scores of 0.85 in dark lighting scenarios, compared to the VWW’s 0.81.
 Collectively, these results underscore the significance of Wake Vision’s large-scale, diverse data and fine-grained benchmarks, enabling the development of more robust and reliable TinyML person detection models across a wide range of realistic and real-world scenarios.
+5 Ethical Considerations
+The Wake Vision dataset and benchmark suite aim to advance TinyML research by providing large-scale datasets and enabling fine-grained analysis of TinyML systems.
+Wake Vision can directly improve the quality of person detection systems that benefit society, such as energy-saving applications, while protecting user privacy through on-device inference.
+However, person detection models can also be used for malicious purposes, such as weapons targeting or large-scale surveillance.
+The images in Wake Vision are sourced from Flickr through Open Images[16, 14] under the CC-BY 2.0 license.
+While we have made efforts to ensure the images are properly licensed, we cannot guarantee the license status of each image.
+It is possible that some images have been uploaded without the right to distribute under the CC-BY 2.0 license.
+Given sufficient resources, individuals in the dataset could potentially be identified based on biometric data.
+The dataset may also contain offensive content, as manually verifying millions of images is infeasible.
+The benchmark suite aims to ensure the fairness and robustness of person detection models.
+However, the demographic benchmarks could potentially be misused to create systems that classify gender and age, breaching personal privacy or perpetuating discrimination.
+To mitigate this risk, we have not released training sets for the fine-grained benchmark datasets.
+6 Conclusions
+The lack of large, high-quality datasets has been a significant bottleneck in the field of TinyML research.
+To address this issue, we introduced Wake Vision, a large-scale dataset and benchmark suite for person detection that surpasses the current state-of-the-art dataset, VWW, in terms of size (~100x larger), ease of use, label accuracy, and real-world relevance.
+These improvements make Wake Vision a drop-in replacement for VWW.
+Recognizing the importance of fairness across demographics and robustness to challenging settings in person detection use cases, we also released a Wake Vision benchmark suite comprising six fine-grained benchmark sets to evaluate the fairness and robustness of person detection models.
+Furthermore, our investigation into the trade-off between dataset quality and quantity for TinyML systems suggests that dataset quality is likely more crucial for TinyML systems compared to conventional ML systems.
+The Wake Vision dataset and benchmark suite aim to facilitate the development of more accurate, fair, and robust person detection models in TinyML.
+7 Acknowledgements
+This work was partially supported by the Google TPU Research Cloud program. We would like to thank Andrew Howard of Google Deepmind for his input and support in this effort.
+Appendix A Code Repository
+The code used to generate the Wake Vision dataset is available at the following GitHub repository: https://github.com/harvard-edge/Wake_Vision
+This repo contains the code to generate Wake Vision and the benchmark suite, as well as the code to train and evaluate models. This code is sufficient to reproduce all results in the paper.
+Appendix B Flowchart of Bounding Box Filtering Process
+fig. 4 illustrates the filtering process for Wake Vision when using Open Images’ bounding box labels as the label source.
+Figure 4: A flowchart of the filtering of an image according to default Bounding Box filtering rules.
 Figure 5: Effects of scaling the image size vs.
@@
 When not specified the image size is 160 and the width multiplier is 1.0.
-5 Model Design Case Study
+Appendix C Model Design Case Study
 Basic test set performance can often misrepresent a model’s performance, given the typical domain shift between images scraped from the internet and real-world use cases.
@@
 While these results can inform ML practitioners, our intention is to demonstrate the usefulness of the benchmark suite.
-5.1 Image Size vs. Model Width Scaling
+C.1 Image Size vs. Model Width Scaling
 We train two series of MobileNetV2 models: one series that sweeps the input image size [64-256] and one that sweeps the width multiplier of a model [0.1-1.5].
@@
 We plot these results against the number of multiply accumulate (MAC) operations in the model as a proxy for on-device latency [2].
-The results in Fig.
-3 (left) show that when looking exclusively at the high-level metric (i.e., test accuracy), scaling the input image size has a similar impact as scaling the model size.
-However, as shown in Fig.
-3 (right), when we consider only samples where the person is far away from the camera (i.e., the distance benchmark), we observe a much more significant impact when scaling the image size.
+The results in Table 5 (left) show that when looking exclusively at the high-level metric (i.e., test accuracy), scaling the input image size has a similar impact as scaling the model size.
+However, as shown in Figure 5 (right), when we consider only samples where the person is far away from the camera (i.e., the distance benchmark), we observe a much more significant impact when scaling the image size.
 In the case of distant subjects, the image size becomes the bottleneck.
@@
 The distance benchmark in Wake Vision effectively unveils the disproportionate impact of image size on model performance for distant subjects, enabling more informed decision-making during model optimization.
-5.2 Quantization
+C.2 Quantization
 Quantization is a crucial technique for deploying efficient TinyML models, offering substantial benefits in terms of reduced latency, memory footprint, and model size.
@@
 To assess the implications of quantization in the context of person detection, we investigate the impacts of int8 quantization on a model’s benchmark results across Wake Vision’s fine-grained benchmarks.
-[…]
-5.3 Grayscale
+Our findings show negligible degradation in performance across all benchmarks (±$\pm 0.004$ F1) when employing int8 quantization, even on outlier sets.
+This result contradicts the previously observed disproportionate impact of quantization on underrepresented subsets.
+We speculate that person detection may be a relatively simple task, potentially explaining why we do not observe this specific property of quantization in our experiments.
+Given the negligible performance degradation and the substantial latency, memory, and model size benefits of quantization, we conclude that quantization is a win for person detection.
+C.3 Grayscale
 Converting a model’s input image channels to grayscale from RGB is a commonly employed optimization in the TinyML field [2] as it can substantially reduce a model’s memory consumption.
-We observed, however, that the grayscale optimization disproportionately impacts images on the brighter end of the spectrum as illustrated in Fig.
-4.
+We observed, however, that the grayscale optimization disproportionately impacts images on the brighter end of the spectrum as illustrated in Figure 6.
 This further demonstrates the importance of fine grained analysis, as some real world deployment environments might be far brighter than the average Wake Vision test sample.
 Figure 6: Impact of grayscale input images on the lighting benchmarks: dark, normal light, and bright. Models that use grayscale input images are more sensitive to bright lighting conditions than RGB.
-6 Ethical Considerations
-The primary purpose of the Wake Vision dataset and benchmark suite is to mature the field of TinyML by giving researchers access to large-scale datasets and enabling fine-grain analysis.
-Beyond this purpose, Wake Vision can directly improve the quality of person detection systems that benefit society, such as saving energy by only turning on the lights in a room when a person is present.
-Additionally, TinyML person detection protects user privacy by running inference on-device and not transmitting private data over the network.
-However, it should be noted that person detection models can also be used for malicious purposes, such as weapons targeting systems or large-scale surveillance.
-The images that make up Wake Vision are entirely sourced from Flickr.
-While we attempt to only use images under the consent given by the CC-BY 2.0 license (by sourcing images through Open Images[21, 19]), we make no representations or warranties regarding the license status of each image.
-It is possible that some of the millions of images have been uploaded to Flickr by a user without the right to distribute the images with a CC-BY 2.0 license.
-Futhermore, the benchmark suite is intended to ensure the fairness and robustness of person detection models.
-While the goal of these benchmark suites is to build more robust and fair models, our demographic benchmarks could be misused to create systems to classify gender and age, potentially breaching personal privacy or perpetuating discrimination.
-We have attempted to mitigate this risk by not releasing training sets for the fine-grained benchmark datasets.
-7 Conclusions
-The field of TinyML research is currently bottlenecked by the lack of large, high-quality datasets.
-To alleviate this bottleneck we introduced Wake Vision—a large scale dataset and benchmark suite for person detection.
-Compared to VWW, the current state of the art dataset for person detection, Wake Vision is ~100x larger, easier to use, has a comparable amount of label errors, and is more reflective of real world use cases.
-These factors contribute to Wake Vision being a drop-in improvement over VWW.
-Fairness across demographics and robustness to challenging settings is essential for many person detection use cases.
-To facilitate evaluation of this, we released a Wake Vision benchmark suite of six fine-grained benchmark sets to test the fairness and robustness of person detection models.
-Lastly we demonstrate how the Wake Vision benchmark suite can be used to investigate the effects of different types of input and model scaling and highlight the importance of fine grain analysis when designing an model.
-Appendix 0.A Anonymized Repo
-The code used for the paper is available at this anonymous repository: https://anonymous.4open.science/r/Wake_Vision-0D15/README.md
-This repo contains the code to generate Wake Vision and the benchmark suite, as well as the code to train and evaluate models. This code is sufficient to reproduce all results in the paper.
-Appendix 0.B Flowchart of Bounding Box Filtering Process
-Fig. 5 illustrates the filtering process for Wake Vision when using Open Images’ bounding box labels as the label source.
-Figure 5: A flowchart of the filtering of an image according to default Bounding Box filtering rules.
-Appendix 0.C Fine Grained Benchmark Images
+Appendix D Inducing Label Errors
+The goal is to make a single pass through the dataset and flip the labels of a binary classification dataset such that the expected label error rate is $d$.
+There is an underlying rate of label errors $e$.
+If we flip one of these underlying errors, we correct it, thereby inadvertently decreasing the label error rate.
+After flipping labels with a probability of $p$ we can claim that the likelihood of a single label being correct is the probability that we flipped the label and it was originally an error plus the probability that we didn’t flip the label and it was originally correct: $1-d=p*e+(1-p)(1-e)$.
+Then solving for $p$ gives $p=(e-d)/(2e-1)$.
+A current flaw of this method is that the injected label errors are not consistent between epochs, which would likely be less destructive to a model’s accuracy since the same errors are not reinforced each epoch.
+This could also potentially explain why large models in Fig.
+3 don’t overfit on training data with higher label errors, as the inconsistent label noise has a regularizing effect.
+Appendix E Fine Grained Benchmark Images
 fig. 7 gives example images for each of the benchmarks in the suite.
@@
 Figure 7: Images from each fine grained benchmark dataset.
-Appendix 0.D Open Image Label Distribution
+Appendix F Open Image Label Distribution
 Wake Vision is derived from Open Images V7, therefore the diversity of subjects in the images should follow the distribution of labels in Open Images.
 The Label distribution of Open images can be found here: https://storage.googleapis.com/openimages/web/factsfigures_v7.html#statistics
-Appendix 0.E Dataset Access and Organization
-Wake Vision is available directly online, via TFDS, or Hugging Face Datasets. In order to maintain author anonymity, we do not link to the dataset.
+Appendix G Manual Labeling
+The validation and test sets were manually labeled through the crowd-sourced labeling platform Scale Rapid(https://scale.com/.
+Figure 8 shows a screenshot of the labeling menu.
+The labelers were instructed to label an image as a "person" if a person was present anywhere in the image and "no person" if no visible person was present.
+The labelers also indicated if the image contained a depiction of a person, which meant that sample was labeled "no person" in Wake Vision, or a "picture of a picture/reflection", which we used for metadata.
+The cost per image was \$0.10 USD and is set based on Scale’s pricing structure.
+Each image is labeled by 3 different labelers to form a consensus.
+The total cost of labeling was \$7089.8 USD.
+The authors do not know the hourly rate paid to each labeler, but Scale lists \$18/hr on job postings at the time of writing.
+For context on the difficulty of the task, the authors were able to average around 500 images per hour at a reasonable pace when estimating error rates.
+Figure 8: A screenshot of the labeling menu used to manually label the validation and test sets.
+Appendix H Dataset Access and Organization
+Wake Vision is available online at the Harvard Dataverse, via TFDS, or Hugging Face Datasets. More information can be found on the https://wakevision.ai webpage.
 The dataset is organized as a set of compressed tar files containing images, and a series of label CSVs.
 The label CSVs are organized such that the file name of the image is the identifying index.
-The person label is provided as is a flag to indicate if the image is of a depiction.
+CSVs for all dataset splits include the person and depiction label.
 The Validation and Test label CSVs also have flags that denote a sample’s inclusion into a fine-grain benchmark set (e.g.
@@
 It also ensures the dataset can be easily updated as new versions are introduced.
-Appendix 0.F Label Generation Details
+Appendix I Label Generation Details
 Person Labels.
-The most straightforward Open Images label classes to label as person in Wake Vision are the "person" label and its subcategories (listed in Appendix 0.I).
+The most straightforward Open Images label classes to label as person in Wake Vision are the "person" label and its subcategories (listed in appendix L).
 All of these are relabelled as persons in Wake Vision.
@@
 For different requirements, users can change a configuration parameter in our open-source dataset creation code.
-Appendix 0.G Standardized Evaluation
+Appendix J Standardized Evaluation
 VWW has no standardized way to evaluate performance on the dataset.
@@
 To allow for both data-centric and model-centric improvements, we provide a standard model for data-centric contributions, and a standard pre-processing pipeline for model-centric contributions.
-Both types of contributions are expected to use accuracy as the primary metric for the overall test and validation set, and F1-score as the primary metric for the fine grained benchmarks introduced in Sec.
-4.
-Therefore, for data-centric contributions we propose to use a Mobilenet v2 model with a width modifier***Also known as alpha.
+Both types of contributions are expected to use accuracy as the primary metric for the overall test and validation set, and F1-score as the primary metric for the fine grained benchmarks introduced in section 4.
+Therefore, for data-centric contributions we propose to use a Mobilenet v2 model with a width modifier [^1] Also known as alpha.
 of 0.25 [26].
@@
     Use image tensor as input features and person label as target feature
-Appendix 0.H Dataset Download
+Appendix K Open Images Download
 The full Open Images v7 dataset is not hosted by the dataset authors.
@@
 As an unfortunate result of this, the dataset is not static over time as image owners can delete their images from the flickr platform.
-As a result we were only able to download a subset of the original Open Images v7 dataset as shown in Tab.
-7.
+As a result we were only able to download a subset of the original Open Images v7 dataset as shown in table 6.
 Table 6: Number of images downloaded from Open Images v7. Download occurred between the 28th of November to the 5th of December
@@
 Errors 1,055,669 5,214 16,131
-Appendix 0.I Person Label Classes
+Appendix L Person Label Classes
 We consider the following Open Images v7 labels to be a person for the Wake Vision dataset:
@@
   • Beard (Part of Person)
-Appendix 0.J Wake Vision Dataset Size
+Appendix M Wake Vision Dataset Size
 Table 7: Amount of images in the Wake Vision dataset
 Person Images Non-Person Images Excluded
-Image Level Label Train Dataset 2,880,214 2,880,214 2,176,551
-Bounding Box Label Train Dataset 624,115 624,115 6,688,749
-Validation Dataset 8,482 8,482 19,442
-Test Dataset 25,641 25,641 58,023
+Wake Vision (Large) Training Dataset 2,880,214 2,880,214 2,176,551
+Wake Vision (Quality) Training Dataset 624,115 624,115 6,688,749
+Validation Dataset 9,291 9,291 17,824
+Test Dataset 27,881 27,881 53,543
+Appendix N Automatic Label Correction
+To address the challenge of correcting Wake Vision validation and test labels we initially employed the Confident Learning technique [23] to intelligently identify potential label errors.
+We selected Confident Learning as prior work has demonstrated its capability to find label errors in large datasets [24].
+Confident Learning flagged suspected mislabeled instances, which we inspected and corrected through a manual verification process.
+Table 8: Amount of label errors identified and corrected using Confident Learning.
+Dataset Total Suggested Corrected Suggestion
+Split Size Errors Errors Accuracy
+Validation 18,582 632 81 12.82%
+Test 51,282 1672 267 15.97%
+As shown in Table 8, the confident learning process identified a large amount of possible label errors in Wake Vision’s Validation and Test sets.
+However, only between 12 and 16% of these possible label errors were legitimate errors.
+We corrected these identified label errors in the Wake Vision validation and test sets.
+Given this low acceptance rate of label issues identified by Confident Learning, we concluded that we could not automate label cleaning to a point where no human in the loop is needed.
+This made this strategy too human-intensive to be applied to the much larger training sets.
+To further correct label errors in the final Wake Vision Validation and Test sets we employed the Scale AI platform to crowd-source manual label corrections.
+These label corrections are described in section 3.2 and deprecate the automatic label corrections described in this section.

### v3

@@
-Wake Vision: A Large-scale, High-Quality Dataset and Benchmark Suite for TinyML Applications
+Wake Vision: A Tailored Dataset and Benchmark Suite for TinyML Computer Vision Applications
 Abstract
-Tiny machine learning (TinyML), which enables machine learning applications on extremely low-power devices, suffers from limited size and quality of relevant datasets.
-To address this issue, we introduce Wake Vision, a large-scale, diverse dataset tailored for person detection—the canonical task for TinyML visual sensing.
-Wake Vision comprises over 6 million images, representing a hundredfold increase compared to the previous standard, and has undergone thorough quality filtering.
-We provide two Wake Vision training sets: Wake Vision (Large) and Wake Vision (Quality), a smaller set with higher-quality labels.
-Our results demonstrate that using the Wake Vision (Quality) training set produces more accurate models than the Wake Vision (Large) training set, strongly suggesting that label quality is more important than quantity in our setting.
-We find use for the large training set for pre-training and knowledge distillation.
-To minimize label errors that can obscure true model performance, we manually label the validation and test sets, improving the test set error rate from 7.8% in the prior standard to only 2.2%.
-In addition to the dataset, we provide a collection of five detailed benchmark sets to facilitate the evaluation of model quality in challenging real-world scenarios that are often ignored when focusing solely on overall accuracy.
-These novel fine-grained benchmarks assess model performance on specific segments of the test data, such as varying lighting conditions, distances from the camera, and demographic characteristics of subjects.
-These benchmarks enable researchers and practitioners to better understand the strengths and limitations of their models in real-world scenarios.
-Our results demonstrate that using Wake Vision for training results in a 2.49% increase in accuracy compared to the established dataset.
-We also show the importance of dataset quality for low-capacity models and the value of dataset size for high-capacity models.
-The dataset, benchmark suite, code, and models are publicly available under the CC-BY 4.0 license at https://wakevision.ai/.
-**footnotetext: These authors contributed equally to this work.11footnotetext: Work done while a visiting researcher at Harvard University.
+Tiny machine learning (TinyML) for low-power devices lacks robust datasets for development.
+We present Wake Vision, a large-scale dataset for person detection that contains over 6 million quality-filtered images.
+We provide two variants: Wake Vision (Large) and Wake Vision (Quality), leveraging the large variant for pretraining and knowledge distillation, while the higher-quality labels drive final model performance.
+The manually labeled validation and test sets reduce error rates from 7.8% to 2.2% compared to previous standards.
+In addition, we introduce five detailed benchmark sets to evaluate model performance in real-world scenarios, including varying lighting, camera distances, and demographic characteristics.
+Training with Wake Vision improves accuracy by 1.93% over existing datasets, demonstrating the importance of dataset quality for low-capacity models and dataset size for high-capacity models.
+The dataset, benchmarks, code, and models are available under the CC-BY 4.0 license, maintained by the Edge AI Foundation.
 1 Introduction
-To reduce the energy consumption and cost of deploying Machine Learning (ML) deployments, researchers have pioneered ultra-low-power machine learning (TinyML) [3, 30, 1, 6].
-TinyML co-locates ML models with the sensors by deploying the model onto always-on microcontrollers (MCUs) and TinyML accelerators, thereby reducing the energy cost of transmitting data or constantly running larger processors.
-Deploying models on these tiny devices imposes severe constraints, confining the model’s size to $\sim$100 kilobytes, nearly four orders of magnitude less memory than mobile phones [2].
+Ultra-low power machine learning (TinyML) has emerged as an important technology to enable efficient ML deployments by co-locating models with sensors on microcontrollers (MCUs) [3, 38, 1, 6].
+Although this approach dramatically reduces energy consumption and deployment costs, it imposes severe constraints.
+For instance, the models must often be within hundreds of kilobytes of memory in size, nearly four orders of magnitude less than mobile phones[2].
 TinyML requires large-scale datasets to facilitate high-quality research.
@@
 As a result, conventional ML datasets like ImageNet [7], designed for more resource-intensive models, are often ill-suited for TinyML research endeavors.
-While existing TinyML datasets [5, 29] have been valuable, they tend to be limited in scale, making it challenging to train production-grade TinyML models on these datasets.
-To enable and mature the field of TinyML research, we present the Wake Vision dataset, a large and high-quality dataset for person detection, the quintessential vision use case for TinyML [1].
-Person detection is a binary image classification task where a model detects whether a person is present.
-While simple, this task has many important use cases, such as occupancy detection [25], smart HVAC/lighting [31], or acting as an always-on ‘wake model’ in a larger ML system [11].
-Wake Vision has around  6M images, ~100x more than the prior state-of-the-art person detection dataset, Visual Wake Words (VWW) [5].
-Wake Vision is derived from the Open Images v7 dataset [16, 14] and is permissively licensed.
-Wake Vision demonstrates a 2.21% accuracy improvement over VWW.
-We leverage the Wake Vision dataset for an insightful study on the trade-offs between data quality and quantity in TinyML.
-We connect the dataset’s unique characteristics with the limitations of TinyML models.
-Prior work suggests that large, overparameterized models, which have more learnable parameters than data samples, can identify and adapt to errors in their training sets [32, 9].
-However, TinyML models are rarely overparameterized, leading us to hypothesize that the trade-off between data quality and quantity may be skewed towards quality for these much smaller ML models.
-To explore this hypothesis and showcase the value of the Wake Vision dataset in the TinyML domain, we release two training sets as part of the Wake Vision visual corpus: Wake Vision (Large), which prioritizes dataset size, and Wake Vision (Quality), which prioritizes data quality.
-Using a naive training script, we find that Wake Vision (Quality) outperforms Wake Vision (Large).
-However, our best-performing Wake Vision model is created by using the two sets in unison, employing Wake Vision (Large) for pre-training and Wake Vision (Quality) for fine-tuning.
-This approach demonstrates the potential of the Wake Vision dataset to advance TinyML research by leveraging both data quantity and quality.
-In addition to presenting these findings, we conduct several experiments designed to investigate the trade-offs between dataset quality and quantity across a range of models.
-To facilitate further research into these trade-offs for TinyML, we release both training sets.
-TinyML person detection systems are deployed and expected to function effectively in challenging settings, such as low lighting conditions.
-However, these challenging settings are often underrepresented in images sourced from the internet, which can lead to high-level metrics, like overall test set accuracy, overestimating a model’s real-world performance.
-To address this issue, we introduce a suite of five fine-grained benchmarks, each containing images of a specific demographic or scenario.
-The benchmarks are designed to provide a more comprehensive evaluation of a person detection model’s performance across a range of real-world conditions.
-Our benchmarks cover several factors, including distance, which evaluates the model’s performance in detecting people at various distances from the camera; lighting, which assesses the model’s accuracy in detecting individuals in both well-lit and poorly-lit environments; depictions, which measures the model’s ability to handle different representations of people, such as drawings or sculptures; perceived gender, which investigates potential biases in the model’s performance when detecting people of different perceived genders; and age, which examines the model’s effectiveness in detecting individuals across various age groups.
-By evaluating models using these fine-grained benchmarks contained within Wake Vision, we can identify potential biases, limitations, and areas for improvement that may not be apparent when relying solely on high-level metrics.
-This approach ensures that the person detection systems we develop are not only accurate on standard test sets but also robust and reliable across a diverse range of real-world settings and demographics.
-In summary, we present the following novel contributions:
-  • A new person detection dataset, called Wake Vision, which is an easily accessible, permissively licensed dataset that is two orders of magnitude larger than the prior standard.
-  • An improved, high-quality test set that reduces the estimated label error rate to only 2.2% from the prior standard of 7.8%
-  • A suite of five fine-grained benchmark sets that evaluate a model’s fairness and robustness in challenging settings.
-2 Related Work
-The current de facto standard dataset for person detection is the VWW dataset [5, 6, 18, 2, 1].
-To our knowledge, this is the only current open-source dataset designed explicitly for person detection.
-Unfortunately, the VWW dataset is small and cannot be directly downloaded.
-Instead, it must be regenerated from MS-COCO [19], thus limiting its accessibility and usefulness.
-Apart from dedicated person detection datasets, some general image classification datasets contain a person label, which makes it possible to create person detection models based on them.
-Examples include the Cifar-100 dataset [15], and the PASCAL Visual Object Classes dataset [8].
-However, using general image classification datasets can lead to poor perceived performance of TinyML models, as they do not adequately represent the open “no-person” class [5].
-Table 1 shows that our Wake Vision dataset contains up to almost two orders of magnitude more images than any other public and permissively licensed dataset.
-Furthermore, Wake Vision is the only person detection dataset to provide a fine-grained benchmark suite and be officially distributed to popular dataset services such as TensorFlow Datasets (TFDS) and Hugging Face Datasets.
-However, recent research suggests that this focus on top-line metrics can hide poor performance on critical subsets of a dataset, such as subsets relating to sensitive information and fairness [12, 13].
+While existing TinyML datasets [5, 37] have been valuable, they tend to be limited in scale, making it challenging to train production-grade TinyML models.
+This creates a fundamental challenge where we can not develop robust TinyML models due to existing datasets being either too complex for tiny architectures or too limited in scale and diversity.
 Table 1: A comparison of person detection and image classification datasets.
@@
 Dataset Images Train Validation Test Filtering TinyML
-Wake Vision (Quality) 1,322,574 624,115 9,2091 27,881
-Wake Vision (Large) 5,760,428 2,880,214 - -
+Our work – Wake Vision (Quality) 1,322,574 624,115 9,291 27,881
+Our work – Wake Vision (Large) 5,760,428 2,880,214 -- --
 Visual Wake Words [5] $123,287$ $36,000$ $3,926$ $19,107$
@@
 PASCAL VOC 2012 [8] $11,530$ $1,994$ $2,093$-
-3 Wake Vision Dataset
-We introduce the Wake Vision dataset, the largest dataset for TinyML person detection.
-The dataset’s size, quality, and detailed metadata enable new avenues of TinyML research.
-Wake Vision is two orders of magnitude larger than prior comparable datasets and can be adapted based on the target use case.
-We also demonstrate how its quality is significantly improved compared to existing datsets.
-Fig.
-1 illustrates examples of images in the dataset.
-We discuss our methodologies for label generation, data filtering, and error correction, as well as discussing the usability of the dataset and comparing it to the Visual Wake Words Dataset.
+To address this critical gap, we present Wake Vision, a large-scale dataset specifically designed for person detection, the quintessential vision use case for TinyML [1].
+Person detection is a binary image classification task that, while simple, enables crucial applications ranging from occupancy detection [31] and smart HVAC/lighting [39] to acting as an always-on ’wake model’ in larger ML systems [15].
+Wake Vision advances the state-of-the-art by providing 6M images, approximately 100x more than the current leading dataset Visual Wake Words (VWW) [5], while achieving a 1.93% accuracy improvement.
+The dataset is derived from Open Images v7 [21, 19] and is released under a permissive CC-BY 4.0 license to facilitate broad adoption.
+A key contribution of our work is the first systematic investigation of how dataset quality versus quantity impacts TinyML model performance.
+This question is particularly relevant for resource-constrained models, as prior work suggests that large, overparameterized models can effectively identify and adapt to training set errors [40, 9].
+TinyML models, with their limited capacity, may exhibit fundamentally different behavior.
+To explore this hypothesis, we introduce two complementary training sets within the Wake Vision visual corpus: Wake Vision (Large), emphasizing scale, and Wake Vision (Quality), prioritizing label accuracy.
+Our experiments reveal that while Wake Vision (Quality) outperforms its larger counterpart in naive training scenarios, optimal performance is achieved through a training strategy that leverages both sets — using Wake Vision (Large) for pre-training followed by using Wake Vision (Quality) fine-tuning.
+To further explore the dataset quality and quantity trade-off for tinyML, we train a range of ResNet [14] style models across Wake Vision datasets of varying error rates.
+Our experiments collectively reveal that small models are more sensitive to higher error rates than large models and that large models can benefit from larger datasets despite a higher error rate.
+These findings provide insight into the unique challenges of training highly constrained models.
+Beyond raw performance metrics, real-world deployment of TinyML systems often demands robustness across diverse and challenging use case conditions.
+Internet-sourced datasets often underrepresent challenging scenarios, such as low-lighting conditions, leading to overly optimistic performance estimates based on standard test set metrics.
+To enable a rigorous evaluation of real-world capabilities, we additionally develop a comprehensive suite of five fine-grained benchmarks based on our Wake Vision dataset.
+Our benchmarks systematically evaluate performance across multiple dimensions: distance (detection at varying ranges), lighting (performance in well-lit and poorly-lit environments), depictions (handling of drawings and sculptures), perceived gender (investigating potential gender biases), and perceived age (examining effectiveness across age groups).
+This benchmarking framework addresses a gap in current TinyML assessment methods and enables researchers to identify potential biases, limitations, and areas for improvement that might be masked by aggregate metrics alone.
+We find that Wake Vision models outperform VWW on 13 of the 16 fine-grained sets, demonstrating robust performance further illustrating the advantages of a large diverse training dataset.
+In summary, we present the following contributions:
+  • A production grade computer vision dataset for TinyML containing over 6M images (100$\times$ larger than state-of-the-art) for person detection, with comprehensive filtering, permissive licensing, and accessibility through standard ML frameworks.
+  • A cleaned test set with 2.2% label error rate, enabling more reliable evaluation of TinyML models.
+  • A benchmark suite for systematically evaluating model robustness across five critical dimensions (perceived gender, age, distance, lighting, depictions), exposing potential failure modes in real-world deployments.
+  • Evidence that TinyML models are more sensitive to dataset quality than larger models, providing guidance for dataset curation in resource-constrained settings.
+  • A complete, maintained dataset infrastructure including benchmarks, evaluation code, and reproducible filtering pipeline to support ongoing TinyML research.
 (a) Older Person
@@
 (e) Depicted Person
-(f) Young Person
 Figure 1: Sample images from the Wake Vision fine-grained benchmark suite.
+2 Background and Related Work
+Person Detection and Its Role in TinyML
+Person detection is a fundamental computer vision task whose job is to determine the presence or absence of humans in visual data.
+Unlike more complex tasks such as pose estimation or instance segmentation, person detection focuses on the binary classification problem.
+This seemingly simple task has emerged as a canonical use case for TinyML systems [3, 1], driven by both theoretical foundations and practical considerations.
+In the context of resource-constrained computing, TinyML systems demand tasks that carefully balance computational feasibility with real-world utility.
+Person detection represents an optimal compromise.
+It is sufficiently lightweight to be implemented on embedded microcontrollers while maintaining enough discriminative power for practical applications.
+The significance of person detection extends beyond its computational characteristics, serving as a critical selective activation mechanism in resource-constrained systems.
+These lightweight models efficiently manage the activation of power-intensive sensors and processors by determining human presence.
+This architectural pattern has demonstrated particular value in diverse applications, including occupancy detection [31], smart HVAC systems [39], and as an always-on "wake model" in larger ML pipelines [15].
+Moreover, this approach enables substantial energy savings while preserving privacy through on-device processing.
+Existing Person Detection Datasets
+In the TinyML computer vision domain, the VWW dataset [5, 6, 24, 2, 1] has established itself as the de facto standard.
+It remains the only open-source dataset specifically designed for person detection with direct commercial applicability and potential for real-world impact.
+However, VWW faces significant limitations: its small size and indirect access requirements (requiring regeneration from MS-COCO [25]) restrict its accessibility and utility.
+Although other data sets contain person labels within general image classification collections, such as Cifar-100 [20] and PASCAL Visual Object Classes [8], these present significant drawbacks for TinyML applications.
+Specifically, their inadequate representation of the open “no-person” class can lead to poor perceived performance of TinyML models [5].
+Our Wake Vision dataset addresses these limitations by providing nearly two orders of magnitude more images than any existing public, permissively licensed dataset (Table 1).
+Furthermore, it distinguishes itself as the only person detection dataset offering a fine-grained benchmark suite and official distribution through popular services like TensorFlow Datasets (TFDS) and Hugging Face Datasets.
+3 Wake Vision: Dataset and Benchmark Suite
+We introduce Wake Vision, a new dataset for TinyML person detection.
+The dataset’s size, quality, and detailed metadata enable new avenues of TinyML research.
+Wake Vision is almost two orders of magnitude larger than prior comparable datasets and can be adapted based on the target use case.
+Fig.
+1 illustrates examples of images in the dataset.
+In this section, we discuss our methodologies for label generation, data filtering, and error correction, as well as discussing the usability of the data set and comparing it with VWW Dataset.
 3.1 Label Generation
-While a large person detection dataset is indispensable for TinyML research, it can be challenging to obtain.
-Manually labelling millions of images would be prohibitively expensive for a nascent field like TinyML.
-Therefore, we use existing large-scale data efforts to generate Wake Vision automatically.
+A large person detection dataset is indispensable for TinyML research.
+It can, however, be challenging to obtain.
+Manually labeling millions of images would be prohibitively expensive for a nascent field like TinyML.
+Therefore, we bootstrap Wake Vision using existing large-scale data efforts.
 The base label in Wake Vision is a binary person/non-person label.
-We derive these labels from Open Images, which contain both image-level and bounding box labels.
+We derive these labels from Open Images [21, 19], which contain both image-level and bounding box labels of various objects.
 Image-level labels describe which objects are present in an image but not localized to a specific point in the image.
@@
 The bounding box label classes are hierarchically structured so that one class can be a subcategory or a part of another class.
-For example a “Woman” is a subcategory of a “Person,” and a “Human Hand” is a part of “Person.”
+For example a “Woman” is a subcategory of a “Person,” and a “Human Hand” is a part of “Person.” A comprehensive list of Open Images labels related to Wake Vision can be found in Appendix L.
+Users can adapt Wake Vision to meet the needs of specific use cases by adding several configurable options to the open-source dataset creation script.
+For example, a configuration option exists to change whether depictions are labeled as non-persons or excluded from the dataset.
+We refer to Appendix I for details about these configuration options.
 We observe that the bounding box labels are generally more accurate and provide more information for data filtering but are less numerous.
 The Open Images training set has ~9 million images with image-level person labels, but only ~1.7 million images with a person bounding box.
-The Open Images validation and test sets are fully labeled with image-level and bounding box labels.
 Consequently, this presents a trade-off between more data (image-level labels) or higher-quality labels (bounding box labels).
@@
 One larger set, labeled via Open Images image-level labels, and one smaller set, labeled via Open Images bounding box labels (Table 1).
+The Open Images validation and test sets are fully labeled with bounding box labels.
+Therefore, both the Wake Vision validation and test sets are created using the higher-quality bounding box labels.
 Large Training Set from Image-Level Labels
-Image-level labels in Open Images contain little additional information in addition to the labels.
-The only additionally relevant information is a confidence property, which represents how certain it is that a label is correct.
-The confidence ranges from 0 to 1 for labels generated by machines, and is either 0 or 1 for labels negatively or positively verified by humans respectively.
-See section 3.2 for more information.
-We use this property to exclude images from Wake Vision if they contain only low confidence person labels.
+Image-level labels in Open Images contain little additional information than the raw labels.
+The only additional relevant information is a confidence property, which represents how certain it is that a label is correct.
+The confidence ranges from 0 to 1 for labels generated by machines and is either 0 or 1 for labels negatively or positively verified by humans, respectively.
+See Sec.
+3.2 for more information.
+We use this property to exclude images from Wake Vision if they contain only low-confidence person labels.
 Quality Training Set From Bounding Box Labels
@@
 This ensures that bounding box labels are less likely to be false positives.
-Bounding box labels can also be used to calculate an approximation of the area taken up by a person.
-We use this as a proxy for the distance of a person in the image, and exclude images where all person labels are far away, i.e., take up only a small portion of the image.
+Bounding box labels can be used to calculate an approximation of the area taken up by a person.
+We use this as a proxy for the distance of a person in the image, and by default exclude images where all person labels are far away, i.e., take up less than 5% of the image.
 An example of an image with a faraway person can be found in Fig.
@@
 2(b).
-By default we use this flag to make sure that we do not consider a depiction a person.
-We allow users to adapt the Wake Vision dataset to meet the needs of specific usecases by adding several configurable options in the open-source dataset creation script.
-For example a configuration option exists to change whether depictions are labelled as non-persons or excluded from the dataset.
-A comprehensive list of Open Images labels related to Wake Vision can be found in Sec.
-L.
+By default, we use this flag to make sure that we do not consider a depiction a person.
+See Appendix C for an illustration of the bounding box filtering process.
 (a) An example of a person far away
@@
 Figure 2: Examples of challenging outlier images.
-Training Set Comparison
-To compare the performance of the two Wake Vision training sets, we train a MobileNetV2-0.25 [26] model on each dataset for an equal number of steps and compare its performance on the common test set (Table 2).
-The filtered training set generated from the bounding box labels outperforms the larger training set generated from the image-level labels by 4.09% test accuracy.
-This indicates that label quality is more important than quantity in this setting.
-The gap between the two is reduced to just 0.17% when training exclusively on soft labels from a teacher model (MobileNetV2-1.0) trained on bounding box labels.
-Combining the two training sets to train a model where Wake Vision (Large) acts as a pre-training set and Wake Vision (Quality) acts as a fine-tuning set achieves our best performance at a mean top-level test accuracy of 85.72%.
-As the images in Wake Vision (Quality) is a subset of the images in Wake Vision (Large), we do not train on more images than for Wake Vision (Large).
-Table 2: Accuracy on test set trained on image-level and bounding box labels, with and without distillation. A MobileNetV2-1.0 model trained on the bounding box set is used as the teacher in KD.
+3.2 Label Correction
+Label errors are a challenge in computer vision that limit progress and obscure true performance [30, 4].
+This is especially prevalent in large datasets that use complex labeling pipelines to scale while managing costs.
+Recognizing the importance of label quality, we take measures to estimate and improve the accuracy of its labels, aligning with best practices in the field of computer vision.
+Wake Vision’s labels are derived from the Open Images labels; therefore, the errors that occur in the Open Images labeling pipeline are inherited by Wake Vision.
+All labels in the Open Images dataset originate from machine-generated candidate labels.
+Many of these labels are later verified by human annotators, and a subset of them are given bounding boxes and additional annotations.
+While the machine-generated label phase aims to identify objects efficiently, any instances missed during this initial step are unlikely to be captured in downstream phases.
+Consequently, the Open Images dataset may contain numerous false negative labels, referring to images where an object is present but lacks the corresponding label.
+Therefore, additional measures are necessary to identify and correct such labeling omissions to ensure the dataset’s integrity.
+Given Wake Vision’s size, manually correcting label errors across the entire dataset becomes an arduous undertaking.
+Therefore, we prioritize labeling the Wake Vision validation and test sets so that the dataset can be used to accurately evaluate a model’s performance.
+We use the Scale Rapid tool from Scale AI (https://scale.com/) to relabel the Wake Vision validation and test sets.
+Details about the relabelling process, including instructions given and total cost, can be found in Appendix N.
+With the ground truth established in the validation and test sets, Wake Vision becomes an ideal foundation for research on automated data cleaning techniques [29].
+We perform an initial exploration of these methods in appendix O.
+We report our estimated label error rate after label corrections versus VWW in Table 2.
+These estimates are all based on a random subset of 500 samples from each dataset, which the authors manually relabelled according to whether we saw a non-depiction person or body part in the image.
+The Wake Vision validation and test set is considerably lower than that of the VWW dataset top-level error rate.
+Table 2: Label error rate of VWW and Wake Vision after label correction.
+The Wake Vision error rate is broken down by set.
+The Wake Vision Train (Quality) and Train (Large) have different error rates due to the different label sources.
+The Wake Vision Validation and Test sets have a lower error rate due to manual relabeling.
+Dataset Wake Vision Set Label Error
+Visual Wake Words Train, Val, & Test 7.8%
+Wake Vision Train (Large) 15.2%
+Wake Vision Train (Quality) 6.8%
+Wake Vision Val & Test 2.2%
+3.3 Fine-grained Benchmark Suite
+While Wake Vision provides a substantial improvement in scale and quality over existing datasets, we recognize that overall test set performance alone may not capture the nuanced challenges of real-world deployment for TinyML applications.
+Internet-sourced images typically exhibit a distinct bias towards well-composed photographs—they are often well-lit, properly framed, and feature clearly visible subjects.
+This stands in stark contrast to the challenging conditions where TinyML person detection systems are typically deployed, such as varying lighting conditions, extreme viewing distances, or partial occlusions [32].
+To bridge this gap between benchmark performance and practical utility, we augment Wake Vision with a comprehensive fine-grained benchmark suite.
+This suite evaluates model robustness across challenging real-world scenarios where traditional accuracy metrics might mask significant failure modes.
+The benchmarks assess performance across critical dimensions, including lighting conditions, subject distance, and demographic attributes, enabling developers to identify potential biases or limitations during the design phase rather than after deployment.
+By providing these targeted evaluation sets, we enable a more nuanced analysis of model behavior beyond aggregate accuracy metrics.
+The suite consists of five fine-grained benchmark sets.
+Each of the five fine-grained benchmarks has been chosen based on a combination of its relevance to TinyML use cases and the availability of requisite metadata to generate the sets.
+Each benchmark set is a subset of the respective validation or test set filtered based on the criteria under test.
+The benchmark determines whether a model is sufficiently accurate in the planned deployment setting.
+For example, a model designer may make different design choices for a use case where the subject is close to the camera and well lit compared to the inverse setting.
+Appendix D contains a case study that shows how fine-grained evaluation sets can be used to design robust models.
+Perceived Gender and Age
+Table 3: Accuracy on the Wake Vision (WV) and VWW test sets by models trained on the VWW, Wake Vision (Quality) and Wake Vision (Combined) training sets.
+Train
+VWW Wake Vision (Quality) WV (Combined)
+Test VWW 88.33$\pm$0.2988.59$\pm$0.1789.34$\pm$0.02%
+WV 83.79$\pm$0.2384.89$\pm$0.1185.72$\pm$0.04%
+Underrepresented sub-groups typically constitute a small portion of a generic test set; therefore, top-line metrics often obscure a bias in a model until it is deployed [16].
+This benchmark, generated from the Open Images More Inclusive Annotations for People (MIAP) extended fairness labels [34], evaluates a model separately on demographics that are underrepresented in the underlying dataset distribution, identifying bias.
+These labels are based on perceived gender and age representation and are not necessarily representative.
+Distance
+This tests how the distance of people in images impacts model performance.
+If a person detection system is intended to recognize subjects at great distances, the system’s performance on the faraway dataset will be more informative than its performance on the top-level test set.
+We created three data sets based on the percentage of the image the subject bounding box covers.
+The three sets are near (>60%), at a medium distance (10-60%), and far away (<10%).
+Lighting
+The lighting data sets test the performance of Machine Learning (ML) systems in different lighting conditions.
+In scenarios like security monitoring, outdoor robotics, or augmented reality applications, models must be robust to varying lighting conditions, including low-light environments.
+We create three fine-grained datasets of this type for dark, normal, and bright lighting conditions, respectively.
+We quantify lighting conditions by the average pixel values of images in greyscale, a simple but effective method for distinguishing lighting conditions [41].
+We define low as an average pixel value less than 85, normal as between 85 and 170, and bright lighting conditions as greater than 170.
+Depictions
+A particularly challenging task for a person detection model is to correctly reject depictions of people.
+In many use cases, a person detection model should not trigger on a depiction.
+For example, a room occupancy detector could incorrectly identify a painting on the wall as a person.
+This benchmark measures a model’s accuracy on three related sets of non-person samples: depictions of people, depictions that are not of people, and images that do not contain a depiction of any kind.
+Depictions of people can range from photo-realistic to crude stick figures.
+3.4 Usability
+Wake Vision and its fine-grained benchmark suite are available through TensorFlow Datasets [36] and HuggingFace Datasets [22] to enable easy access and use by the community.
+The images and labels are rehosted to ensure the dataset will not be changed over time due to dead links, making it a more stable benchmark.
+The rehosted labels are generated according to our default dataset configuration, further described in Appendix I.
+4 Results
+In this section, we present a comprehensive evaluation of Wake Vision through five complementary analyses.
+First, we provide the first insights into best practices for Wake Vision by comparing its two training sets, which also provide initial insights into the trade-off between dataset quality and size in a TinyML setting.
+Second, we show that Wake Vision serves as an effective drop-in replacement for Visual Wake Words by conducting cross-evaluation studies using identical MobileNetV2 models.
+Third, we validate Wake Vision’s benefits across diverse model architectures commonly used in the TinyML community, from MCUNet to ColabNAS families.
+Fourth, we thoroughly investigate the critical trade-off between dataset quality and size for TinyML models, providing insights into how model capacity influences sensitivity to label noise versus dataset scale.
+Finally, we present fine-grained benchmark results across demographics, environmental conditions, and visual contexts to assess model robustness in real-world scenarios.
+Throughout these analyses, we demonstrate that Wake Vision not only improves upon VWW’s performance but also enables a more thorough evaluation of TinyML computer vision models.
+4.1 Training Set Evaluation
+Table 4: Accuracy on test set of a MobileNetV2-0.25 trained on image-level and bounding box labels, with and without distillation.
+A MobileNetV2-1.0 model trained on the bounding box set is used as the teacher in KD.
 Training set Label Source Training Set Size Accuracy Distilled Accuracy
@@
 Wake Vision (Combined) Both 5,760,428 85.72$\pm$0.04-
-As our distillation results indicate, further data-centric filtering techniques may be needed to fully leverage the image-level training set.
-We experimented with tripling the model size and doubling the training time, but the difference in accuracy remained relatively consistent.
-The two training sets provide a useful foundation for future work on the importance of data quality vs.
-quantity.
-Wake Vision can enable further research on data-centric TinyML [21, 22], specifically uncovering the relationship between model capacity and the sensitivity to training data quality.
-3.2 Label Correction
-Label errors are a challenge in computer vision [24, 4] that limit progress and obscures true performance.
-This is especially prevalent in large datasets that use complex labeling pipelines to scale while managing costs.
-Recognizing the importance of label quality, we take measures to estimate and improve the accuracy of its labels, aligning with best practices in the field of computer vision.
-Wake Vision’s labels are derived from the Open Images labels; therefore, the errors that occur in the Open Images labeling pipeline are inherited by Wake Vision.
-All labels in the Open Images dataset are originally machine-generated candidate labels.
-Many of these labels are later verified by human annotators, and a subset of them are given bounding boxes and additional annotations.
-While the machine-generated label phase aims to identify objects efficiently, any instances missed during this initial step are unlikely to be captured in downstream phases.
-Consequently, the Open Images dataset may contain numerous false negative labels, referring to images where an object is present, but lacks the corresponding label.
-Therefore, additional measures are necessary to identify and correct such labeling omissions to ensure the dataset’s integrity.
-Given Wake Vision’s size, manually correcting label errors across the entire dataset becomes an arduous undertaking.
-Therefore, we prioritize labeling the Wake Vision validation and test sets so that the dataset accurately represents the model’s performance.
-To do so we make use of the Scale AI platform to crowd-source manual label corrections (see Appendix).
-With the ground truth established in the validation and test sets, Wake Vision becomes an ideal foundation for research on automated data cleaning techniques [23].
-We perform an initial exploration of these methods in appendix N.
-We report our estimated label error rate after label corrections versus VWW in Table 3.
-The Wake Vision validation and test set is considerably lower than that of the VWW dataset top-level error rate.
-Table 3: Estimated label error rate of VWW and Wake Vision, based on a 500 sample subset.
-The Wake Vision estimated error rate is broken down by set.
-The Wake Vision Train (Quality) and Train (Large) have different error rates due to the different sources of their labels.
-The Wake Vision Validation and Test sets have a lower error rate due to manual relabeling.
-Dataset Wake Vision Set Label Error Rate
-Visual Wake Words Train, Val, & Test 7.8%
-Wake Vision Train (Large) 15.2%
-Wake Vision Train (Quality) 6.8%
-Wake Vision Val & Test 2.2%
-3.3 Usability
-Wake Vision is available through TensorFlow Datasets [28] and HuggingFace Datasets [17] to enable easy access and use by the community.
-The images and labels are rehosted to ensure the dataset will not be changed over time due to dead links, making it a more stable benchmark.
-The rehosted labels are generated according to our default dataset configuration, further described in Sec.
-I.
-3.4 Comparison to Visual Wake Words
-Wake Vision serves as a drop-in improvement over VWW, thanks to its size and our extensive filtering.
-To compare the benefits of Wake Vision against VWW, we train two identical MobileNetv2 [26] models with a width modifier of 0.25 for 200,000 steps on 224x224x3 images using AdamW [20].
+To evaluate the performance of the two Wake Vision training sets, we train identical models on each set.
+We use MobileNetV2 [33] models with a width modifier of 0.25 for 200,000 steps on 224x224x3 images using AdamW [26].
 We use a lr of 0.002 with a cosine decay and a weight decay of 4E-6.
-One model uses VWW’s training set, and the other employs Wake Vision’s Quality training set.
-After training these identical models using the respective datasets’ training sets with the same recipe for an equal number of steps, we cross-evaluate their performance on the corresponding test sets.
-Table 4 shows a 0.26% improvement on VWW’s own test set, which illustrates that Wake Vision is a direct improvement over VWW, not simply a domain shift.
-We additionally achieve a 1.1% improvement over VWW on Wake Vision’s Test set, indicating that the new test set is more challenging.
-Table 3 also shows the performance of a the Wake Vision (Combined) model described in section 3.1.
-This model achieves an average 0.75% improvement over the Wake Vision (Quality) trained model on the VWW test set and an average 0.83% improvement on the Wake Vision test set.
-The Wake Vision (Combined) model is trained for longer than the VWW and Wake Vision (Quality) models, but shows the value of using the training sets in unison even on the VWW test set.
-Table 4: Accuracy on the Wake Vision and VWW test sets by models trained on the VWW, Wake Vision (Quality) and Wake Vision (Combined) training sets.
-Train
-VWW Wake Vision (Quality) Wake Vision (Combined)
-Test VWW 88.33$\pm$0.29 88.59$\pm$0.1789.34$\pm$0.02%
-Wake Vision 83.79$\pm$0.2384.89$\pm$0.1185.72$\pm$0.04%
-3.5 Dataset Quality Vs. Size
-We investigate the impact of dataset quality and size of Resnet [10] style models of varying capacity.
-We measure dataset quality by the approximate rate of label errors.
+After training, we compare each model’s performance on the common test set.
+The result of this comparison is shown in Tab.
+4.
+These results show that the Wake Vision (Quality) training set outperforms the Wake Vision (Large) training set by 4.09% test accuracy.
+This indicates that label quality is more important than quantity in this setting.
+The gap between the two is reduced to just 0.17% when training exclusively on soft labels from a teacher model (MobileNetV2-1.0) trained on bounding box labels.
+Using the two training sets in unison to train a model where Wake Vision (Large) acts as a pre-training set, and Wake Vision (Quality) acts as a fine-tuning set achieves our best performance at a mean top-level test accuracy of 85.72%.
+The two training sets together provide a foundation for further research on data-centric TinyML [27, 28].
+4.2 Wake Vision & VWW Cross Evaluation
+A key challenge in advancing TinyML research is ensuring that improvements can be readily adopted by the community without requiring significant modifications to existing workflows or architectures.
+While novel datasets can offer better quality or scale, their practical impact is limited if they require substantial changes to model architectures, training pipelines, or deployment processes.
+Therefore, we designed Wake Vision to serve as a drop-in replacement for VWW.
+To validate this drop-in capability and demonstrate Wake Vision’s effectiveness, we do a cross-evaluation study.
+We train two identical MobileNetv2 [33] models according to the same training recipe as in Sec.
+4.1.
+One model uses VWW’s training set, and the other uses Wake Vision’s Quality training set.
+After training these identical models using the respective datasets’ training sets with the same recipe for an equal number of steps, we evaluate their performance on the corresponding test sets.
+Creftab:wv-vww-cross-val shows the result of the cross-evaluation.
+The Wake Vision trained model outperforms the VWW trained model with a 0.26% improvement on VWW’s own test set, which indicates that Wake Vision is a direct improvement over VWW, not simply a domain shift.
+Also, we achieve a 1.1% improvement over VWW on Wake Vision’s Test set, indicating that the new test set is more challenging.
+Table 3 also shows the performance of the Wake Vision (Combined) model described in Sec.
+3.1.
+This model achieves 0.75% improvement over the Wake Vision (Quality) trained model on the VWW test set and 0.83% improvement on the Wake Vision test set.
+The Wake Vision (Combined) model is trained for longer than the VWW and Wake Vision (Quality) models but shows the value of using the training sets in unison even on the VWW test set.
+4.3 Comprehensive Model Architecture Evaluation
+To show that the benefits of using Wake Vision over VWW are not limited to MobileNetv2 architectures, we also train several other models on both training sets, Wake Vision (Quality) and VWW, and compare their performance.
+For this experiment, we use a collection of models commonly used in the TinyML community, such as models from the MCUNet [24] and Micronets [2] families.
+Furthermore, we include models from ColabNAS [11] to provide evaluation results on even smaller models.
+Figure 3: Cross-evaluation results on the Wake Vision test set.
+Each model is trained on each dataset using the Adam optimizer with a learning rate of 0.001 for 243,800 steps at a batch size of 512, equivalent to 100 epochs on the Wake Vision (Quality) training set.
+In addition, we apply random horizontal flips rotations and use 8-bit quantization-aware training.
+We use the checkpoint with the best validation score for testing.
+In our comparison, we cross-evaluate each trained model’s performance on the test sets of both VWW and Wake Vision.
+Results on the Wake Vision test set can be found in Fig.
+3, while the results for the VWW test set can be found in Fig.
+5 alongside the complete data in Appendix A.
+For the majority of models, the one trained on the Wake Vision (Quality) set achieves a higher accuracy on both the VWW and Wake Vision test set, which supports our claim that Wake Vision is a direct improvement over VWW.
+In some models, the VWW trained variant performs better on the VWW test set.
+This is only the case for the smallest models in our evaluation suite, which we believe may be due to these models being too small to generalize past the slight domain shift between the VWW and Wake Vision test sets.
+We additionally see that the accuracy discrepancy between identical models is often highest on Wake Vision’s Test set, further suggesting that the new test set is more challenging.
+The boost in accuracy achieved by using the Wake Vision (Quality) Training set can be traded off for smaller model architectures (in some cases 3X fewer MACs) while preserving test accuracy.
+Accuracy and performance metrics, such as latency or model size, are in constant tension in TinyML models.
+Therefore, any boost in accuracy that does not impact the model’s architecture enables further efficiency.
+4.4 Impact of Dataset Quality & Size in TinyML
+Understanding the relationship between dataset quality and size is important for training efficient models in real-world scenarios.
+While gathering large amounts of data has become easier, ensuring high-label quality remains costly and time-consuming.
+This creates an important trade-off: is it better to have a smaller dataset with high-quality labels or a larger dataset with more label noise?
+The answer likely depends on the capacity of the model and the specific task.
+We investigate the impact of dataset quality and size on Resnet [14] style models of varying capacity.
+We measure quality by the approximate rate of label errors.
 The Wake Vision (Quality) training set has an estimated error rate of around 7%.
 We simulate higher error rate datasets (15% and 30%) by flipping the binary label with a certain probability.
-Appendix D gives the derivation for the flip probability.
+Appendix E gives the derivation for the flip probability.
 We also sweep the dataset size by taking a slice of Wake Vision (Quality).
 Each model is trained for 50,000 steps on 224x224x3 images using AdamW [26], a learning rate of 0.001 with a cosine schedule and a weight decay of 0.004.
-Figure 3 illustrates the experiment’s results.
-We observe that smaller models (leftmost figure) are more sensitive to higher error rates than large models.
-The smallest model’s accuracy drop is 1.3%, going from the base train (quality) error rate of around 7% to 15% compared to only a 0.5% accuracy loss for the largest model.
-In contrast, large models benefit more from big datasets and start to over fit on the smaller slices of the training data.
-Figure 3: Comparing the impact of dataset error rate and size on ResNet models of varying capacity.
+Figure 4 shows that smaller models (leftmost figure) are more sensitive to higher error rates than large models, similar to prior findings that model capacity can help memorize and ignore errors and outliers [40, 9, 13].
+The lowest accuracy drop in the smallest model is 1.
+3%, going from the base train (quality) error rate of around 7% to 15% compared to only a 0.5% accuracy loss for the largest model.
+In contrast, large models benefit more from big datasets and start to overfit on the smaller slices of the training data [18].
+We also see the large slice of the 15% error rate dataset outperforming a smaller slice of the 7% error rate dataset.
+Figure 4: Impact of dataset error rate and size on models of varying capacity.
 The $y$-axis is the Wake Vision test accuracy and the $x$-axis is the percentage of Wake Vision Train (Quality) used in training.
 The error % indicates the expected rate of incorrectly labeled samples.
-4 Fine Grain Benchmark Suite
-Test sets generated directly from images posted to the internet are not reflective of real world use cases as the distribution is biased towards images people deems as worth sharing (e.g.
-well lit and framed).
-In contrast, person detection systems are deployed, and expected to work, in more challenging settings, such as low lighting.
-Due to this gap, a model that achieves high test accuracy may perform poorly in real-world scenarios after the design stage has ended.
-To address this issue, we present a benchmark suite for person detection that tests a model’s robustness in challenging settings, enabling better analysis at the design stage.
-Additionally, the suite measures the model’s performance across demographics to ensure it does not exhibit clear bias.
-The suite consists of five fine-grained benchmark sets.
-Each of the five fine-grained benchmarks have been chosen based on a combination of its relevance to TinyML use cases and the availability of requisite metadata to generate the sets.
-Each benchmark set is a subset of the respective validation or test set filtered based on the criteria under test.
-The benchmark determines whether a model is sufficiently accurate in the planned deployment setting.
-For example, a model designer may make different design choices for a use case where the subject is close to the camera and well-lit compared to the inverse setting.
-Appendix C contains a case study showing how the fine grained evaluation sets can be used to design robust models.
-Table 5 reports the number of samples in each of the individual benchmark sets and the F1 scores of a Wake Vision and VWW model on each set.
-Perceived Gender and Age
-These fine grained datasets are generated from the Open Images More Inclusive Annotations for People (MIAP) extended fairness labels [27].
-Underrepresented sub-groups typically constitute a small portion of a generic test set; therefore, top-line metrics often obscure a bias in a model until it is deployed [12].
-This benchmark aims to evaluate a model separately on demographics that are underrepresented in the underlying dataset distribution, identifying bias.
-These labels are based on a perceived gender and age representation and are not necessarily representative.
+4.5 Fine-Grained Benchmark Evaluation
+Although aggregate metrics like overall accuracy or F1 score provide a high-level view of model performance, they mask critical failure modes and biases that emerge in real-world scenarios.
+For this reason, we included fine-grained benchmarks in Wake Vision (Section 3.3).
+Fine-grained evaluation across demographics, environmental conditions, and visual contexts helps us develop reliable TinyML Computer Vision systems for bespoke use cases.
+This is particularly important for person detection models deployed in real-world applications, where failures on certain subgroups or conditions could have serious implications for fairness and safety.
+In Table 5, we compare the models from Sec.
+4.2, trained on Wake Vision and VWW, across various scenarios to understand their relative strengths and limitations.
+The Wake Vision model exhibits superior robustness across the challenging settings exercised by our benchmarking suite, further demonstrating the effectiveness of the dataset.
+For instance, on the “Depictions” benchmark, which evaluates performance on images containing persons, non-person objects, or no depictions, the Wake Vision model achieves an F1 score of 0.71 for person depictions, outperforming the VWW model’s 0.66.
+Similarly, the Wake Vision model scores 0.86 for non-person depictions compared to the VWW model’s 0.82.
 Table 5: Wake Vision Fine Grain Benchmark Suite. We report the samples in each set and the average F1 score across three Wake Vision models and three VWW models on the benchmark set.
@@
 Benchmark Val Test Wake Vision VWW
-Gender Female 684 2,157 0.93 0.89
-Male 1,310 3,918 0.91 0.88
-Unknown 1,612 4,940 0.77 0.78
+Gender Female $684$ $2157$ $0.93$ $0.89$
+Male $1310$ $3918$ $0.91$ $0.88$
+Unknown $1612$ $4940$ $0.77$ $0.78$
 Age Young $275$ $884$ $0.94$ $0.90$
-Middle 2,133 6,595 0.91 0.88
+Middle $2133$ $6595$ $0.91$ $0.88$
 Older $90$ $276$ $0.94$ $0.89$
-Unknown 1,299 3,837 0.71 0.74
-Distance Near 5,457 16,333 0.91 0.89
-Medium 2,213 6,876 0.85 0.84
-Far 398 1,140 0.59 0.67
-Lighting Dark 3,255 9,420 0.85 0.81
-Normal 14,315 43,010 0.82 0.82
-Bright 1,012 3,332 0.82 0.82
+Unknown $1299$ $3837$ $0.71$ $0.74$
+Distance Near $5457$ 16 3331633316\,33316 333 $0.91$ $0.89$
+Medium $2213$ $6876$ $0.85$ $0.84$
+Far $398$ $1140$ $0.59$ $0.67$
+Lighting Dark $3255$ $9420$ $0.85$ $0.81$
+Normal 14 3151431514\,31514 315 43 0104301043\,01043 010 $0.82$ $0.82$
+Bright $1012$ $3332$ $0.82$ $0.82$
 Depictions Person $356$ $978$ $0.71$ $0.66$
 Non-Person $352$ $1101$ $0.86$ $0.82$
-No Depiction 8,583 25,802 0.87 0.85
-Distance
-The distance datasets test how the distance of persons in images impact the performance of ML systems.
-This benchmark is critical for correctly predicting a model’s deployed performance.
-If a person detection system is intended to recognize subjects at great distances, the system’s performance on the far away dataset will be more informative than its performance on the top-level test set.
-We create three datasets based on the percentage of the image the subject bounding box covers.
-The three sets are near (>60%), at a medium distance (10-60%), and far away (<10%).
-Lighting
-The lighting datasets aim to test the performance of ML systems across different lighting conditions.
-In scenarios like security cameras, outdoor robotics, or augmented reality applications, models must be robust to varying lighting conditions, including low-light environments.
-We create three fine grained datasets of this type for dark, normal, and bright lighting conditions, respectively.
-We quantify lighting conditions by the average pixel values of images in greyscale, which is a simple but effective method for distinguishing lighting conditions [33].
-We define low as an average pixel value less than 85, normal as between 85 and 170, and bright lighting conditions as greater than 170.
-Depictions
-A particularly challenging task for a person detection model is to correctly reject depictions of people.
-In many use cases a person detection model can not falsely trigger on a depiction.
-For example, a room occupancy detector that incorrectly identifies a painting on the wall as a person.
-This benchmark measures the models accuracy on three related sets of non-person samples: depictions of people, depictions that are not of people, and images do not contain a depiction of any kind.
-Depictions of people can range from photo-realistic to crude stick figures.
-4.1 Benchmark Results
-Table 5 compares the results of a model trained on the Wake Vision (Quality) training set to an identical model trained on VWW.
-The Wake Vision model exhibits superior robustness across the challenging settings exercised by our benchmarking suite, further demonstrating the effectiveness of the dataset.
-For instance, on the “Depictions” benchmark, which evaluates performance on images containing persons, non-person objects, or no depictions, the Wake Vision model achieves an F1 score of 0.71 for person depictions, outperforming the VWW model’s 0.66.
-Similarly, the Wake Vision model scores 0.86 for non-person depictions compared to the VWW model’s 0.82.
+No Depiction $8583$ 25 8022580225\,80225 802 $0.87$ $0.85$
 The Wake Vision model also showcases improved performance on the “Age” benchmark, with F1 scores of 0.94, 0.91, and 0.94 for young, middle, and older individuals, surpassing the VWW model’s scores of 0.90, 0.88, and 0.89, respectively.
@@
 In addition, the Wake Vision model demonstrates resilience to challenging lighting conditions, achieving F1 scores of 0.85 in dark lighting scenarios, compared to the VWW’s 0.81.
+Notably, a VWW model only significantly outperforms a Wake Vision model on the far distance dataset at a score of 0.67 and 0.59, respectively.
+This is likely caused by a difference in how the two datasets are created.
+Where VWW filters out persons smaller than 0.5% of the image, this filter is set to 5% for Wake Vision.
+Therefore, the VWW dataset contains images with significantly smaller persons than the Wake Vision dataset, which can improve its performance on far away persons.
+We did experiment with setting the limit to 0.5% for Wake Vision however, we experienced that it disproportionately decreased the performance of models trained on the dataset in other settings.
 Collectively, these results underscore the significance of Wake Vision’s large-scale, diverse data and fine-grained benchmarks, enabling the development of more robust and reliable TinyML person detection models across a wide range of realistic and real-world scenarios.
 5 Ethical Considerations
-The Wake Vision dataset and benchmark suite aim to advance TinyML research by providing large-scale datasets and enabling fine-grained analysis of TinyML systems.
-Wake Vision can directly improve the quality of person detection systems that benefit society, such as energy-saving applications, while protecting user privacy through on-device inference.
-However, person detection models can also be used for malicious purposes, such as weapons targeting or large-scale surveillance.
+The Wake Vision dataset and benchmark suite are designed to advance TinyML research by providing data resources and enabling detailed evaluation of person detection systems.
+While this work has significant potential benefits, we acknowledge several important ethical considerations:
+Intended Benefits
+Wake Vision can improve the quality of person detection systems that serve beneficial societal purposes, for example, an occupancy detection system that conserves energy by turning off the lights/HVAC/TV when a person leaves.
+The focus on TinyML and on-device inference helps protect user privacy by avoiding cloud-based processing of sensitive data.
+Potential Misuse
+We recognize that person detection technology can be misappropriated for harmful purposes, including weapons targeting or mass surveillance systems.
+While we cannot completely prevent misuse, we implement several safeguards, including releasing only evaluation sets, not training sets, for our fine-grained benchmark datasets.
+Data Rights and Privacy
 The images in Wake Vision are sourced from Flickr through Open Images [21, 19] under the CC-BY 2.0 license.
-While we have made efforts to ensure the images are properly licensed, we cannot guarantee the license status of each image.
-It is possible that some images have been uploaded without the right to distribute under the CC-BY 2.0 license.
-Given sufficient resources, individuals in the dataset could potentially be identified based on biometric data.
-The dataset may also contain offensive content, as manually verifying millions of images is infeasible.
-The benchmark suite aims to ensure the fairness and robustness of person detection models.
-However, the demographic benchmarks could potentially be misused to create systems that classify gender and age, breaching personal privacy or perpetuating discrimination.
-To mitigate this risk, we have not released training sets for the fine-grained benchmark datasets.
-6 Conclusions
-The lack of large, high-quality datasets has been a significant bottleneck in the field of TinyML research.
-To address this issue, we introduced Wake Vision, a large-scale dataset and benchmark suite for person detection that surpasses the current state-of-the-art dataset, VWW, in terms of size (~100x larger), ease of use, label accuracy, and real-world relevance.
-These improvements make Wake Vision a drop-in replacement for VWW.
-Recognizing the importance of fairness across demographics and robustness to challenging settings in person detection use cases, we also released a Wake Vision benchmark suite comprising six fine-grained benchmark sets to evaluate the fairness and robustness of person detection models.
-Furthermore, our investigation into the trade-off between dataset quality and quantity for TinyML systems suggests that dataset quality is likely more crucial for TinyML systems compared to conventional ML systems.
-The Wake Vision dataset and benchmark suite aim to facilitate the development of more accurate, fair, and robust person detection models in TinyML.
-7 Acknowledgements
-This work was partially supported by the Google TPU Research Cloud program. We would like to thank Andrew Howard of Google Deepmind for his input and support in this effort.
-Appendix A Code Repository
-The code used to generate the Wake Vision dataset is available at the following GitHub repository: https://github.com/harvard-edge/Wake_Vision
+While we have made extensive efforts to ensure proper licensing, we acknowledge the possibility that some images may have been uploaded without proper distribution rights.
+Furthermore, the dataset may contain potentially identifying biometric information, and given sufficient resources, individuals could potentially be identified.
+Due to the scale of the dataset, exhaustive manual verification for offensive content or privacy concerns is infeasible.
+Fairness and Representation
+Our benchmark suite is specifically designed to evaluate model fairness and robustness across different demographics and conditions.
+However, we acknowledge that these same demographic benchmarks could be misused to develop systems that classify gender and age without consent, potentially enabling privacy violations or discriminatory practices.
+We have carefully balanced these concerns against the importance of ensuring that person detection systems work equitably for all populations.
+6 Conclusion
+The lack of large, high-quality datasets has been a significant bottleneck in TinyML research.
+Through Wake Vision, we address this challenge by providing a comprehensive dataset and benchmark suite that advances the field in several key dimensions.
+Wake Vision surpasses the current state-of-the-art VWW dataset by offering  100x more images while maintaining ease of use and improving label accuracy.
+Our experiments show that Wake Vision serves as an effective drop-in replacement for VWW, consistently improving performance across various model architectures and training scenarios.
+Moreover, our investigation into dataset quality versus quantity trade-offs reveals important insights specific to TinyML systems.
+Unlike large models that can effectively learn from noisy data, we find that TinyML models are particularly sensitive to label quality because of their limited capacity.
+However, we also demonstrate that carefully designed training strategies—such as our approach of pre-training on Wake Vision (Large) followed by fine-tuning on Wake Vision (Quality)—can effectively leverage both quantity and quality to achieve superior performance.
+Finally, the Wake Vision benchmark suite, with its fine-grained evaluation sets, establishes new standards to evaluate the robustness of the TinyML model in demographics, environmental conditions, and visual contexts.
+These benchmarks enable researchers to identify and address potential biases and failure modes that could otherwise go undetected using traditional evaluation metrics.
+Appendix A Additional Benchmark Results
+Table Tab.
+6 provides input shape, Flash, RAM, and MACs for each model presented in section 4.2 as well as the mean and standard deviation of test accuracies for both Wake Vision and VWW datasets.
+RAM and Flash have been measured using the “stm32tflm” utility of X-CUBE-AI 8.1.0, whereas MACs with “stm32ai”.
+Links pointing to the original models used for training are also present for each model family involved in the experiment.
+Figure Fig.
+5 shows cross-evaluation results on the Visual Wake Words test set.
+The majority of models trained on the Wake Vision dataset obtain better test accuracy on the VWW than the ones trained on the VWW.
+Only the smallest models do not achieve a higher test accuracy when trained on the Wake Vision, likely due to the lack of enough resources to overcome the slight domain shift between the two datasets.
+Model input RAM Flash MACs Test Set Training Set
+.tflite volume [kiB] [kiB] [MM] WV (Qual.) VWW
+MobileNetV2  [^1] https://keras.io/api/applications/mobilenet/[33]
+MobileNetV2_0.25 (224,224,3) 1,244.5 410.55 36,453,732 WV $84.9\pm 0.11$ $83.8\pm 0.23$
+VWW $88.6\pm 0.17$ $88.3\pm 0.29$
+MCUNet  [^2] https://github.com/mit-han-lab/mcunet/blob/master/mcunet/model_zoo.py[24]
+10fps_vww (64,64,3) 168.5 533.84 5,998,334 WV $81.7\pm 0.28$ $77.6\pm 0.31$
+VWW $81\pm 0.12$ $80.8\pm 0.1$
+5fps_vww (80,80,3) 226.5 624.55 11,645,502 WV $82.9\pm 0.29$ $79.6\pm 0.26$
+VWW $83.2\pm 0.57$ $82.6\pm 0.16$
+320kb-1mb_vww (144,144,3) 393 923.76 56,022,934 WV $85.6\pm 0.34$ $82.4\pm 0.62$
+VWW $86.5\pm 1.01$ $86\pm 0.25$
+Micronets [^3] https://github.com/ARM-software/ML-zoo/tree/master/models/visual_wake_words[2]
+vww2_50_50_INT8 (50,50,1) 71.50 225.54 3,167,382 WV $71.9\pm 0.67$ $66.2\pm 0.88$
+VWW $70.8\pm 0.68$ $70.6\pm 0.64$
+vww3_128_128_INT8 (128,128,1) 137.50 463.73 22,690,291 WV $77.8\pm 0.56$ $72.6\pm 1.01$
+VWW $78.3\pm 0.91$ $77.6\pm 1.13$
+vww4_128_128_INT8 (128,128,1) 123.50 417.03 18,963,302 WV $77.9\pm 0.6$ $71.3\pm 1.03$
+VWW $78.4\pm 0.51$ $76.5\pm 0.37$
+ColabNAS444 https://github.com/AndreaMattiaGaravagno/ColabNAS/tree/main/models [11]
+k_2_c_3 (50,50,3) 18.5 7.66 250,256 WV $70.6\pm 0.96$ $69.3\pm 0.97$
+VWW $65.6\pm 0.66$ $70.7\pm 0.08$
+k_4_c_5 (50,50,3) 22 18.49 688,790 WV $75.7\pm 0.18$ $74\pm 0.23$
+VWW $69.9\pm 0.26$ $75.5\pm 0,64$
+k_8_c_5 (50,50,3) 32.5 44.56 2,135,476 WV $77.3\pm 0.37$ $75\pm 0.15$
+VWW $73\pm 0.91$ $77.3\pm 0.57$
+Table 6: Raw data of reconstructed models.
+Figure 5: Cross-evaluation results on the Visual Wake Words test set
+Appendix B Code Repository
+The code used to generate the Wake Vision dataset is available at the following GitHub repository: https://anonymous.4open.science/r/Wake_Vision-DD19/README.md
 This repo contains the code to generate Wake Vision and the benchmark suite, as well as the code to train and evaluate models. This code is sufficient to reproduce all results in the paper.
-Appendix B Flowchart of Bounding Box Filtering Process
+Appendix C Flowchart of Bounding Box Filtering Process
 Fig. 6 illustrates the filtering process for Wake Vision when using Open Images’ bounding box labels as the label source.
@@
 When not specified the image size is 160 and the width multiplier is 1.0.
-Appendix C Model Design Case Study
+Appendix D Model Design Case Study
 Basic test set performance can often misrepresent a model’s performance, given the typical domain shift between images scraped from the internet and real-world use cases.
@@
 A design decision might have seemingly little impact on the test accuracy but may destroy real-world performance depending on the deployment environment.
-For example, a person detection system may only operate in dark lighting conditions, but the test dataset has an insignificant number of dark samples; therefore the test accuracy will not reflect the real world accuracy.
+For example, a person detection system may only operate in dark lighting conditions, but the test dataset has an insignificant number of dark samples; therefore, the test accuracy will not reflect the real-world accuracy.
 The benchmark suite enables more holistic analysis during the design phase.
@@
 While these results can inform ML practitioners, our intention is to demonstrate the usefulness of the benchmark suite.
-C.1 Image Size vs. Model Width Scaling
+D.1 Image Size vs. Model Width Scaling
 We train two series of MobileNetV2 models: one series that sweeps the input image size [64-256] and one that sweeps the width multiplier of a model [0.1-1.5].
 We then benchmark these models on the Wake Vision test set as well as the far distance benchmark.
-We plot these results against the number of multiply accumulate (MAC) operations in the model as a proxy for on-device latency [2].
+We plot these results against the number of multiply-accumulate (MAC) operations in the model as a proxy for on-device latency [2].
 The results in Table 7 (left) show that when looking exclusively at the high-level metric (i.e., test accuracy), scaling the input image size has a similar impact as scaling the model size.
@@
 The distance benchmark in Wake Vision effectively unveils the disproportionate impact of image size on model performance for distant subjects, enabling more informed decision-making during model optimization.
-C.2 Quantization
+D.2 Quantization
 Quantization is a crucial technique for deploying efficient TinyML models, offering substantial benefits in terms of reduced latency, memory footprint, and model size.
@@
 Given the negligible performance degradation and the substantial latency, memory, and model size benefits of quantization, we conclude that quantization is a win for person detection.
-C.3 Grayscale
+D.3 Grayscale
 Converting a model’s input image channels to grayscale from RGB is a commonly employed optimization in the TinyML field [2] as it can substantially reduce a model’s memory consumption.
 We observed, however, that the grayscale optimization disproportionately impacts images on the brighter end of the spectrum as illustrated in Figure 8.
-This further demonstrates the importance of fine grained analysis, as some real world deployment environments might be far brighter than the average Wake Vision test sample.
+This further demonstrates the importance of fine-grained analysis, as some real-world deployment environments might be far brighter than the average Wake Vision test sample.
 Figure 8: Impact of grayscale input images on the lighting benchmarks: dark, normal light, and bright. Models that use grayscale input images are more sensitive to bright lighting conditions than RGB.
-Appendix D Inducing Label Errors
+Appendix E Inducing Label Errors
 The goal is to make a single pass through the dataset and flip the labels of a binary classification dataset such that the expected label error rate is $d$.
@@
 4 don’t overfit on training data with higher label errors, as the inconsistent label noise has a regularizing effect.
-Appendix E Fine Grained Benchmark Images
+Appendix F Fine Grained Benchmark Images
 Fig. 9 gives example images for each of the benchmarks in the suite.
@@
 Figure 9: Images from each fine grained benchmark dataset.
-Appendix F Open Image Label Distribution
+Appendix G Open Image Label Distribution
 Wake Vision is derived from Open Images V7, therefore the diversity of subjects in the images should follow the distribution of labels in Open Images.
 The Label distribution of Open images can be found here: https://storage.googleapis.com/openimages/web/factsfigures_v7.html#statistics
-Appendix G Manual Labeling
-The validation and test sets were manually labeled through the crowd-sourced labeling platform Scale Rapid(https://scale.com/.
-Figure 8 shows a screenshot of the labeling menu.
-The labelers were instructed to label an image as a "person" if a person was present anywhere in the image and "no person" if no visible person was present.
-The labelers also indicated if the image contained a depiction of a person, which meant that sample was labeled "no person" in Wake Vision, or a "picture of a picture/reflection", which we used for metadata.
-The cost per image was \$0.10 USD and is set based on Scale’s pricing structure.
-Each image is labeled by 3 different labelers to form a consensus.
-The total cost of labeling was \$7089.8 USD.
-The authors do not know the hourly rate paid to each labeler, but Scale lists \$18/hr on job postings at the time of writing.
-For context on the difficulty of the task, the authors were able to average around 500 images per hour at a reasonable pace when estimating error rates.
-Figure 8: A screenshot of the labeling menu used to manually label the validation and test sets.
 Appendix H Dataset Access and Organization
-Wake Vision is available online at the Harvard Dataverse, via TFDS, or Hugging Face Datasets. More information can be found on the https://wakevision.ai webpage.
-The dataset is organized as a set of compressed tar files containing images, and a series of label CSVs.
+Wake Vision is available online via TFDS and Hugging Face Datasets. We exclude direct links to the dataset during the review period to avoid breaking anonymity.
+The dataset is organized as a set of compressed tar files containing images and a series of label CSVs.
 The label CSVs are organized such that the file name of the image is the identifying index.
@@
 Person Labels.
-The most straightforward Open Images label classes to label as person in Wake Vision are the "person" label and its subcategories (listed in appendix L).
+The most straightforward Open Images label classes to label as a person in Wake Vision are the "person" label and its subcategories (Configurable, but defaults listed in Appendix L).
 All of these are relabelled as persons in Wake Vision.
 These labels are present as both image-level labels and bounding box labels.
-We furthermore inspect the image level label classes for synonyms and umbrella terms for all the person related labels.
+We furthermore inspect the image-level label classes for synonyms and umbrella terms for all the person related labels.
 This search resulted in an additional six person related label classes.
 These person related label classes will only be used in the image level label configuration.
+Lastly, the image level person labels in Open Images have an associated confidence score.
+This score ranges from 0-10.
+Labels that have been verified by humans to be absent from an image have a score of 0, while labels that have been verified present by humans have a score of 10.
+Purely machine-generated labels have a fractional confidence score that is generally $>=5$ [21, 19].
+By default, we only respect labels that have a minimum confidence of 7.
+Labels below this threshold are ignored.
 Person Body Part Labels. Body parts are more challenging to relabel, as it is dependent on the use case whether a body part should be considered a person.
@@
 Bounding Box Size.
-The VWW dataset only considered a Common Objects in Context (COCO) person to be a person if the bounding box around the person took up at least 5% of the image [5].
-If the person took up less than 5% of the image, the image was excluded from the dataset.
-To make our dataset work as a plug in replacement for VWW, we adopt the same defaults in Wake Vision.
+The VWW dataset only considered a Common Objects in Context (COCO) person to be a person if the bounding box around the person took up at least 0.5% of the image [5].
+If the person took up less than 0.5% of the image, the image was excluded from the dataset.
+For Wake Vision, we observed a decrease in performance at the 0.5% limit.
+We theorize that the cause of this decrease is that a person becomes indistinguishable at such a low percentage of already small images.
+Therefore we adopted a 5% threshold for the Wake Vision dataset instead.
 For different requirements, users can change a configuration parameter in our open-source dataset creation code.
@@
 To allow for both data-centric and model-centric improvements, we provide a standard model for data-centric contributions, and a standard pre-processing pipeline for model-centric contributions.
-Both types of contributions are expected to use accuracy as the primary metric for the overall test and validation set, and F1-score as the primary metric for the fine grained benchmarks introduced in section 4.
+Both types of contributions are expected to use accuracy as the primary metric for the overall test and validation set, and F1-score as the primary metric for the fine grained benchmarks introduced in Sec.
+3.3.
 Therefore, for data-centric contributions we propose to use a Mobilenet v2 model with a width modifier [^5] Also known as alpha.
@@
 As an unfortunate result of this, the dataset is not static over time as image owners can delete their images from the flickr platform.
-As a result we were only able to download a subset of the original Open Images v7 dataset as shown in table 6.
+As a result we were only able to download a subset of the original Open Images v7 dataset as shown in Tab.
+7.
 Table 7: Number of images downloaded from Open Images v7. Download occurred between the 28th of November to the 5th of December
@@
 Appendix L Person Label Classes
-We consider the following Open Images v7 labels to be a person for the Wake Vision dataset:
+In our default configuration, we consider the following Open Images v7 labels to be a person for the Wake Vision dataset:
   • Person
@@
   • Youth (Umbrella term for Girl & Boy - Only in Image Level Label Configuration)
-Images containing the following Open Images v7 labels and no other person related labels are excluded from the Wake Vision dataset:
+The following labels are considered body parts and are labelled according to the body part configuration.:
   • Human eye (Part of Person)
@@
 Appendix M Wake Vision Dataset Size
+See Tab. 8.
 Table 8: Amount of images in the Wake Vision dataset
@@
 Test Dataset 27,881 27,881 53,543
-Appendix N Automatic Label Correction
+Appendix N Manual Label Correction
+We use the Scale Rapid tool from Scale AI to relabel the Wake Vision validation and test sets.
+Figure 10 shows a screenshot of the labeling menu presented to the human labelers.
+The labelers were instructed to label an image as a "person" if a person was present anywhere in the image and "no person" if no visible person was present.
+The labelers also indicated if the image contained a depiction of a person or a "picture of a picture/reflection", which we use for metadata.
+The cost per image was \$0.10 USD and is set based on Scale’s pricing structure.
+Each image is labeled by 3 different labelers to form a consensus.
+The total cost of labeling was \$7089.8 USD.
+The authors do not know the hourly rate paid to each labeler, but Scale lists \$18/hr on job postings at the time of writing.
+For context on the difficulty of the task, the authors were able to average around 500 images per hour at a reasonable pace when estimating error rates.
+Figure 10: A screenshot of the labeling menu used to manually label the validation and test sets.
+Appendix O Automatic Label Correction
 To address the challenge of correcting Wake Vision validation and test labels we initially employed the Confident Learning technique [29] to intelligently identify potential label errors.
@@
 To further correct label errors in the final Wake Vision Validation and Test sets we employed the Scale AI platform to crowd-source manual label corrections.
-These label corrections are described in section 3.2 and deprecate the automatic label corrections described in this section.
+These label corrections are described in Sec.
+3.2 and deprecate the automatic label corrections described in this section.
+Appendix P Author Statement
+The authors of this paper hereby confirm that we bear all responsibility in case of violation of rights, etc, and confirm the data license to be CC BY 4.0 for our labels, and that all images in the dataset have been uploaded to Flickr with a CC BY 2.0 license.
+Note: while the Open Images authors tried to identify images that are licensed under a Creative Commons Attribution license, we make no representations or warranties regarding the license status of each image and users should verify the license for each image themselves.
+Appendix Q Hosting, License and Maintenance Plan
+The dataset is hosted on HuggingFace Datasets.
+Additionally, the dataset is available through TensorFlow datasets and can be regenerated from Open Images using our open-source filtering code.
+The Edge AI Foundation will assume responsibility for hosting and maintaining the dataset long-term.
+Wake Vision’s labels and metadata are licensed under a CC BY 4.0 license.
+Images in the dataset have all been uploaded to Flickr with a CC BY 2.0 license.
+Note: while the Open Images authors tried to identify images that are licensed under a Creative Commons Attribution license, we make no representations or warranties regarding the license status of each image and users should verify the license for each image themselves.
+Appendix R DOI and Croissant URLs
+The DOI and URLs will be provided after the review period to avoid breaking anonymity.
+Appendix S Datasheet for Wake Vision
+This document is based on Datasheets for Datasets by Gebru et al. [12]. Please see the most updated version here.
+{mdframed}
+[linecolor=black]
+MOTIVATION
+For what purpose was the dataset created? Was there a specific task in mind? Was there a specific gap that needed to be filled? Please provide a description.
+Wake Vision was created as a prototype for a new generation of datasets for TinyML.
+Where previous TinyML datasets were small due to limited resources, Wake Vision is a production grade dataset consisting of almost 6 million images and with an extensively cleaned validation and test set.
+Wake Vision specifically targets the binary person classification task.
+Such a large dataset enables TinyML researchers to explore questions previously impossible.
+For example, the size of the Wake Vision dataset enables new research into the tradeoff that often exist between dataset size and sample quality.
+Who created this dataset (e.g., which team, research group) and on behalf of which entity (e.g., company, institution, organization)?
+These details will be provided after the review period to avoid breaking anonymity.
+What support was needed to make this dataset?
+(e.g.who funded the creation of the dataset?
+If there is an associated grant, provide the name of the grantor and the grant name and number, or if it was supported by a company or government agency, give those details.)
+These details will be provided after the review period to avoid breaking anonymity.
+Any other comments?
+No additional comments
+{mdframed}
+[linecolor=black]
+COMPOSITION
+What do the instances that comprise the dataset represent (e.g., documents, photos, people, countries)?
+Are there multiple types of instances (e.g., movies, users, and ratings; people and interactions between them; nodes and edges)?
+Please provide a description.
+The dataset is composed of images, labels, and metadata.
+The images are sourced from Open Images.
+The labels are binary, with ’1’ indicating there is a person somewhere in the image and ’0’ indicating that there is not a person in the image.
+The metadata is used to form the benchmark suite where we can evaluate models in specific challenging settings.
+Each image has a binary variable for each fine-grained benchmark where a ’1’ indicates that the image belongs to the fine-grained benchmark set.
+How many instances are there in total (of each type, if appropriate)?
+There are over 6 million labeled images in the dataset.
+Does the dataset contain all possible instances or is it a sample (not necessarily random) of instances from a larger set?
+If the dataset is a sample, then what is the larger set?
+Is the sample representative of the larger set (e.g., geographic coverage)?
+If so, please describe how this representativeness was validated/verified.
+If it is not representative of the larger set, please describe why not (e.g., to cover a more diverse range of instances, because instances were withheld or unavailable).
+The dataset uses a large subset of the Open Images dataset (6 million out of a possible 9 million images).
+We use a subset in order to have a balanced number of person samples and non-person samples.
+We also filter the dataset for quality.
+The filtering removes subjects that are too far and outside of a center crop.
+Open Images sources images from Flickr that are licensed as CC-BY 2.0.
+This is likely not a fully representative sample of the overall world population and likely skews heavily towards a European or North American population.
+See the Open Images paper [21] for more information.
+What data does each instance consist of? “Raw” data (e.g., unprocessed text or images) or features? In either case, please provide a description.
+The dataset is composed of images, labels, and metadata.
+Is there a label or target associated with each instance? If so, please provide a description.
+Yes, every image has a binary label to indicate if a person is visible in the image.
+Images in the some dataset splits include more metadata labels, e.g., about distance of subject or whether the subject is a depiction.
+Is any information missing from individual instances?
+If so, please provide a description, explaining why this information is missing (e.g., because it was unavailable).
+This does not include intentionally removed information, but might include, e.g., redacted text.
+The training set does not contain the metadata used for the fine grained benchmark suite.
+The dataset contains two training sets, one of higher quality and one larger.
+The higher quality dataset and other sets contain information about depictions and body parts.
+The large training set does not contain this information.
+Are relationships between individual instances made explicit (e.g., users’ movie ratings, social network links)? If so, please describe how these relationships are made explicit.
+No
+Are there recommended data splits (e.g., training, development/validation, testing)? If so, please provide a description of these splits, explaining the rationale behind them.
+Yes there are separate splits for training, validation, and testing.
+These splits are generated from equivalent splits in Open Images.
+The dataset contains two training splits, one focusing on quality of samples and one on number of samples.
+Are there any errors, sources of noise, or redundancies in the dataset? If so, please provide a description.
+Yes there are label errors in the dataset due to the labeling process. We discuss this at length in the paper.
+Is the dataset self-contained, or does it link to or otherwise rely on external resources (e.g., websites, tweets, other datasets)?
+If it links to or relies on external resources, a) are there guarantees that they will exist, and remain constant, over time; b) are there official archival versions of the complete dataset (i.e., including the external resources as they existed at the time the dataset was created); c) are there any restrictions (e.g., licenses, fees) associated with any of the external resources that might apply to a future user?
+Please provide descriptions of all external resources and any restrictions associated with them, as well as links or other access points, as appropriate.
+The dataset is self-contained.
+We re-host the images, labels, and metadata.
+The images may be subject to different licenses than our labels and metadata.
+All images have originally been hosted on Flickr under a permissive CC BY license.
+Does the dataset contain data that might be considered confidential (e.g., data that is protected by legal privilege or by doctor-patient confidentiality, data that includes the content of individuals’ non-public communications)?
+If so, please provide a description.
+No, the images were posted online with the CC-BY 2.0 license.
+It is possible an image was posted without the individual’s consent.
+While we tried to identify images that are licensed under a Creative Commons Attribution license, we make no representations or warranties regarding the license status of each image, and a user should verify the license for each image themselves.
+Does the dataset contain data that, if viewed directly, might be offensive, insulting, threatening, or might otherwise cause anxiety? If so, please describe why.
+No, not to our knowledge. Due to the size of the dataset we have not had the resources to manually verify this.
+Does the dataset relate to people? If not, you may skip the remaining questions in this section.
+Yes, the dataset contains images of people.
+Does the dataset identify any subpopulations (e.g., by age, gender)?
+If so, please describe how these subpopulations are identified and provide a description of their respective distributions within the dataset.
+Yes the benchmark suite contains subsets that focus on perceived age and gender.
+These labels have been sourced from the More Inclusive Annotations of People [34].
+The purpose of these benchmarks are to test the fairness of a trained model across these subpopulation and we do not condone using these labels to train a model.
+The size of these sets is listed in Table 5.
+Is it possible to identify individuals (i.e., one or more natural persons), either directly or indirectly (i.e., in combination with other data) from the dataset? If so, please describe how.
+It could be possible to identify a person by their face or other visible characteristics in the image. We do not share any other personal information besides the images.
+Does the dataset contain data that might be considered sensitive in any way (e.g., data that reveals racial or ethnic origins, sexual orientations, religious beliefs, political opinions or union memberships, or locations; financial or health data; biometric or genetic data; forms of government identification, such as social security numbers; criminal history)?
+If so, please provide a description.
+No, not to our knowledge.
+Some personal or sensitive information may be visible in the images.
+However, the images were shared publicly under a CC-BY 2.0 license.
+We do not include any personal information beyond the images.
+Any other comments?
+The images used were published publicly under a CC-BY 2.0 license and are already used in the Open Images [19] dataset and its derivatives.
+The authors of Open images tried to identify images that are licensed under a Creative Commons Attribution license but make no representations or warranties regarding the license status of each image and a user should verify the license for each image themselves.
+{mdframed}
+[linecolor=black]
+COLLECTION
+How was the data associated with each instance acquired?
+Was the data directly observable (e.g., raw text, movie ratings), reported by subjects (e.g., survey responses), or indirectly inferred/derived from other data (e.g., part-of-speech tags, model-based guesses for age or language)?
+If data was reported by subjects or indirectly inferred/derived from other data, was the data validated/verified?
+If so, please describe how.
+The dataset is derived from Open Images [19], which in turn acquired its images from Flickr. Section 3.1 describes the process.
+Over what timeframe was the data collected?
+Does this timeframe match the creation timeframe of the data associated with the instances (e.g., recent crawl of old news articles)?
+If not, please describe the timeframe in which the data associated with the instances was created.
+Finally, list when the dataset was first published.
+The original data collection was conducted in Noember of 2015. The Open Images paper [21] details the image acquisition process.
+What mechanisms or procedures were used to collect the data (e.g., hardware apparatus or sensor, manual human curation, software program, software API)?
+How were these mechanisms or procedures validated?
+The images were sources from Flickr in an automated pipeline.
+The original labeling pipeline was a combination of automated labeling and manual verification and labeling.
+The Open Images paper details this process [21].
+We detail our process for relabeling and filtering the dataset in Section 3.1.
+What was the resource cost of collecting the data?
+(e.g.
+what were the required computational resources, and the associated financial costs, and energy consumption - estimate the carbon footprint.
+See Strubell et al.[35] for approaches in this area.)
+We used cloud TPU credits provided by the Google Cloud TRC program.
+The primary source of required resources was model training to evaluate the dataset and the storage and bandwidth required to process and upload the dataset to hosting locations.
+The total size of the dataset is approximately 2 TB.
+To the authors knowledge, there is no public information about the power draw of TPUs.
+This matches the conclusions drawn by Strubell et al.[35].
+This makes it impossible to disclose the energy consumption and estimated carbon footprint of the project.
+We used paid crowdsource labelers to relabel the validation and test sets. The total cost of labeling was approximately \$7,000.
+If the dataset is a sample from a larger set, what was the sampling strategy (e.g., deterministic, probabilistic with specific sampling probabilities)?
+The dataset is a large subset of the total Open Images dataset. The subset is generated by a deterministic filtering process described in Section 3.1.
+Who was involved in the data collection process (e.g., students, crowdworkers, contractors) and how were they compensated (e.g., how much were crowdworkers paid)?
+We used paid crowdsource labelers to relabel the validation and test sets.
+The total cost of labeling was approximately \$7,000.
+We used the Scale Rapid labeling service.
+A total of 70,898 images were labeled at a cost of \$0.1 per image and \$0.033 per individual label.
+Appendix 3.2 gives additional detail.
+Were any ethical review processes conducted (e.g., by an institutional review board)?
+If so, please provide a description of these review processes, including the outcomes, as well as a link or other access point to any supporting documentation.
+No
+Does the dataset relate to people? If not, you may skip the remainder of the questions in this section.
+Yes the dataset contains images of people.
+Did you collect the data from the individuals in question directly, or obtain it via third parties or other sources (e.g., websites)?
+We obtained the images via the Open Images dataset [21] which originally sourced its images from Flickr [10].
+Were the individuals in question notified about the data collection?
+If so, please describe (or show with screenshots or other information) how notice was provided, and provide a link or other access point to, or otherwise reproduce, the exact language of the notification itself.
+No the individuals were not notified to our knowledge. The images were posted publicly under the CC-BY 2.0 license.
+Did the individuals in question consent to the collection and use of their data?
+If so, please describe (or show with screenshots or other information) how consent was requested and provided, and provide a link or other access point to, or otherwise reproduce, the exact language to which the individuals consented.
+Yes, the images were posted publicly under the CC-BY 2.0 license. It is possible that the images were uploaded by a third party without the consent of the individual in the image.
+If consent was obtained, were the consenting individuals provided with a mechanism to revoke their consent in the future or for certain uses?
+If so, please provide a description, as well as a link or other access point to the mechanism (if appropriate)
+No we do not have a mechanism for an individual the revoke their consent.
+As with Open Images, we make no representations or warranties regarding the license status of each image, and a user should verify the license for each image themselves via the Flickr link in the metadata.
+Note that it is impossible to revoke a CC BY license after publication [23].
+Has an analysis of the potential impact of the dataset and its use on data subjects (e.g., a data protection impact analysis)been conducted?
+If so, please provide a description of this analysis, including the outcomes, as well as a link or other access point to any supporting documentation.
+No formal analysis has been conducted. We describe the ethical considerations of the dataset in Section 5.
+Any other comments?
+No additional comments.
+{mdframed}
+[linecolor=black]
+PREPROCESSING / CLEANING / LABELING
+Was any preprocessing/cleaning/labeling of the data done(e.g.,discretization or bucketing, tokenization, part-of-speech tagging, SIFT feature extraction, removal of instances, processing of missing values)?
+If so, please provide a description.
+If not, you may skip the remainder of the questions in this section.
+Yes we filtered the data as shown in Figure 6.
+Was the “raw” data saved in addition to the preprocessed/cleaned/labeled data (e.g., to support unanticipated future uses)? If so, please provide a link or other access point to the “raw” data.
+The "raw" data is accessible in the Open Images dataset.
+Is the software used to preprocess/clean/label the instances available? If so, please provide a link or other access point.
+Yes, we open-source our filtering code. https://anonymous.4open.science/r/Wake_Vision-DD19/README.md
+Any other comments?
+No additional comments
+{mdframed}
+[linecolor=black]
+USES
+Has the dataset been used for any tasks already? If so, please provide a description.
+The dataset has been used for the experiments reported in the paper.
+Is there a repository that links to any or all papers or systems that use the dataset? If so, please provide a link or other access point.
+The usage of the dataset can be tracked via the citations of this paper. We plan to host competitions using the dataset and will track participants.
+What (other) tasks could the dataset be used for?
+The dataset can be used as a benchmark for TinyML models and as the basis for data-centric AI research given the prevalence of label errors in the training set and the cleaned nature of the validation and test sets.
+Is there anything about the composition of the dataset or the way it was collected and preprocessed/cleaned/labeled that might impact future uses?
+For example, is there anything that a future user might need to know to avoid uses that could result in unfair treatment of individuals or groups (e.g., stereotyping, quality of service issues) or other undesirable harms (e.g., financial harms, legal risks) If so, please provide a description.
+Is there anything a future user could do to mitigate these undesirable harms?
+Yes, the estimated error rates are reported in Table 2.
+We also discuss the high likelihood of false negatives in Wake Vision’s training sets in Section 3.2.
+These errors are due to the labeling and filtering process.
+We mitigate the impact of these errors by manually labeling the validation and test sets as well as study their impact for TinyML models in Section 4.4.
+Are there tasks for which the dataset should not be used? If so, please provide a description.
+The perceived age and gender benchmarks should not be used to train a gender or age classifier.
+Any other comments?
+No additional comments
+{mdframed}
+[linecolor=black]
+DISTRIBUTION
+Will the dataset be distributed to third parties outside of the entity (e.g., company, institution, organization) on behalf of which the dataset was created? If so, please provide a description.
+The dataset is hosted on Harvard Dataverse and HuggingFace Datasets.
+Additionally, the dataset is available through TensorFlow datasets and can be regenerated from Open Images with our open source filtering code.
+We plan for the TinyML foundation (https://www.tinyml.org/) to assume responsibility for hosting and maintaining the dataset long-term.
+How will the dataset will be distributed (e.g., tarball on website, API, GitHub)? Does the dataset have a digital object identifier (DOI)?
+Yes, the dataset is hosted on Harvard Dataverse and HuggingFace Datasets.
+Additionally, the dataset is available through TensorFlow datasets and can be regenerated from Open Images with our open source filtering code.
+When will the dataset be distributed?
+The Dataset is publicly accessible at the time of writing.
+Will the dataset be distributed under a copyright or other intellectual property (IP) license, and/or under applicable terms of use (ToU)?
+If so, please describe this license and/or ToU, and provide a link or other access point to, or otherwise reproduce, any relevant licensing terms or ToU, as well as any fees associated with these restrictions.
+The dataset is published under the CC-BY 4.0 license.
+Have any third parties imposed IP-based or other restrictions on the data associated with the instances?
+If so, please describe these restrictions, and provide a link or other access point to, or otherwise reproduce, any relevant licensing terms, as well as any fees associated with these restrictions.
+No
+Do any export controls or other regulatory restrictions apply to the dataset or to individual instances?
+If so, please describe these restrictions, and provide a link or other access point to, or otherwise reproduce, any supporting documentation.
+No
+Any other comments?
+No additional comments.
+{mdframed}
+[linecolor=black]
+MAINTENANCE
+Who is supporting/hosting/maintaining the dataset?
+The dataset is hosted on HuggingFace Datasets.
+Additionally, the dataset is available through TensorFlow datasets and can be regenerated from Open Images with our open-source filtering code.
+The Edge AI Foundation will assume responsibility for hosting and maintaining the dataset long-term.
+How can the owner/curator/manager of the dataset be contacted (e.g., email address)?
+Removed during review to maintain anonymity.
+Is there an erratum? If so, please provide a link or other access point.
+No
+Will the dataset be updated (e.g., to correct labeling errors, add new instances, delete instances)?
+If so, please describe how often, by whom, and how updates will be communicated to users (e.g., mailing list, GitHub)?
+The dataset may be updated if any problematic instances are discovered or the label errors are corrected.
+In this case, we will issue a new version of the dataset to ensure past results are contextualized under the version of the dataset that was used.
+If the dataset relates to people, are there applicable limits on the retention of the data associated with the instances (e.g., were individuals in question told that their data would be retained for a fixed period of time and then deleted)?
+If so, please describe these limits and explain how they will be enforced.
+There is no limit.
+Will older versions of the dataset continue to be supported/hosted/maintained? If so, please describe how. If not, please describe how its obsolescence will be communicated to users.
+Yes, we will continue to host older versions of the dataset. Details will be provided after review to preserve anonymity.
+If others want to extend/augment/build on/contribute to the dataset, is there a mechanism for them to do so?
+If so, please provide a description.
+Will these contributions be validated/verified?
+If so, please describe how.
+If not, why not?
+Is there a process for communicating/distributing these contributions to other users?
+If so, please provide a description.
+Yes, the dataset is licensed permissively, and the code used to create Wake Vision is open source.
+Any other comments?
+No additional comments

### v5

@@
 Abstract
-Tiny machine learning (TinyML) for low-power devices lacks robust datasets for development.
-We present Wake Vision, a large-scale dataset for person detection that contains over 6 million quality-filtered images.
-We provide two variants: Wake Vision (Large) and Wake Vision (Quality), leveraging the large variant for pretraining and knowledge distillation, while the higher-quality labels drive final model performance.
-The manually labeled validation and test sets reduce error rates from 7.8% to 2.2% compared to previous standards.
-In addition, we introduce five detailed benchmark sets to evaluate model performance in real-world scenarios, including varying lighting, camera distances, and demographic characteristics.
-Training with Wake Vision improves accuracy by 1.93% over existing datasets, demonstrating the importance of dataset quality for low-capacity models and dataset size for high-capacity models.
-The dataset, benchmarks, code, and models are available under the CC-BY 4.0 license, maintained by the Edge AI Foundation.
+Tiny machine learning (TinyML) for low-power devices lacks systematic methodologies for creating large, high-quality datasets suitable for production-grade systems.
+We present a novel automated pipeline for generating binary classification datasets that addresses this critical gap through several algorithmic innovations: intelligent multi-source label fusion, confidence-aware filtering, automated label correction, and systematic fine-grained benchmark generation.
+Crucially, automation is not merely convenient but necessary to cope with TinyML’s diverse applications.
+TinyML requires bespoke datasets tailored to specific deployment constraints and use cases, making manual approaches prohibitively expensive and impractical for widespread adoption.
+Using our pipeline, we create Wake Vision, a large-scale binary classification dataset of almost 6 million images that demonstrates our methodology through person detection—the canonical vision task for TinyML.
+Wake Vision achieves up to a 6.6% accuracy improvement over existing datasets via a carefully designed two-stage training strategy and provides 100× more images.
+We demonstrate our broad applicability for automated large-scale TinyML dataset generation across two additional target categories, and show our label error rates are substantially lower than prior work.
+Our comprehensive fine-grained benchmark suite evaluates model robustness across five critical dimensions, revealing failure modes masked by aggregate metrics.
+To ensure continuous improvement, we establish ongoing community engagement through competitions hosted by the Edge AI Foundation.
+All datasets, benchmarks, and code are available under CC-BY 4.0 license, providing a systematic foundation for advancing TinyML research.
+**footnotetext: These authors contributed equally to this work.
 1 Introduction
 Ultra-low power machine learning (TinyML) has emerged as an important technology to enable efficient ML deployments by co-locating models with sensors on microcontrollers (MCUs) [3, 47, 1, 8].
-Although this approach dramatically reduces energy consumption and deployment costs, it imposes severe constraints.
-For instance, the models must often be within hundreds of kilobytes of memory in size, nearly four orders of magnitude less than mobile phones[2].
-TinyML requires large-scale datasets to facilitate high-quality research.
-However, the extreme constraints imposed by the hardware necessitate that TinyML models be compact and efficient, rendering them unsuitable for traditional, complex ML tasks.
-As a result, conventional ML datasets like ImageNet [7], designed for more resource-intensive models, are often ill-suited for TinyML research endeavors.
-While existing TinyML datasets [5, 37] have been valuable, they tend to be limited in scale, making it challenging to train production-grade TinyML models.
-This creates a fundamental challenge where we can not develop robust TinyML models due to existing datasets being either too complex for tiny architectures or too limited in scale and diversity.
-Table 1: A comparison of person detection and image classification datasets.
+Although this approach dramatically reduces energy consumption and deployment costs, it imposes severe constraints that fundamentally differ from standard machine learning scenarios.
+For instance, the models must often fit within hundreds of kilobytes of memory, nearly four orders of magnitude less than common model sizes on smartphones[2].
+These extreme hardware-imposed constraints necessitate compact and efficient TinyML models, rendering traditional Machine Learning (ML) benchmarks and datasets unsuitable.
+For example, ImageNet [10] models must output a thousand classes, which is infeasible on MCUs and even smartphone models struggle to support [7].
+Existing TinyML datasets such as Visual Wake Words (VWW) [7] and Google Speech Commands [46] target simpler applications, requiring support for binary classification and tens of classes, respectively.
+While these datasets are valuable, their limited scale and label error rates make them unsuitable for pushing the frontier of TinyML research and training production-grade TinyML models.
+This creates a challenge where the development of robust TinyML systems is hindered by existing datasets being either too complex for tiny architectures or too limited in scale and diversity.
+However, this challenge extends beyond simply collecting more data—it requires systematic methodologies for rapidly creating, curating, and evaluating datasets that meet TinyML’s unique constraints at scale.
+Table 1: A comparison of Wake Vision for person detection and image classification datasets.
 Total # of Person Images Fine-Grained Suitable for
 Dataset Images Train Validation Test Filtering TinyML
-Our work – Wake Vision (Quality) 1,322,574 624,115 9,291 27,881
-Our work – Wake Vision (Large) 5,760,428 2,880,214 -- --
+Ours – WV (Quality) 1,322,574 624,115 9,291 27,881
+Ours – WV (Large) 5,760,428 2,880,214 -- --
 Visual Wake Words [7] $123,287$ $36,000$ $3,926$ $19,107$
@@
 PASCAL VOC 2012 [15] $11,530$ $1,994$ $2,093$-
-To address this critical gap, we present Wake Vision, a large-scale dataset specifically designed for person detection, the quintessential vision use case for TinyML [1].
-Person detection is a binary image classification task that, while simple, enables crucial applications ranging from occupancy detection [31] and smart HVAC/lighting [39] to acting as an always-on ’wake model’ in larger ML systems [15].
-Wake Vision advances the state-of-the-art by providing 6M images, approximately 100x more than the current leading dataset Visual Wake Words (VWW) [5], while achieving a 1.93% accuracy improvement.
-The dataset is derived from Open Images v7 [21, 19] and is released under a permissive CC-BY 4.0 license to facilitate broad adoption.
-A key contribution of our work is the first systematic investigation of how dataset quality versus quantity impacts TinyML model performance.
-This question is particularly relevant for resource-constrained models, as prior work suggests that large, overparameterized models can effectively identify and adapt to training set errors [40, 9].
-TinyML models, with their limited capacity, may exhibit fundamentally different behavior.
-To explore this hypothesis, we introduce two complementary training sets within the Wake Vision visual corpus: Wake Vision (Large), emphasizing scale, and Wake Vision (Quality), prioritizing label accuracy.
-Our experiments reveal that while Wake Vision (Quality) outperforms its larger counterpart in naive training scenarios, optimal performance is achieved through a training strategy that leverages both sets — using Wake Vision (Large) for pre-training followed by using Wake Vision (Quality) fine-tuning.
-To further explore the dataset quality and quantity trade-off for tinyML, we train a range of ResNet [14] style models across Wake Vision datasets of varying error rates.
-Our experiments collectively reveal that small models are more sensitive to higher error rates than large models and that large models can benefit from larger datasets despite a higher error rate.
-These findings provide insight into the unique challenges of training highly constrained models.
-Beyond raw performance metrics, real-world deployment of TinyML systems often demands robustness across diverse and challenging use case conditions.
-Internet-sourced datasets often underrepresent challenging scenarios, such as low-lighting conditions, leading to overly optimistic performance estimates based on standard test set metrics.
-To enable a rigorous evaluation of real-world capabilities, we additionally develop a comprehensive suite of five fine-grained benchmarks based on our Wake Vision dataset.
-Our benchmarks systematically evaluate performance across multiple dimensions: distance (detection at varying ranges), lighting (performance in well-lit and poorly-lit environments), depictions (handling of drawings and sculptures), perceived gender (investigating potential gender biases), and perceived age (examining effectiveness across age groups).
-This benchmarking framework addresses a gap in current TinyML assessment methods and enables researchers to identify potential biases, limitations, and areas for improvement that might be masked by aggregate metrics alone.
-We find that Wake Vision models outperform VWW on 13 of the 16 fine-grained sets, demonstrating robust performance further illustrating the advantages of a large diverse training dataset.
-In summary, we present the following contributions:
-  • A production grade computer vision dataset for TinyML containing over 6M images (100$\times$ larger than state-of-the-art) for person detection, with comprehensive filtering, permissive licensing, and accessibility through standard ML frameworks.
-  • A cleaned test set with 2.2% label error rate, enabling more reliable evaluation of TinyML models.
-  • A benchmark suite for systematically evaluating model robustness across five critical dimensions (perceived gender, age, distance, lighting, depictions), exposing potential failure modes in real-world deployments.
-  • Evidence that TinyML models are more sensitive to dataset quality than larger models, providing guidance for dataset curation in resource-constrained settings.
-  • A complete, maintained dataset infrastructure including benchmarks, evaluation code, and reproducible filtering pipeline to support ongoing TinyML research.
+To address this gap, we present an automated dataset generation pipeline for creating and benchmarking large-scale binary classification datasets suitable for TinyML that incorporates several algorithmic innovations: intelligent fusion of multiple label sources, confidence-aware filtering strategies, automated label correction techniques, and systematic generation of fine-grained benchmark suites.
+TinyML dataset automation is not merely a convenience but a necessity—manual labeling approaches are prohibitively expensive (our manual validation of 70K images cost \$7,000, which would extrapolate to \$600,000 for a 6M image dataset) and suffer from inconsistency across labelers.
+Our dataset generation pipeline derives datasets from Open Images v7 [29, 27], which enables created datasets to be licensed under a permissive CC-BY 4.0 license to facilitate broad adoption.
+We demonstrate the effectiveness of our pipeline by creating Wake Vision, a large-scale open-source binary classification dataset that uses person detection to validate our methodology.
+Person detection serves as the quintessential vision use case for TinyML [1] and is an ideal testbed because it enables several important applications ranging from occupancy detection [39] and smart HVAC/lighting [48] to acting as an always-on ‘wake model’ in larger ML systems [23].
+As shown in Table 1, Wake Vision advances the state-of-the-art TinyML datasets by providing almost 6M images, close to 100$\times$ more than the current leading dataset VWW [7], while achieving up to a 6.6% accuracy improvement.
+We use Wake Vision to derive insights about TinyML-specific design principles: small models benefit more from smaller, high-quality training sets compared to noisier, larger sets, but if designed carefully, two-stage training (pre-training on Wake Vision Large followed by fine-tuning on Wake Vision Quality) can leverage both scale and quality to achieve optimal performance (Section 4.1).
+Furthermore, to enable rigorous evaluation of real-world capabilities, we develop a comprehensive suite of five fine-grained benchmarks that systematically evaluate performance across critical dimensions including perceived gender and age, distance, lighting, and depictions (Section 3.3).
+This benchmarking framework addresses a gap in current TinyML assessment methods, revealing that Wake Vision trained models outperform VWW trained models on 13 of 16 fine-grained sets, demonstrating robust performance beyond aggregate metrics alone.
+Beyond person detection, we show that our dataset generation pipeline generalizes to all 9.6K trainable classes provided by Open Images v7, enabling the systematic creation of more production-grade datasets across diverse TinyML applications (Section 3.5).
+We provide evidence with two additional domains for TinyML deployments: bird detection for smart nests [9] and automated feeders [20] (12$\times$ more images with label error rates improving from 6.4% to 4.8%), and car detection for parking monitoring [6] and blind-spot systems [43] (27$\times$ more images with error rates improving from 6.6% to 0.6%).
+Additionally, we host public challenges for long-term, community-driven enhancements to Wake Vision.
+Our dataset, benchmarks, code, and models, along with past and upcoming challenges, can be found at https://wakevision.ai.
 (a) Older Person
@@
 (e) Depicted Person
-Figure 1: Sample images from the Wake Vision fine-grained benchmark suite.
+Figure 1: Sample images from the Wake Vision fine-grained benchmark suite (for person detection).
 2 Background and Related Work
-Person Detection and Its Role in TinyML
-Person detection is a fundamental computer vision task whose job is to determine the presence or absence of humans in visual data.
-Unlike more complex tasks such as pose estimation or instance segmentation, person detection focuses on the binary classification problem.
-This seemingly simple task has emerged as a canonical use case for TinyML systems [3, 1], driven by both theoretical foundations and practical considerations.
+Generating Binary Classification Datasets for TinyML
+Automatically generated datasets can significantly lower costs associated with deploying TinyML systems.
+Prior to our work, the primary way to generate binary classification datasets for TinyML applications was through the code used to create VWW [7], which can be applied to other MS-COCO [33] classes.
+While we offer the same capability for Open Images, Wake Vision goes beyond dataset generation, offering customizable curation properties comprising of label correction, fine-grained benchmark suite generation, and community-driven enhancements through competitions, addressing the whole dataset life-cycle.
+Person Detection in TinyML Applications
+Person detection has emerged as the canonical computer vision use case for TinyML systems [3, 1].
 In the context of resource-constrained computing, TinyML systems demand tasks that carefully balance computational feasibility with real-world utility.
@@
 It is sufficiently lightweight to be implemented on embedded microcontrollers while maintaining enough discriminative power for practical applications.
-The significance of person detection extends beyond its computational characteristics, serving as a critical selective activation mechanism in resource-constrained systems.
-These lightweight models efficiently manage the activation of power-intensive sensors and processors by determining human presence.
-This architectural pattern has demonstrated particular value in diverse applications, including occupancy detection [31], smart HVAC systems [39], and as an always-on "wake model" in larger ML pipelines [15].
-Moreover, this approach enables substantial energy savings while preserving privacy through on-device processing.
+In particular, person detection serves as a critical activation mechanism in resource-constrained systems.
+These lightweight models support always-on deployment on low-powered MCUs to selectively “wake” power-intensive sensors, processors, and larger, more capable ML models upon detecting human presence [23].
+This architectural pattern enables substantial energy savings while preserving privacy through on-device processing, and supports diverse applications ranging from occupancy detection [39], smart HVAC systems [48], or poacher detection [11].
 Existing Person Detection Datasets
 In the TinyML computer vision domain, the VWW dataset [7, 8, 32, 2, 1] has established itself as the de facto standard.
-It remains the only open-source dataset specifically designed for person detection with direct commercial applicability and potential for real-world impact.
+Before Wake Vision, it was the only open-source dataset specifically designed for person detection with direct commercial applicability.
 However, VWW faces significant limitations: its small size and indirect access requirements (requiring regeneration from MS-COCO [33]) restrict its accessibility and utility.
@@
 Furthermore, it distinguishes itself as the only person detection dataset offering a fine-grained benchmark suite and official distribution through popular services like TensorFlow Datasets (TFDS) and Hugging Face Datasets.
-3 Wake Vision: Dataset and Benchmark Suite
-We introduce Wake Vision, a new dataset for TinyML person detection.
+3 Dataset and Benchmark Generation
+With Wake Vision, we introduce a new dataset for TinyML person detection which is two orders of magnitude larger than prior datasets.
 The dataset’s size, quality, and detailed metadata enable new avenues of TinyML research.
-Wake Vision is almost two orders of magnitude larger than prior comparable datasets and can be adapted based on the target use case.
-Fig.
-1 illustrates examples of images in the dataset.
-In this section, we discuss our methodologies for label generation, data filtering, and error correction, as well as discussing the usability of the data set and comparing it with VWW Dataset.
+Figure 1 illustrates example images in our dataset.
+This section presents our methodologies for dataset generation and curation, with a focus on person detection, including label generation, data filtering, error correction, and community-driven enhancements, as well as discussing the usability and generalization capabilities of our dataset pipeline in comparison with VWW.
 3.1 Label Generation
 A large person detection dataset is indispensable for TinyML research.
-It can, however, be challenging to obtain.
-Manually labeling millions of images would be prohibitively expensive for a nascent field like TinyML.
+However, manually labeling millions of images would be prohibitively expensive for a nascent field like TinyML.
 Therefore, we bootstrap Wake Vision using existing large-scale data efforts.
 The base label in Wake Vision is a binary person/non-person label.
-We derive these labels from Open Images [21, 19], which contain both image-level and bounding box labels of various objects.
-Image-level labels describe which objects are present in an image but not localized to a specific point in the image.
-Bounding box labels provide a set of four coordinates forming a box around the object.
+We derive these labels from Open Images [29, 27], which contain both image-level and bounding box labels of 9.6K trainable classes and 600 objects, respectively.
+Image-level labels are unlocalized, describing objects present in an image while bounding box labels localize an object by four coordinates.
 The bounding box label classes are hierarchically structured so that one class can be a subcategory or a part of another class.
 For example a “Woman” is a subcategory of a “Person,” and a “Human Hand” is a part of “Person.” A comprehensive list of Open Images labels related to Wake Vision can be found in Appendix L.
-Users can adapt Wake Vision to meet the needs of specific use cases by adding several configurable options to the open-source dataset creation script.
-For example, a configuration option exists to change whether depictions are labeled as non-persons or excluded from the dataset.
+Users can adapt Wake Vision to meet the needs of specific use cases using the many configuration options available in our binary classification dataset generation pipeline.
+For example, a configuration option exists to change whether artistic depictions of humans are labeled as non-persons or excluded from the dataset.
 We refer to Appendix I for details about these configuration options.
-We observe that the bounding box labels are generally more accurate and provide more information for data filtering but are less numerous.
+We observe that bounding box labels are generally more accurate and provide more information for data filtering while being less numerous.
 The Open Images training set has ~9 million images with image-level person labels, but only ~1.7 million images with a person bounding box.
 Consequently, this presents a trade-off between more data (image-level labels) or higher-quality labels (bounding box labels).
-Because of this trade-off, we provide two Wake Vision training sets: Wake Vision (Large) and Wake Vision (Quality).
-One larger set, labeled via Open Images image-level labels, and one smaller set, labeled via Open Images bounding box labels (Table 1).
-The Open Images validation and test sets are fully labeled with bounding box labels.
-Therefore, both the Wake Vision validation and test sets are created using the higher-quality bounding box labels.
-Large Training Set from Image-Level Labels
-Image-level labels in Open Images contain little additional information than the raw labels.
-The only additional relevant information is a confidence property, which represents how certain it is that a label is correct.
-The confidence ranges from 0 to 1 for labels generated by machines and is either 0 or 1 for labels negatively or positively verified by humans, respectively.
-See Sec.
-3.2 for more information.
-We use this property to exclude images from Wake Vision if they contain only low-confidence person labels.
-Quality Training Set From Bounding Box Labels
-Bounding Box labels in Open Images, in contrast to image-level labels, are all verified and localized by humans.
-This ensures that bounding box labels are less likely to be false positives.
-Bounding box labels can be used to calculate an approximation of the area taken up by a person.
-We use this as a proxy for the distance of a person in the image, and by default exclude images where all person labels are far away, i.e., take up less than 5% of the image.
-An example of an image with a faraway person can be found in Fig.
-2(a).
-Finally, bounding box attributes include a depiction flag, which represents whether a label is a depiction.
-See an example of an image containing a person with the depiction flag set in Fig.
-2(b).
-By default, we use this flag to make sure that we do not consider a depiction a person.
-See Appendix C for an illustration of the bounding box filtering process.
+In response, we provide two Wake Vision training sets: Wake Vision (Large), labeled via Open Images image-level labels, and Wake Vision (Quality), a smaller set labeled via Open Images bounding boxes (Table 1).
+Since the Open Images validation and test sets are fully labeled with higher-quality bounding box labels, we derive the Wake Vision validation and test sets from the bounding box labels in these Open Images splits.
+Wake Vision (Large) Training Set
+Image-level labels in Open Images include a confidence property, which represents how certain it is that a label is correct.
+This confidence ranges from 0 to 1 for machine-generated labels and is strictly either 0 or 1 for human-verified labels.
+See Section 3.2 for more information.
+We exclude low-confidence machine-labeled person images from Wake Vision.
+Wake Vision (Quality) Training Set
+Bounding box labels in Open Images, in contrast to image-level labels, are all verified and localized by humans, minimizing false positive labels.
+Bounding box labels can be used to calculate an approximation of the image area taken up by a person.
+We use area as a proxy for the distance of a person to the camera, and by default exclude far away persons, i.e., take up less than 5% of the image (see  Figure 2(a) for an example).
+Finally, bounding box attributes include an artistic depiction flag (an example is provided in  Figure 2(b)).
+By default, Wake Vision does not consider a depiction a person.
+See Appendix D for an illustration of the bounding box filtering process.
 (a) An example of a person far away
@@
 This is especially prevalent in large datasets that use complex labeling pipelines to scale while managing costs.
-Recognizing the importance of label quality, we take measures to estimate and improve the accuracy of its labels, aligning with best practices in the field of computer vision.
-Wake Vision’s labels are derived from the Open Images labels; therefore, the errors that occur in the Open Images labeling pipeline are inherited by Wake Vision.
-All labels in the Open Images dataset originate from machine-generated candidate labels.
+Recognizing the importance of label quality, we take measures to estimate and improve the accuracy of Wake Vision’s labels.
+As Wake Vision’s labels are derived from Open Images v7; errors in these labels are inherited by Wake Vision.
+All image-level labels in the Open Images dataset originate from machine-generated candidate labels.
 Many of these labels are later verified by human annotators, and a subset of them are given bounding boxes and additional annotations.
@@
 Consequently, the Open Images dataset may contain numerous false negative labels, referring to images where an object is present but lacks the corresponding label.
-Therefore, additional measures are necessary to identify and correct such labeling omissions to ensure the dataset’s integrity.
+Therefore, additional measures are necessary to identify and correct such labeling omissions.
 Given Wake Vision’s size, manually correcting label errors across the entire dataset becomes an arduous undertaking.
@@
 We report our estimated label error rate after label corrections versus VWW in Table 2.
-These estimates are all based on a random subset of 500 samples from each dataset, which the authors manually relabelled according to whether we saw a non-depiction person or body part in the image.
+These estimates are all based on a random subset of 500 samples from each dataset, manually evaluated by the authors.
 The Wake Vision validation and test set is considerably lower than that of the VWW dataset top-level error rate.
 Table 2: Label error rate of VWW and Wake Vision after label correction.
-The Wake Vision error rate is broken down by set.
 The Wake Vision Train (Quality) and Train (Large) have different error rates due to the different label sources.
@@
 By providing these targeted evaluation sets, we enable a more nuanced analysis of model behavior beyond aggregate accuracy metrics.
-The suite consists of five fine-grained benchmark sets.
-Each of the five fine-grained benchmarks has been chosen based on a combination of its relevance to TinyML use cases and the availability of requisite metadata to generate the sets.
+The suite consists of five fine-grained benchmark sets, three of which are applicable to any dataset created by our dataset generation pipeline, while two are specific to Wake Vision.
+The three generally applicable fine-grained benchmarks include Distance, Lighting, and Depictions, whereas the Wake Vision-specific criteria include Perceived Gender and Perceived Age.
+Appendix F details all the adopted criteria.
+Each of the five fine-grained benchmarks has been picked based on a combination of its relevance to TinyML use cases and the availability of requisite metadata to generate the sets.
 Each benchmark set is a subset of the respective validation or test set filtered based on the criteria under test.
 The benchmark determines whether a model is sufficiently accurate in the planned deployment setting.
-For example, a model designer may make different design choices for a use case where the subject is close to the camera and well lit compared to the inverse setting.
-Appendix D contains a case study that shows how fine-grained evaluation sets can be used to design robust models.
+For example, a model designer may make different design choices for a use case where the subject is close to the camera and well-lit compared to the inverse setting.
+Appendix E contains a case study that shows how fine-grained evaluation sets can be used to design robust models.
+Table 3: Accuracy on the Wake Vision and VWW test sets by models trained on the VWW, Wake Vision (Quality) and Wake Vision (Combined) training sets.
+Train
+VWW Wake Vision (Quality) Wake Vision (Combined)
+Test VWW 88.33$\pm$0.2988.59$\pm$0.1789.34$\pm$0.02%
+Wake Vision 83.79$\pm$0.2384.89$\pm$0.1185.72$\pm$0.04%
+3.4 Community Accessibility and Involvement
+Wake Vision and its fine-grained benchmark suite are available through TensorFlow Datasets [45] and HuggingFace Datasets [30] to enable easy access and use by the community.
+The images and labels are rehosted to ensure the dataset will not shrink due to dead links.
+The rehosted labels are generated according to our default dataset configuration, further described in Appendix I.
+Most datasets remain static after the initial release.
+To continuously improve Wake Vision we partner with the Edge AI Foundation [14] to gather community involvement through recurrent competitions (Sec.
+4.5).
+In our vision, the iterative enhancements brought by the community are integrated into the dataset leading to a continuous improvement cycle.
+3.5 Generating Binary Image Classification Datasets for TinyML
+Beyond person detection, many other TinyML applications can benefit from large-scale binary classification datasets.
+For example, bird detection enables smart nest systems [9] and automated feeding  [20], while vehicle detection enables automated parking space occupancy monitoring [6] and blind spot monitoring [43].
+We demonstrate that our binary dataset generation pipeline is adaptable for constructing image classification datasets for these and other classes of interest among the 9.6K trainable classes or 600 boxable objects present in Open Images v7 [29, 27].
+Constructing binary classification datasets using our pipeline yields significantly larger datasets compared to the prior VWW [7] methodology applied to the same class.
+For the bird category, the dataset generated via our pipeline contains 12×12$\times$12 × more images than VWW, with a decrease in label error rate from 6.4% to 4.8%.
+Similarly, for car detection, we provide 27×27$\times$27 × more images with a decrease from 6.6% to 0.6% in label error rate.
+Further details are available in Appendix S.
+4 Results
+We present a comprehensive evaluation of Wake Vision through five complementary analyses.
+First, we compare our two training sets to derive insights on best practices when training.
+Second, we show that Wake Vision is an effective drop-in replacement for Visual Wake Words by conducting cross-evaluation studies using identical MobileNetV2 models.
+Third, we validate Wake Vision’s benefits across diverse model architectures used in the TinyML community, from MCUNet to ColabNAS families.
+Fourth, we present fine-grained benchmark results across demographics, environmental conditions, and visual contexts to assess model robustness in real-world scenarios.
+Fifth, we present the results of the first Wake Vision Challenge and its resulting community-driven improvements to the Wake Vision dataset.
+Through these analyses, we demonstrate that Wake Vision not only improves upon VWW’s performance but also enables a more thorough evaluation of TinyML computer vision models.
+4.1 Training Set Evaluation
+Table 4: Accuracy on test set of a MobileNetV2-0.25 trained on image-level and bounding box labels, with and without KD. A MobileNetV2-1.0 model trained on the bounding box set is the teacher.
+Training set Label Source Training Set Size Accuracy Distilled Accuracy
+Wake Vision (Large) Image-Level 5,760,428 80.80$\pm$0.1885.00$\pm$0.20%
+Wake Vision (Quality) Bounding Box 1,248,230 84.89$\pm$0.1185.17$\pm$0.18%
+Wake Vision (Combined) Both 5,760,428 85.72$\pm$0.04-
+To evaluate the performance of the two Wake Vision training sets, we train identical models on each set.
+We use MobileNetV2 [41] models with a width modifier of 0.25 for 200,000 steps on 224x224x3 images using AdamW [34], a learning rate of 2e-3 with cosine decay and a weight decay of 4e-6.
+After training, we compare each model’s performance on the common test set.
+The result of this comparison is shown in Table 4.
+The Wake Vision (Quality) training set outperforms the Wake Vision (Large) training set by 4.09% test accuracy, indicating that label quality is more important than quantity in this evaluation setting.
+The gap between the two is reduced to just 0.17% when training exclusively on soft labels from a teacher model (MobileNetV2-1.0) trained on bounding box labels.
+We further study how training set size and quality impact accuracy across model sizes, and observe that smaller models are more sensitive to errors in the training set, dropping accuracy by a delta of 0.8% compared to larger models (Appendix C).
+Using the two training sets in unison to train a model where Wake Vision (Large) acts as a pre-training set, and Wake Vision (Quality) acts as a fine-tuning set achieves our best performance at a mean top-level test accuracy of 85.72%.
+The two training sets together provide a foundation for further research on data-centric TinyML [35, 36].
+4.2 Wake Vision & VWW Cross Evaluation
+A key challenge in advancing TinyML research is ensuring that improvements can be readily adopted by the community without requiring significant modifications to existing workflows or architectures.
+While novel datasets can offer better quality or scale, their practical impact is limited if they require substantial changes to model architectures, training pipelines, or deployment processes.
+Therefore, we designed Wake Vision to serve as a drop-in replacement for VWW.
+To validate this drop-in capability, we do a cross-evaluation study.
+We train two identical MobileNetv2 [41] models according to the same training recipe as in Section 4.1.
+One model uses VWW’s training set, and the other uses Wake Vision’s Quality training set.
+After training these identical models using the respective datasets’ training sets with the same recipe for an equal number of steps, we evaluate their performance on the corresponding test sets.
+Table 3 shows the result of the cross-evaluation.
+The Wake Vision trained model outperforms the VWW trained model with a 0.26% improvement on VWW’s own test set, which indicates that Wake Vision is a direct improvement over VWW, not simply a domain shift.
+Also, we achieve a 1.1% improvement over VWW on Wake Vision’s Test set, indicating that the new test set is more challenging.
+Table 3 also shows the performance of the Wake Vision (Combined) model described in Section 3.1.
+This model achieves 0.75% improvement over the Wake Vision (Quality) trained model on the VWW test set and 0.83% improvement on the Wake Vision test set.
+The Wake Vision (Combined) model is trained for longer than the VWW and Wake Vision (Quality) models but shows the value of using the training sets in unison even on the VWW test set.
+4.3 Comprehensive Model Architecture Evaluation
+To show that the benefits of using Wake Vision over VWW are not limited to MobileNetv2 architectures, we also train several other models on both training sets, Wake Vision (Quality) and VWW, and compare their performance.
+For this experiment, we use a collection of models commonly used in the TinyML community, such as models from the MCUNet [32] and Micronets [2] families.
+Furthermore, we include models from ColabNAS [18] to provide evaluation results on even smaller models.
+(a) Initial Pareto frontier for Wake Vision vs VWW
+(b) Advanced Pareto frontier by challenge submissions
+Figure 3: The Wake Vision Challenge submissions advanced our initial Pareto frontiers.
+Each model is trained on each dataset using the Adam optimizer with a learning rate of 1e-3 for 243,800 steps at a batch size of 512, equivalent to 100 epochs on the Wake Vision (Quality) training set.
+In addition, we apply random horizontal flips rotations and use 8-bit quantization-aware training.
+We use the checkpoint with the best validation score for testing.
+In our comparison, we cross-evaluate each trained model’s performance on the test sets of both VWW and Wake Vision.
+Results on the Wake Vision test set can be found in Figure 3(a), while further details and the results for the VWW test set can be found in Appendix A.
+For the majority of models, the one trained on the Wake Vision (Quality) set achieves a higher accuracy on both the VWW and Wake Vision test set, which supports our claim that Wake Vision is a direct improvement over VWW.
+At the extreme, for one Micronets [2] model (middle right of Figure 3(a)), switching the training set to Wake Vision increased the test accuracy by as much as 6.6%.
+The only instance where Wake Vision isn’t a complete improvement over VWW is in Figure 4 in Appendix A, where small models perform better when trained on the same dataset they are tested on (i.e., trained on VWW and tested on VWW), likely due to ultra-low-capacity models failing to generalize past the slight domain shift between the VWW and Wake Vision test sets.
+Accuracy and performance metrics, such as latency or model size, are in constant tension in TinyML models.
+Therefore, any boost in accuracy achieved through Wake Vision can also be traded off for more efficient architectures (in some cases 3X fewer MACs) while preserving test accuracy.
+4.4 Fine-Grained Benchmark Evaluation
+Aggregate metrics like accuracy or F1-score provide high-level insight but can mask critical failure modes and biases that emerge in real-world scenarios.
+In response, we include fine-grained benchmarks in Wake Vision (Section 3.3) across demographics, environmental conditions, and visual contexts.
+This is particularly important for person detection models deployed in real-world applications, where failures on certain subgroups or conditions could adversely affect fairness and safety.
+In Table 5, we compare the models from Section 4.2, trained on Wake Vision and VWW, across various scenarios to understand their relative strengths and limitations.
+The Wake Vision model exhibits superior robustness across the challenging settings exercised by our benchmarking suite.
+For instance, on the “Depictions” benchmark, which evaluates performance on images containing persons, non-person objects, or no depictions, the Wake Vision model achieves an F1 score of 0.71 for person depictions, outperforming the VWW model’s 0.66.
+The Wake Vision model also showcases improved performance on the “Age” benchmark, with F1 scores of 0.94, 0.91, and 0.94 for young, middle, and older individuals, surpassing the VWW model’s scores of 0.90, 0.88, and 0.89, respectively.
+This highlights the dataset’s effectiveness in enhancing model robustness for detecting people across various age groups, particularly the elderly demographic.
+In addition, the Wake Vision model demonstrates resilience to challenging lighting conditions, achieving F1 scores of 0.85 in dark lighting scenarios, compared to the VWW’s 0.81.
+Notably, a VWW model only significantly outperforms a Wake Vision model on the far distance benchmark at a score of 0.67 and 0.59, respectively.
+During dataset generation,VWW filters out persons smaller than 0.5% of the image, versus 5% for Wake Vision.
+Therefore, VWW contains images with significantly smaller persons.
+We did experiment with adopting a default of 0.5% for Wake Vision, however this disproportionately decreased the performance of models in other settings.
+Table 5: Wake Vision Fine-Grained Benchmark Suite. We report the samples in each set and the average F1 score across three Wake Vision models and three VWW models on each benchmark.
+Gender Age Distance Lighting Depictions
+Female
+Male
+Unknown
+Young
+Middle
+Older
+Unknown
+Near
+Medium
+Far
+Dark
+Normal
+Bright
+Person
+Non-Person
+No Depiction
+Size Val 684 1310 1612 275 2133 90 1299 5457 2213 398 3255 14315 1012 356 352 8583
+Test 2157 3918 4940 884 6595 276 3837 16333 6876 1140 9420 43010 3332 978 1101 25802
+F1 Wake Vision 0.93 0.91 0.77 0.94 0.91 0.94 0.71 0.91 0.85 0.59 0.85 0.82 0.82 0.71 0.86 0.87
+VWW 0.89 0.88 0.78 0.90 0.88 0.89 0.74 0.89 0.84 0.67 0.81 0.82 0.82 0.66 0.82 0.85
+Collectively, these results underscore the significance of Wake Vision’s large-scale, diverse data and fine-grained benchmarks, enabling the development of more robust and reliable TinyML person detection models across a wide range of realistic and real-world scenarios.
+4.5 The Wake Vision Flywheel: Community-Driven Continuous Improvement
+Unlike most datasets which remain static, Wake Vision employs a community-driven flywheel where each competition generates improvements that fuel greater participation and innovation in subsequent rounds, thereby advancing TinyML research and commercial applications.
+We partner with the Edge AI Foundation [14] to host Wake Vision competitions.
+The inaugural competition in early 2025 featured two tracks, one focused on developing high-performance, resource-efficient models, and the other on enhancing the Wake Vision person detection dataset.
+The event garnered significant support from the TinyML community, with 48 participants across both tracks [12].
+Successful challenge outcomes included an advanced pareto frontier (Figure 3(b)), containing 3 new models (orange), and a technique to enhance the quality of the Wake Vision large split [4], which in our experiments led to a reduction in the label error rate from 15.2% to 9.8%.
+Further details can be found in Appendix T.
+A new competition is currently being launched [13] to further community-led advancements to Wake Vision, and the best results will be integrated in a new version of our dataset.
+This competition-driven advancement exemplifies our vision of continuous dataset improvement.
+5 Ethical Considerations and Limitations
+We acknowledge several ethical considerations.
+Potential Misuse: Our dataset and benchmark suite are designed to advance TinyML research and applications.
+Intended benefits include occupancy detection systems that conserve energy by turning off the lights/HVAC/TV when a person leaves, and privacy preservation through on-device inference, avoiding cloud processing of camera data.
+We acknowledge person detection technology can be misappropriated for harmful purposes such as weapons targeting or mass surveillance systems.
+Data Rights and Privacy: Images are sourced from Flickr through Open Images under the CC-BY 2.0 license.
+Some images may have been uploaded without proper distribution rights, or may contain potentially identifying biometric information.
+Fairness and Representation: Our benchmark suite is designed to evaluate model fairness and robustness across different demographics and conditions, ensuring person detection systems work equitably for all populations.
+However, demographic benchmarks can be misused towards classifying gender and age without consent, potentially enabling privacy violations or discriminatory practices.
+6 Conclusion
+We introduced an automated binary dataset generation pipeline that addresses a fundamental challenge in TinyML: the systematic creation of large-scale, high-quality datasets tailored to diverse deployment constraints.
+Our approach combines intelligent multi-source label fusion, confidence-aware filtering, automated label correction, and systematic benchmark generation—innovations necessitated by the prohibitive cost and inconsistency of manual approaches for TinyML’s bespoke dataset requirements.
+Through Wake Vision, a large-scale binary classification dataset demonstrating our methodology via person detection, we achieved 100$\times$ scale improvement and 6.6% accuracy gains over existing benchmarks.
+More importantly, we uncovered critical TinyML-specific insights: smaller models exhibit greater sensitivity to label quality than their large-scale counterparts, and optimal performance requires carefully orchestrated two-stage training that balances scale and quality.
+Our fine-grained evaluation framework revealed performance disparities across demographic and environmental conditions that aggregate metrics fail to capture.
+The generality of our approach extends beyond person detection, successfully creating production-grade datasets for bird and car detection with similar improvements.
+By enabling systematic dataset creation across Open Images’ 20K classes, our work provides a methodological foundation that can accelerate TinyML research across diverse application domains.
+All resources are open-sourced under CC-BY 4.0 license to maximize community impact.
+Appendix A Additional Benchmark Results
+Table 6 provides input shape, Flash, RAM, and MACs for each model presented in section 4.2 as well as the mean and standard deviation of test accuracies for both Wake Vision and VWW datasets.
+RAM and Flash have been measured using the “stm32tflm” utility of X-CUBE-AI 8.1.0, whereas MACs with “stm32ai”.
+Links pointing to the original models used for training are also present for each model family involved in the experiment.
+Figure 4 shows cross-evaluation results on the Visual Wake Words test set.
+The majority of models trained on the Wake Vision dataset obtain better test accuracy on the VWW than the ones trained on the VWW.
+Only the smallest models do not achieve a higher test accuracy when trained on the Wake Vision, likely due to the lack of enough resources to overcome the slight domain shift between the two datasets.
+Table 6: Raw data of reconstructed models.
+Model RAM Flash MACs Test Set Training Set
+.tflite [kiB] [kiB] [MM] WV (Qual.) VWW
+MobileNetV2 [41]
+MobileNetV2_0.25 1,244.5 410.55 36,453,732 WV $84.9\pm 0.11$ $83.8\pm 0.23$
+VWW $88.6\pm 0.17$ $88.3\pm 0.29$
+MCUNet [32]
+10fps_vww 168.5 533.84 5,998,334 WV $81.7\pm 0.28$ $77.6\pm 0.31$
+VWW $81\pm 0.12$ $80.8\pm 0.1$
+5fps_vww 226.5 624.55 11,645,502 WV $82.9\pm 0.29$ $79.6\pm 0.26$
+VWW $83.2\pm 0.57$ $82.6\pm 0.16$
+320kb-1mb_vww 393 923.76 56,022,934 WV $85.6\pm 0.34$ $82.4\pm 0.62$
+VWW $86.5\pm 1.01$ $86\pm 0.25$
+Micronets [2]
+vww2_50_50 71.50 225.54 3,167,382 WV $71.9\pm 0.67$ $66.2\pm 0.88$
+VWW $70.8\pm 0.68$ $70.6\pm 0.64$
+vww3_128_128 137.50 463.73 22,690,291 WV $77.8\pm 0.56$ $72.6\pm 1.01$
+VWW $78.3\pm 0.91$ $77.6\pm 1.13$
+vww4_128_128 123.50 417.03 18,963,302 WV $77.9\pm 0.6$ $71.3\pm 1.03$
+VWW $78.4\pm 0.51$ $76.5\pm 0.37$
+ColabNAS [18]
+k_2_c_3 18.5 7.66 250,256 WV $70.6\pm 0.96$ $69.3\pm 0.97$
+VWW $65.6\pm 0.66$ $70.7\pm 0.08$
+k_4_c_5 22 18.49 688,790 WV $75.7\pm 0.18$ $74\pm 0.23$
+VWW $69.9\pm 0.26$ $75.5\pm 0,64$
+k_8_c_5 32.5 44.56 2,135,476 WV $77.3\pm 0.37$ $75\pm 0.15$
+VWW $73\pm 0.91$ $77.3\pm 0.57$
+Figure 4: Cross-evaluation results on the Visual Wake Words test set
+Appendix B Code Repository
+The code used to generate the Wake Vision dataset is available at the following GitHub repository: https://github.com/harvard-edge/Wake_Vision
+This repo contains the code to generate Wake Vision and the benchmark suite, as well as the code to train and evaluate models. This code is sufficient to reproduce all results in the paper.
+Appendix C Impact of Dataset Quality & Size in TinyML
+Understanding the relationship between dataset quality and size is important for training efficient models in real-world scenarios.
+While gathering large amounts of data has become easier, ensuring high-label quality remains costly and time-consuming.
+This creates an important trade-off: is it better to have a smaller dataset with high-quality labels or a larger dataset with more label noise?
+The answer likely depends on the capacity of the model and the specific task.
+We investigate the impact of dataset quality and size on Resnet [22] style models of varying capacity.
+We measure quality by the approximate rate of label errors.
+The Wake Vision (Quality) training set has an estimated error rate of around 7%.
+We simulate higher error rate datasets (15% and 30%) by flipping the binary label with a certain probability.
+Appendix C.1 gives the derivation for the flip probability.
+We also sweep the dataset size by taking a slice of Wake Vision (Quality).
+Each model is trained for 50,000 steps on 224x224x3 images using AdamW [34], a learning rate of 0.001 with a cosine schedule and a weight decay of 0.004.
+Figure 5 shows that smaller models (leftmost figure) are more sensitive to higher error rates than large models, similar to prior findings that model capacity can help memorize and ignore errors and outliers [49, 16, 21].
+The lowest accuracy drop in the smallest model is 1.
+3%, going from the base train (quality) error rate of around 7% to 15% compared to only a 0.5% accuracy loss for the largest model.
+In contrast, large models benefit more from big datasets and start to overfit on the smaller slices of the training data [26].
+We also see the large slice of the 15% error rate dataset outperforming a smaller slice of the 7% error rate dataset.
+Figure 5: Impact of dataset error rate and size on models of varying capacity.
+The $y$-axis is the Wake Vision test accuracy and the $x$-axis is the percentage of Wake Vision Train (Quality) used in training.
+The error % indicates the expected rate of incorrectly labeled samples.
+C.1 Inducing Label Errors
+The goal is to make a single pass through the dataset and flip the labels of a binary classification dataset such that the expected label error rate is $d$.
+There is an underlying rate of label errors $e$.
+If we flip one of these underlying errors, we correct it, thereby inadvertently decreasing the label error rate.
+After flipping labels with a probability of $p$ we can claim that the likelihood of a single label being correct is the probability that we flipped the label and it was originally an error plus the probability that we didn’t flip the label and it was originally correct: $1-d=p*e+(1-p)(1-e)$.
+Then solving for $p$ gives $p=(e-d)/(2e-1)$.
+A current flaw of this method is that the injected label errors are not consistent between epochs, which would likely be less destructive to a model’s accuracy since the same errors are not reinforced each epoch.
+This could also potentially explain why large models in Fig.
+5 don’t overfit on training data with higher label errors, as the inconsistent label noise has a regularizing effect.
+Appendix D Flowchart of Bounding Box Filtering Process
+Figure 6 illustrates the filtering process for Wake Vision when using Open Images’ bounding box labels as the label source.
+Figure 6: A flowchart of the filtering of an image according to default Bounding Box filtering rules.
+Figure 7: Effects of scaling the image size vs.
+the model width on Wake Vision test accuracy and the Distance-Far benchmark set.
+The Distance-Far F1 score (right) is far more sensitive to changes in the image size than the overall test accuracy (left).
+Image sizes refer to square image resolutions (e.g.
+96x96) and model size refers to the MobileNetV2 width multiplier.
+When not specified the image size is 160 and the width multiplier is 1.0.
+Appendix E Model Design Case Study
+Basic test set performance can often misrepresent a model’s performance, given the typical domain shift between images scraped from the internet and real-world use cases.
+This issue is exacerbated when ML practitioners must trade off accuracy for model performance and size, which is necessary for TinyML use cases.
+A design decision might have seemingly little impact on the test accuracy but may destroy real-world performance depending on the deployment environment.
+For example, a person detection system may only operate in dark lighting conditions, but the test dataset has an insignificant number of dark samples; therefore, the test accuracy will not reflect the real-world accuracy.
+The benchmark suite enables more holistic analysis during the design phase.
+To show this use we perform a series of scaling experiments employing typical TinyML compression techniques and identifying under which circumstances these techniques are appropriate.
+While these results can inform ML practitioners, our intention is to demonstrate the usefulness of the benchmark suite.
+E.1 Image Size vs. Model Width Scaling
+We train two series of MobileNetV2 models: one series that sweeps the input image size [64-256] and one that sweeps the width multiplier of a model [0.1-1.5].
+We then benchmark these models on the Wake Vision test set as well as the far distance benchmark.
+We plot these results against the number of multiply-accumulate (MAC) operations in the model as a proxy for on-device latency [2].
+The results in Table 7 (left) show that when looking exclusively at the high-level metric (i.e., test accuracy), scaling the input image size has a similar impact as scaling the model size.
+However, as shown in Figure 7 (right), when we consider only samples where the person is far away from the camera (i.e., the distance benchmark), we observe a much more significant impact when scaling the image size.
+In the case of distant subjects, the image size becomes the bottleneck.
+These findings suggest that for ML developers targeting use cases where the subject is likely to be far from the camera, prioritizing larger input image sizes over wider models may be more beneficial.
+However, this critical design consideration could be obscured when solely relying on high-level metrics like overall test accuracy.
+The distance benchmark in Wake Vision effectively unveils the disproportionate impact of image size on model performance for distant subjects, enabling more informed decision-making during model optimization.
+E.2 Quantization
+Quantization is a crucial technique for deploying efficient TinyML models, offering substantial benefits in terms of reduced latency, memory footprint, and model size.
+However, prior work has suggested that quantization can disproportionately impact the performance of models on underrepresented subsets of data [25].
+To assess the implications of quantization in the context of person detection, we investigate the impacts of int8 quantization on a model’s benchmark results across Wake Vision’s fine-grained benchmarks.
+Our findings show negligible degradation in performance across all benchmarks (±$\pm 0.004$ F1) when employing int8 quantization, even on outlier sets.
+This result contradicts the previously observed disproportionate impact of quantization on underrepresented subsets.
+We speculate that person detection may be a relatively simple task, potentially explaining why we do not observe this specific property of quantization in our experiments.
+Given the negligible performance degradation and the substantial latency, memory, and model size benefits of quantization, we conclude that quantization is a win for person detection.
+E.3 Grayscale
+Converting a model’s input image channels to grayscale from RGB is a commonly employed optimization in the TinyML field [2] as it can substantially reduce a model’s memory consumption.
+We observed, however, that the grayscale optimization disproportionately impacts images on the brighter end of the spectrum as illustrated in Figure 8.
+This further demonstrates the importance of fine-grained analysis, as some real-world deployment environments might be far brighter than the average Wake Vision test sample.
+Figure 8: Impact of grayscale input images on the lighting benchmarks: dark, normal light, and bright. Models that use grayscale input images are more sensitive to bright lighting conditions than RGB.
+Appendix F Fine Grained Benchmark Suite
+This section details the fine grained benchmark suite proposed.
 Perceived Gender and Age
-Table 3: Accuracy on the Wake Vision (WV) and VWW test sets by models trained on the VWW, Wake Vision (Quality) and Wake Vision (Combined) training sets.
-Train
-VWW Wake Vision (Quality) WV (Combined)
-Test VWW 88.33$\pm$0.2988.59$\pm$0.1789.34$\pm$0.02%
-WV 83.79$\pm$0.2384.89$\pm$0.1185.72$\pm$0.04%
 Underrepresented sub-groups typically constitute a small portion of a generic test set; therefore, top-line metrics often obscure a bias in a model until it is deployed [24].
@@
 Lighting
-The lighting data sets test the performance of Machine Learning (ML) systems in different lighting conditions.
+The lighting data sets test the performance of ML systems in different lighting conditions.
 In scenarios like security monitoring, outdoor robotics, or augmented reality applications, models must be robust to varying lighting conditions, including low-light environments.
@@
 Depictions of people can range from photo-realistic to crude stick figures.
-3.4 Usability
-Wake Vision and its fine-grained benchmark suite are available through TensorFlow Datasets [36] and HuggingFace Datasets [22] to enable easy access and use by the community.
-The images and labels are rehosted to ensure the dataset will not be changed over time due to dead links, making it a more stable benchmark.
-The rehosted labels are generated according to our default dataset configuration, further described in Appendix I.
-4 Results
-In this section, we present a comprehensive evaluation of Wake Vision through five complementary analyses.
-First, we provide the first insights into best practices for Wake Vision by comparing its two training sets, which also provide initial insights into the trade-off between dataset quality and size in a TinyML setting.
-Second, we show that Wake Vision serves as an effective drop-in replacement for Visual Wake Words by conducting cross-evaluation studies using identical MobileNetV2 models.
-Third, we validate Wake Vision’s benefits across diverse model architectures commonly used in the TinyML community, from MCUNet to ColabNAS families.
-Fourth, we thoroughly investigate the critical trade-off between dataset quality and size for TinyML models, providing insights into how model capacity influences sensitivity to label noise versus dataset scale.
-Finally, we present fine-grained benchmark results across demographics, environmental conditions, and visual contexts to assess model robustness in real-world scenarios.
-Throughout these analyses, we demonstrate that Wake Vision not only improves upon VWW’s performance but also enables a more thorough evaluation of TinyML computer vision models.
-4.1 Training Set Evaluation
-Table 4: Accuracy on test set of a MobileNetV2-0.25 trained on image-level and bounding box labels, with and without distillation.
-A MobileNetV2-1.0 model trained on the bounding box set is used as the teacher in KD.
-Training set Label Source Training Set Size Accuracy Distilled Accuracy
-Wake Vision (Large) Image-Level 5,760,428 80.80$\pm$0.1885.00$\pm$0.20%
-Wake Vision (Quality) Bounding Box 1,248,230 84.89$\pm$0.1185.17$\pm$0.18%
-Wake Vision (Combined) Both 5,760,428 85.72$\pm$0.04-
-To evaluate the performance of the two Wake Vision training sets, we train identical models on each set.
-We use MobileNetV2 [33] models with a width modifier of 0.25 for 200,000 steps on 224x224x3 images using AdamW [26].
-We use a lr of 0.002 with a cosine decay and a weight decay of 4E-6.
-After training, we compare each model’s performance on the common test set.
-The result of this comparison is shown in Tab.
-4.
-These results show that the Wake Vision (Quality) training set outperforms the Wake Vision (Large) training set by 4.09% test accuracy.
-This indicates that label quality is more important than quantity in this setting.
-The gap between the two is reduced to just 0.17% when training exclusively on soft labels from a teacher model (MobileNetV2-1.0) trained on bounding box labels.
-Using the two training sets in unison to train a model where Wake Vision (Large) acts as a pre-training set, and Wake Vision (Quality) acts as a fine-tuning set achieves our best performance at a mean top-level test accuracy of 85.72%.
-The two training sets together provide a foundation for further research on data-centric TinyML [27, 28].
-4.2 Wake Vision & VWW Cross Evaluation
-A key challenge in advancing TinyML research is ensuring that improvements can be readily adopted by the community without requiring significant modifications to existing workflows or architectures.
-While novel datasets can offer better quality or scale, their practical impact is limited if they require substantial changes to model architectures, training pipelines, or deployment processes.
-Therefore, we designed Wake Vision to serve as a drop-in replacement for VWW.
-To validate this drop-in capability and demonstrate Wake Vision’s effectiveness, we do a cross-evaluation study.
-We train two identical MobileNetv2 [33] models according to the same training recipe as in Sec.
-4.1.
-One model uses VWW’s training set, and the other uses Wake Vision’s Quality training set.
-After training these identical models using the respective datasets’ training sets with the same recipe for an equal number of steps, we evaluate their performance on the corresponding test sets.
-Creftab:wv-vww-cross-val shows the result of the cross-evaluation.
-The Wake Vision trained model outperforms the VWW trained model with a 0.26% improvement on VWW’s own test set, which indicates that Wake Vision is a direct improvement over VWW, not simply a domain shift.
-Also, we achieve a 1.1% improvement over VWW on Wake Vision’s Test set, indicating that the new test set is more challenging.
-Table 3 also shows the performance of the Wake Vision (Combined) model described in Sec.
-3.1.
-This model achieves 0.75% improvement over the Wake Vision (Quality) trained model on the VWW test set and 0.83% improvement on the Wake Vision test set.
-The Wake Vision (Combined) model is trained for longer than the VWW and Wake Vision (Quality) models but shows the value of using the training sets in unison even on the VWW test set.
-4.3 Comprehensive Model Architecture Evaluation
-To show that the benefits of using Wake Vision over VWW are not limited to MobileNetv2 architectures, we also train several other models on both training sets, Wake Vision (Quality) and VWW, and compare their performance.
-For this experiment, we use a collection of models commonly used in the TinyML community, such as models from the MCUNet [24] and Micronets [2] families.
-Furthermore, we include models from ColabNAS [11] to provide evaluation results on even smaller models.
-Figure 3: Cross-evaluation results on the Wake Vision test set.
-Each model is trained on each dataset using the Adam optimizer with a learning rate of 0.001 for 243,800 steps at a batch size of 512, equivalent to 100 epochs on the Wake Vision (Quality) training set.
-In addition, we apply random horizontal flips rotations and use 8-bit quantization-aware training.
-We use the checkpoint with the best validation score for testing.
-In our comparison, we cross-evaluate each trained model’s performance on the test sets of both VWW and Wake Vision.
-Results on the Wake Vision test set can be found in Fig.
-3, while the results for the VWW test set can be found in Fig.
-5 alongside the complete data in Appendix A.
-For the majority of models, the one trained on the Wake Vision (Quality) set achieves a higher accuracy on both the VWW and Wake Vision test set, which supports our claim that Wake Vision is a direct improvement over VWW.
-In some models, the VWW trained variant performs better on the VWW test set.
-This is only the case for the smallest models in our evaluation suite, which we believe may be due to these models being too small to generalize past the slight domain shift between the VWW and Wake Vision test sets.
-We additionally see that the accuracy discrepancy between identical models is often highest on Wake Vision’s Test set, further suggesting that the new test set is more challenging.
-The boost in accuracy achieved by using the Wake Vision (Quality) Training set can be traded off for smaller model architectures (in some cases 3X fewer MACs) while preserving test accuracy.
-Accuracy and performance metrics, such as latency or model size, are in constant tension in TinyML models.
-Therefore, any boost in accuracy that does not impact the model’s architecture enables further efficiency.
-4.4 Impact of Dataset Quality & Size in TinyML
-Understanding the relationship between dataset quality and size is important for training efficient models in real-world scenarios.
-While gathering large amounts of data has become easier, ensuring high-label quality remains costly and time-consuming.
-This creates an important trade-off: is it better to have a smaller dataset with high-quality labels or a larger dataset with more label noise?
-The answer likely depends on the capacity of the model and the specific task.
-We investigate the impact of dataset quality and size on Resnet [14] style models of varying capacity.
-We measure quality by the approximate rate of label errors.
-The Wake Vision (Quality) training set has an estimated error rate of around 7%.
-We simulate higher error rate datasets (15% and 30%) by flipping the binary label with a certain probability.
-Appendix E gives the derivation for the flip probability.
-We also sweep the dataset size by taking a slice of Wake Vision (Quality).
-Each model is trained for 50,000 steps on 224x224x3 images using AdamW [26], a learning rate of 0.001 with a cosine schedule and a weight decay of 0.004.
-Figure 4 shows that smaller models (leftmost figure) are more sensitive to higher error rates than large models, similar to prior findings that model capacity can help memorize and ignore errors and outliers [40, 9, 13].
-The lowest accuracy drop in the smallest model is 1.
-3%, going from the base train (quality) error rate of around 7% to 15% compared to only a 0.5% accuracy loss for the largest model.
-In contrast, large models benefit more from big datasets and start to overfit on the smaller slices of the training data [18].
-We also see the large slice of the 15% error rate dataset outperforming a smaller slice of the 7% error rate dataset.
-Figure 4: Impact of dataset error rate and size on models of varying capacity.
-The $y$-axis is the Wake Vision test accuracy and the $x$-axis is the percentage of Wake Vision Train (Quality) used in training.
-The error % indicates the expected rate of incorrectly labeled samples.
-4.5 Fine-Grained Benchmark Evaluation
-Although aggregate metrics like overall accuracy or F1 score provide a high-level view of model performance, they mask critical failure modes and biases that emerge in real-world scenarios.
-For this reason, we included fine-grained benchmarks in Wake Vision (Section 3.3).
-Fine-grained evaluation across demographics, environmental conditions, and visual contexts helps us develop reliable TinyML Computer Vision systems for bespoke use cases.
-This is particularly important for person detection models deployed in real-world applications, where failures on certain subgroups or conditions could have serious implications for fairness and safety.
-In Table 5, we compare the models from Sec.
-4.2, trained on Wake Vision and VWW, across various scenarios to understand their relative strengths and limitations.
-The Wake Vision model exhibits superior robustness across the challenging settings exercised by our benchmarking suite, further demonstrating the effectiveness of the dataset.
-For instance, on the “Depictions” benchmark, which evaluates performance on images containing persons, non-person objects, or no depictions, the Wake Vision model achieves an F1 score of 0.71 for person depictions, outperforming the VWW model’s 0.66.
-Similarly, the Wake Vision model scores 0.86 for non-person depictions compared to the VWW model’s 0.82.
-Table 5: Wake Vision Fine Grain Benchmark Suite. We report the samples in each set and the average F1 score across three Wake Vision models and three VWW models on the benchmark set.
-Set Size F1-Score
-Benchmark Val Test Wake Vision VWW
-Gender Female $684$ $2157$ $0.93$ $0.89$
-Male $1310$ $3918$ $0.91$ $0.88$
-Unknown $1612$ $4940$ $0.77$ $0.78$
-Age Young $275$ $884$ $0.94$ $0.90$
-Middle $2133$ $6595$ $0.91$ $0.88$
-Older $90$ $276$ $0.94$ $0.89$
-Unknown $1299$ $3837$ $0.71$ $0.74$
-Distance Near $5457$ 16 3331633316\,33316 333 $0.91$ $0.89$
-Medium $2213$ $6876$ $0.85$ $0.84$
-Far $398$ $1140$ $0.59$ $0.67$
-Lighting Dark $3255$ $9420$ $0.85$ $0.81$
-Normal 14 3151431514\,31514 315 43 0104301043\,01043 010 $0.82$ $0.82$
-Bright $1012$ $3332$ $0.82$ $0.82$
-Depictions Person $356$ $978$ $0.71$ $0.66$
-Non-Person $352$ $1101$ $0.86$ $0.82$
-No Depiction $8583$ 25 8022580225\,80225 802 $0.87$ $0.85$
-The Wake Vision model also showcases improved performance on the “Age” benchmark, with F1 scores of 0.94, 0.91, and 0.94 for young, middle, and older individuals, surpassing the VWW model’s scores of 0.90, 0.88, and 0.89, respectively.
-This highlights the dataset’s effectiveness in enhancing model robustness for detecting people across various age groups, particularly the elderly demographic.
-In addition, the Wake Vision model demonstrates resilience to challenging lighting conditions, achieving F1 scores of 0.85 in dark lighting scenarios, compared to the VWW’s 0.81.
-Notably, a VWW model only significantly outperforms a Wake Vision model on the far distance dataset at a score of 0.67 and 0.59, respectively.
-This is likely caused by a difference in how the two datasets are created.
-Where VWW filters out persons smaller than 0.5% of the image, this filter is set to 5% for Wake Vision.
-Therefore, the VWW dataset contains images with significantly smaller persons than the Wake Vision dataset, which can improve its performance on far away persons.
-We did experiment with setting the limit to 0.5% for Wake Vision however, we experienced that it disproportionately decreased the performance of models trained on the dataset in other settings.
-Collectively, these results underscore the significance of Wake Vision’s large-scale, diverse data and fine-grained benchmarks, enabling the development of more robust and reliable TinyML person detection models across a wide range of realistic and real-world scenarios.
-5 Ethical Considerations
-The Wake Vision dataset and benchmark suite are designed to advance TinyML research by providing data resources and enabling detailed evaluation of person detection systems.
-While this work has significant potential benefits, we acknowledge several important ethical considerations:
-Intended Benefits
-Wake Vision can improve the quality of person detection systems that serve beneficial societal purposes, for example, an occupancy detection system that conserves energy by turning off the lights/HVAC/TV when a person leaves.
-The focus on TinyML and on-device inference helps protect user privacy by avoiding cloud-based processing of sensitive data.
-Potential Misuse
-We recognize that person detection technology can be misappropriated for harmful purposes, including weapons targeting or mass surveillance systems.
-While we cannot completely prevent misuse, we implement several safeguards, including releasing only evaluation sets, not training sets, for our fine-grained benchmark datasets.
-Data Rights and Privacy
-The images in Wake Vision are sourced from Flickr through Open Images [21, 19] under the CC-BY 2.0 license.
-While we have made extensive efforts to ensure proper licensing, we acknowledge the possibility that some images may have been uploaded without proper distribution rights.
-Furthermore, the dataset may contain potentially identifying biometric information, and given sufficient resources, individuals could potentially be identified.
-Due to the scale of the dataset, exhaustive manual verification for offensive content or privacy concerns is infeasible.
-Fairness and Representation
-Our benchmark suite is specifically designed to evaluate model fairness and robustness across different demographics and conditions.
-However, we acknowledge that these same demographic benchmarks could be misused to develop systems that classify gender and age without consent, potentially enabling privacy violations or discriminatory practices.
-We have carefully balanced these concerns against the importance of ensuring that person detection systems work equitably for all populations.
-6 Conclusion
-The lack of large, high-quality datasets has been a significant bottleneck in TinyML research.
-Through Wake Vision, we address this challenge by providing a comprehensive dataset and benchmark suite that advances the field in several key dimensions.
-Wake Vision surpasses the current state-of-the-art VWW dataset by offering  100x more images while maintaining ease of use and improving label accuracy.
-Our experiments show that Wake Vision serves as an effective drop-in replacement for VWW, consistently improving performance across various model architectures and training scenarios.
-Moreover, our investigation into dataset quality versus quantity trade-offs reveals important insights specific to TinyML systems.
-Unlike large models that can effectively learn from noisy data, we find that TinyML models are particularly sensitive to label quality because of their limited capacity.
-However, we also demonstrate that carefully designed training strategies—such as our approach of pre-training on Wake Vision (Large) followed by fine-tuning on Wake Vision (Quality)—can effectively leverage both quantity and quality to achieve superior performance.
-Finally, the Wake Vision benchmark suite, with its fine-grained evaluation sets, establishes new standards to evaluate the robustness of the TinyML model in demographics, environmental conditions, and visual contexts.
-These benchmarks enable researchers to identify and address potential biases and failure modes that could otherwise go undetected using traditional evaluation metrics.
-Appendix A Additional Benchmark Results
-Table Tab.
-6 provides input shape, Flash, RAM, and MACs for each model presented in section 4.2 as well as the mean and standard deviation of test accuracies for both Wake Vision and VWW datasets.
-RAM and Flash have been measured using the “stm32tflm” utility of X-CUBE-AI 8.1.0, whereas MACs with “stm32ai”.
-Links pointing to the original models used for training are also present for each model family involved in the experiment.
-Figure Fig.
-5 shows cross-evaluation results on the Visual Wake Words test set.
-The majority of models trained on the Wake Vision dataset obtain better test accuracy on the VWW than the ones trained on the VWW.
-Only the smallest models do not achieve a higher test accuracy when trained on the Wake Vision, likely due to the lack of enough resources to overcome the slight domain shift between the two datasets.
-Model input RAM Flash MACs Test Set Training Set
-.tflite volume [kiB] [kiB] [MM] WV (Qual.) VWW
-MobileNetV2  [^1] https://keras.io/api/applications/mobilenet/[33]
-MobileNetV2_0.25 (224,224,3) 1,244.5 410.55 36,453,732 WV $84.9\pm 0.11$ $83.8\pm 0.23$
-VWW $88.6\pm 0.17$ $88.3\pm 0.29$
-MCUNet  [^2] https://github.com/mit-han-lab/mcunet/blob/master/mcunet/model_zoo.py[24]
-10fps_vww (64,64,3) 168.5 533.84 5,998,334 WV $81.7\pm 0.28$ $77.6\pm 0.31$
-VWW $81\pm 0.12$ $80.8\pm 0.1$
-5fps_vww (80,80,3) 226.5 624.55 11,645,502 WV $82.9\pm 0.29$ $79.6\pm 0.26$
-VWW $83.2\pm 0.57$ $82.6\pm 0.16$
-320kb-1mb_vww (144,144,3) 393 923.76 56,022,934 WV $85.6\pm 0.34$ $82.4\pm 0.62$
-VWW $86.5\pm 1.01$ $86\pm 0.25$
-Micronets [^3] https://github.com/ARM-software/ML-zoo/tree/master/models/visual_wake_words[2]
-vww2_50_50_INT8 (50,50,1) 71.50 225.54 3,167,382 WV $71.9\pm 0.67$ $66.2\pm 0.88$
-VWW $70.8\pm 0.68$ $70.6\pm 0.64$
-vww3_128_128_INT8 (128,128,1) 137.50 463.73 22,690,291 WV $77.8\pm 0.56$ $72.6\pm 1.01$
-VWW $78.3\pm 0.91$ $77.6\pm 1.13$
-vww4_128_128_INT8 (128,128,1) 123.50 417.03 18,963,302 WV $77.9\pm 0.6$ $71.3\pm 1.03$
-VWW $78.4\pm 0.51$ $76.5\pm 0.37$
-ColabNAS444 https://github.com/AndreaMattiaGaravagno/ColabNAS/tree/main/models [11]
-k_2_c_3 (50,50,3) 18.5 7.66 250,256 WV $70.6\pm 0.96$ $69.3\pm 0.97$
-VWW $65.6\pm 0.66$ $70.7\pm 0.08$
-k_4_c_5 (50,50,3) 22 18.49 688,790 WV $75.7\pm 0.18$ $74\pm 0.23$
-VWW $69.9\pm 0.26$ $75.5\pm 0,64$
-k_8_c_5 (50,50,3) 32.5 44.56 2,135,476 WV $77.3\pm 0.37$ $75\pm 0.15$
-VWW $73\pm 0.91$ $77.3\pm 0.57$
-Table 6: Raw data of reconstructed models.
-Figure 5: Cross-evaluation results on the Visual Wake Words test set
-Appendix B Code Repository
-The code used to generate the Wake Vision dataset is available at the following GitHub repository: https://anonymous.4open.science/r/Wake_Vision-DD19/README.md
-This repo contains the code to generate Wake Vision and the benchmark suite, as well as the code to train and evaluate models. This code is sufficient to reproduce all results in the paper.
-Appendix C Flowchart of Bounding Box Filtering Process
-Fig. 6 illustrates the filtering process for Wake Vision when using Open Images’ bounding box labels as the label source.
-Figure 6: A flowchart of the filtering of an image according to default Bounding Box filtering rules.
-Figure 7: Effects of scaling the image size vs.
-the model width on Wake Vision test accuracy and the Distance-Far benchmark set.
-The Distance-Far F1 score (right) is far more sensitive to changes in the image size than the overall test accuracy (left).
-Image sizes refer to square image resolutions (e.g.
-96x96) and model size refers to the MobileNetV2 width multiplier.
-When not specified the image size is 160 and the width multiplier is 1.0.
-Appendix D Model Design Case Study
-Basic test set performance can often misrepresent a model’s performance, given the typical domain shift between images scraped from the internet and real-world use cases.
-This issue is exacerbated when ML practitioners must trade off accuracy for model performance and size, which is necessary for TinyML use cases.
-A design decision might have seemingly little impact on the test accuracy but may destroy real-world performance depending on the deployment environment.
-For example, a person detection system may only operate in dark lighting conditions, but the test dataset has an insignificant number of dark samples; therefore, the test accuracy will not reflect the real-world accuracy.
-The benchmark suite enables more holistic analysis during the design phase.
-To show this use we perform a series of scaling experiments employing typical TinyML compression techniques and identifying under which circumstances these techniques are appropriate.
-While these results can inform ML practitioners, our intention is to demonstrate the usefulness of the benchmark suite.
-D.1 Image Size vs. Model Width Scaling
-We train two series of MobileNetV2 models: one series that sweeps the input image size [64-256] and one that sweeps the width multiplier of a model [0.1-1.5].
-We then benchmark these models on the Wake Vision test set as well as the far distance benchmark.
-We plot these results against the number of multiply-accumulate (MAC) operations in the model as a proxy for on-device latency [2].
-The results in Table 7 (left) show that when looking exclusively at the high-level metric (i.e., test accuracy), scaling the input image size has a similar impact as scaling the model size.
-However, as shown in Figure 7 (right), when we consider only samples where the person is far away from the camera (i.e., the distance benchmark), we observe a much more significant impact when scaling the image size.
-In the case of distant subjects, the image size becomes the bottleneck.
-These findings suggest that for ML developers targeting use cases where the subject is likely to be far from the camera, prioritizing larger input image sizes over wider models may be more beneficial.
-However, this critical design consideration could be obscured when solely relying on high-level metrics like overall test accuracy.
-The distance benchmark in Wake Vision effectively unveils the disproportionate impact of image size on model performance for distant subjects, enabling more informed decision-making during model optimization.
-D.2 Quantization
-Quantization is a crucial technique for deploying efficient TinyML models, offering substantial benefits in terms of reduced latency, memory footprint, and model size.
-However, prior work has suggested that quantization can disproportionately impact the performance of models on underrepresented subsets of data [17].
-To assess the implications of quantization in the context of person detection, we investigate the impacts of int8 quantization on a model’s benchmark results across Wake Vision’s fine-grained benchmarks.
-Our findings show negligible degradation in performance across all benchmarks (±$\pm 0.004$ F1) when employing int8 quantization, even on outlier sets.
-This result contradicts the previously observed disproportionate impact of quantization on underrepresented subsets.
-We speculate that person detection may be a relatively simple task, potentially explaining why we do not observe this specific property of quantization in our experiments.
-Given the negligible performance degradation and the substantial latency, memory, and model size benefits of quantization, we conclude that quantization is a win for person detection.
-D.3 Grayscale
-Converting a model’s input image channels to grayscale from RGB is a commonly employed optimization in the TinyML field [2] as it can substantially reduce a model’s memory consumption.
-We observed, however, that the grayscale optimization disproportionately impacts images on the brighter end of the spectrum as illustrated in Figure 8.
-This further demonstrates the importance of fine-grained analysis, as some real-world deployment environments might be far brighter than the average Wake Vision test sample.
-Figure 8: Impact of grayscale input images on the lighting benchmarks: dark, normal light, and bright. Models that use grayscale input images are more sensitive to bright lighting conditions than RGB.
-Appendix E Inducing Label Errors
-The goal is to make a single pass through the dataset and flip the labels of a binary classification dataset such that the expected label error rate is $d$.
-There is an underlying rate of label errors $e$.
-If we flip one of these underlying errors, we correct it, thereby inadvertently decreasing the label error rate.
-After flipping labels with a probability of $p$ we can claim that the likelihood of a single label being correct is the probability that we flipped the label and it was originally an error plus the probability that we didn’t flip the label and it was originally correct: $1-d=p*e+(1-p)(1-e)$.
-Then solving for $p$ gives $p=(e-d)/(2e-1)$.
-A current flaw of this method is that the injected label errors are not consistent between epochs, which would likely be less destructive to a model’s accuracy since the same errors are not reinforced each epoch.
-This could also potentially explain why large models in Fig.
-4 don’t overfit on training data with higher label errors, as the inconsistent label noise has a regularizing effect.
-Appendix F Fine Grained Benchmark Images
-Fig. 9 gives example images for each of the benchmarks in the suite.
+Figure 9 gives example images for each of the benchmarks in the suite.
 Gender
@@
 Appendix H Dataset Access and Organization
-Wake Vision is available online via TFDS and Hugging Face Datasets. We exclude direct links to the dataset during the review period to avoid breaking anonymity.
+Wake Vision is available online at the Harvard Dataverse, via TFDS and Hugging Face Datasets.
+The version on Harvard Dataverse has been assigned the following DOI: https://doi.org/10.7910/DVN/1HOPXC More information can be found on the https://wakevision.ai webpage.
 The dataset is organized as a set of compressed tar files containing images and a series of label CSVs.
@@
 To allow for both data-centric and model-centric improvements, we provide a standard model for data-centric contributions, and a standard pre-processing pipeline for model-centric contributions.
-Both types of contributions are expected to use accuracy as the primary metric for the overall test and validation set, and F1-score as the primary metric for the fine grained benchmarks introduced in Sec.
-3.3.
+Both types of contributions are expected to use accuracy as the primary metric for the overall test and validation set, and F1-score as the primary metric for the fine grained benchmarks introduced in Section 3.3.
 Therefore, for data-centric contributions we propose to use a Mobilenet v2 model with a width modifier [^1] Also known as alpha.
@@
 As an unfortunate result of this, the dataset is not static over time as image owners can delete their images from the flickr platform.
-As a result we were only able to download a subset of the original Open Images v7 dataset as shown in Tab.
-7.
+As a result we were only able to download a subset of the original Open Images v7 dataset as shown in Table 7.
 Table 7: Number of images downloaded from Open Images v7. Download occurred between the 28th of November to the 5th of December
@@
   • Youth (Umbrella term for Girl & Boy - Only in Image Level Label Configuration)
-The following labels are considered body parts and are labelled according to the body part configuration.:
+The following labels are considered body parts and are labelled according to the body part configuration:
   • Human eye (Part of Person)
@@
 Appendix M Wake Vision Dataset Size
-See Tab. 8.
+See Table 8.
 Table 8: Amount of images in the Wake Vision dataset
@@
 To further correct label errors in the final Wake Vision Validation and Test sets we employed the Scale AI platform to crowd-source manual label corrections.
-These label corrections are described in Sec.
-3.2 and deprecate the automatic label corrections described in this section.
+These label corrections are described in Section 3.2 and deprecate the automatic label corrections described in this section.
 Appendix P Author Statement
@@
 Appendix R DOI and Croissant URLs
-The DOI and URLs will be provided after the review period to avoid breaking anonymity.
-Appendix S Datasheet for Wake Vision
+The version of Wake Vision on Harvard Dataverse has been assigned the following DOI: https://doi.org/10.7910/DVN/1HOPXC The following two croissant files describe the Wake Vision (Quality, Validation, and Test) and Wake Vision (Large) respectively: https://huggingface.co/api/datasets/Harvard-Edge/Wake-Vision/croissant https://huggingface.co/api/datasets/Harvard-Edge/Wake-Vision-Train-Large/croissant
+Appendix S Car and Bird Datasets
+Table 10 compares the number of images as well as the label error rate for the car and bird datasets, when generated using both our methodology and the VWW one.
+All datasets are balanced, i.e., have the same number of images for the target and background class.
+Label error rates have been manually computed on a subset of 500 randomly selected images for each generated dataset.
+Examples of positive class images from the Wake Vision test splits are shown in Figure 11
+Table 10: Raw data of Car and Bird datasets.
+Methodology # of Images Label Error Rate
+Car
+Wake Vision (Ours) 157,668 4.8%
+VWW [7] 12,902 6.4%
+Bird
+Wake Vision (Ours) 31,642 0.6%
+VWW [7] 1,148 6.6%
+(a) Wake Vision “Car” samples
+(b) Wake Vision “Bird” samples
+Figure 11: Positive class examples from the test splits of our Wake Vision generated datasets for “Car” and “Bird” objects, demonstrating our ability to easily prototype large-scale TinyML vision datasets (Section 3.5).
+Appendix T Wake Vision Challenge Results
+Submissions to the model-centric track have been trained using the same recipe adopted in Section 4.3.
+Table 11 provides input shape, Flash, RAM, and MACs for each model presented in Section 4.5 as well as the mean and standard deviation of test accuracies for the Wake Vision dataset.
+RAM and Flash have been measured using the “stm32tflm” utility of X-CUBE-AI 8.1.0, whereas MACs with “stm32ai”.
+Table 11: Raw data of model-centric track.
+Model input RAM Flash MACs Test Acc
+.tflite volume [kiB] [kiB] [MM] [%]
+anas-benalla (50,50,3) 48.5 104.55 4,899,292 $79.7\pm 0.28$
+samy (80,80,3) 73.5 34.55 5,718,064 $79.5\pm 0.61$
+benx13 (48,48,3) 63.5 94.95 1,693,474 $79.6\pm 0.22$
+cezar (50,50,3) 245.5 57.23 20,435,868 $79.6\pm 0.42$
+mohammad_hallaq (80,80,3) 129.5 128.32 3,887,331 $77.3\pm 0.5$
+ymac (50,50,3) 34.5 27.77 431,985 $76.7\pm 0.51$
+apighetti (50,50,3) 48.5 278.36 693,818 $74.3\pm 74.3$
+Table 12 shows test accuracy, precision, and recall of a MobileNetV2 having a width modifier of 0.25, trained on the original Wake Vision large split and on the same split enhanced with the technique proposed by the winner of the data-centric track [4].
+The training recipe is the same reported in Section 4.3.
+Table 12: Main improvement of the data-centric track.
+Metrics Train Large Split
+Original Enhanced
+MobileNetV2_0.25
+Test Accuracy $82.3\pm 0.2$ $87.1\pm 0.81$
+Precision $86.4\pm 1.49$ $87.6\pm 1.78$
+Recall $76.7\pm 1.83$ $86.5\pm 1.32$
+Appendix U Training Resources
+A desktop computer featuring a 13th Gen Intel® Core™ i9-13900K × 32, a 32 GB of RAM, and an NVIDIA GeForce RTX 4080 with 16 GB of VRAM has been employed to train the TinyML models.
+Appendix V Datasheet for Wake Vision
 This document is based on Datasheets for Datasets by Gebru et al. [19]. Please see the most updated version here.
@@
 Who created this dataset (e.g., which team, research group) and on behalf of which entity (e.g., company, institution, organization)?
-These details will be provided after the review period to avoid breaking anonymity.
+The authors of this paper created the dataset (Harvard Edge Computing Lab, Technical University of Denmark, and Useful Sensors).
 What support was needed to make this dataset?
@@
 If there is an associated grant, provide the name of the grantor and the grant name and number, or if it was supported by a company or government agency, give those details.)
-These details will be provided after the review period to avoid breaking anonymity.
+The Google TPU Research Cloud program partially supported this work via cloud computing credits.
+The PhD stipend of Emil Njor was supported by the Innovation Fund Denmark DIREC project (9142-00001B).
+Emil Njor’s external research stay was furthermore supported by Fulbright Denmark, Stibo Fonden, Thomas.
+B.
+Thriges Fond, Otto Mønsteds Fond and Kaj og Hermilla Ostenfelds Fond.
+This work was also support by the National Science Foundation (NSF) and the Semiconductor Research Corporation (SRC).
 Any other comments?
@@
 Is the software used to preprocess/clean/label the instances available? If so, please provide a link or other access point.
-Yes, we open-source our filtering code. https://anonymous.4open.science/r/Wake_Vision-DD19/README.md
+Yes, we open-source our filtering code. https://github.com/harvard-edge/Wake_Vision
 Any other comments?
@@
 These errors are due to the labeling and filtering process.
-We mitigate the impact of these errors by manually labeling the validation and test sets as well as study their impact for TinyML models in Section 4.4.
+We mitigate the impact of these errors by manually labeling the validation and test sets as well as study their impact for TinyML models in Section C.
 Are there tasks for which the dataset should not be used? If so, please provide a description.
@@
 How can the owner/curator/manager of the dataset be contacted (e.g., email address)?
-Removed during review to maintain anonymity.
+At the time of writing, the authors can be contacted at cbanbury@g.harvard.edu. Once maintenance is transferred to the Edge AI Foundation, there will be a point of contact there.
 Is there an erratum? If so, please provide a link or other access point.
@@
 Will older versions of the dataset continue to be supported/hosted/maintained? If so, please describe how. If not, please describe how its obsolescence will be communicated to users.
-Yes, we will continue to host older versions of the dataset. Details will be provided after review to preserve anonymity.
+Yes, we will continue to host older versions of the dataset. This is an inherent feature of the Harvard Dataverse.
 If others want to extend/augment/build on/contribute to the dataset, is there a mechanism for them to do so?